### v1

# Quantitative surgery and total mean curvature

Abstract.
We develop quantitative surgery, which extends the classical constructions of Gromov–Lawson and Lawson–Michelsohn. As an application, we prove a conjecture of Gromov on the total mean curvature of fill-ins.
1. Introduction
Let $(M,g_{M})$ be a closed Riemannian manifold, possibly disconnected, with scalar curvature $\mathrm{scal}_{g_{M}}\colon M\to\mathbb{R}$. A fill-in of $(M,g_{M})$ is a compact Riemannian manifold $(\Omega,g_{\Omega})$ such that $(\partial\Omega,g_{\partial\Omega})$ is isometric to $(M,g_{M})$, where $g_{\partial\Omega}$ denotes the metric induced by $g_{\Omega}$ on $\partial\Omega$. We write $H_{g_{\Omega}}\colon\partial\Omega\to\mathbb{R}$ for the mean curvature of $\partial\Omega\subset(\Omega,g_{\Omega})$, where we adopt the convention that the mean curvature of the closed unit ball in flat Euclidean space $\mathbb{R}^{n}$ is equal to n1n-1.
Gromov formulated the following conjecture:
Conjecture 1.1 ([18, p. 232]).
Let $(M,g_{M})$ be a closed Riemannian manifold. Then for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda(M,g_{M},\sigma)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$, one has
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma).$
The following is our main result.
Theorem 1.2.
Let $(M,g_{M})$ be a closed Riemannian manifold admitting a spin structure. Then, for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying:
  • $\mathrm{scal}_{g_{\Omega}}\geq\sigma$,
  • $H_{g_{\Omega}}\geq 0$,
  • $\Omega$ admits a spin structure in case $\dim M\geq 7$,
one has:
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma).$
Theorem 1.2 has been proven before in the following cases: By Shi–Wang–Wei [25] if $M$ is diffeomorphic to the $n$-sphere, based upon Shi–Tam’s proof for the positivity of the Brown–York mass [24] and the hyperbolic positive mass theorem (PMT) [29, 14, 2]; by Wang [30] for tori $T^{n}$, using Brendle–Hung’s [11] work on the Horowitz–Myers PMT; and by Bär in [5] concurrently under the additional assumption that $\Omega$ admits a spin structure.
Our approach, which differs from the previously mentioned works, relies on performing surgeries of sufficiently high codimension to reduce Theorem 1.2 to the case where $M$ is a sphere, which is covered in [25]. In particular, no spin assumption on the fill-in manifold $\Omega$ is required provided the hyperbolic positive mass theorem holds for orientable manifolds in dimension $\dim M+1$.
Remark 1.3.
The spin assumption on $M$ in Theorem 1.2 is used solely to guarantee the existence of the required surgeries transforming $M$ into a sphere, whereas the spin assumption on $\Omega$ is only needed for the validity of the hyperbolic positive mass theorem. In particular, our argument does not rely on index theory for a spin Dirac operator on either $M$ or $\Omega$. In fact, if 1.1 is proven for one simply connected, totally non-spin manifold in dimension $n$, e.g. $\mathbb{CP}^{2}\times S^{n-4}$ or the non-trivial linear $S^{n-2}$-bundle over $S^{2}$, then our proof of Theorem 1.2 shows that 1.1 holds true for every orientable manifold $M$ of dimension $n$.
1.1. Quantitative surgery
Our main tool for establishing Theorem 1.2 is a new quantitative surgery procedure. This encompasses and strengthens the classical results of Gromov–Lawson [17] and Lawson–Michelsohn [20].
To explain this construction, let $S$ be a compact submanifold in $(M,g_{M})$ of codimension $k$, and let $\varrho>0$ be a small parameter. A Gromov–Lawson handle of size $\varrho$ is a hypersurface $\Sigma_{\varrho}$ contained in $(M\times[0,1],dt^{2}+g_{M})$ which coincides with $M=M\times\{0\}$ outside the $2\sqrt{\varrho}$-tubular neighbourhood of $S$ in $M$, becomes cylindrical of width $\varrho^{4}$ at height $2\varrho$ and is invariant under the fibrewise $O(k)$-symmetry of the normal bundle of $S$ in $M$, see Fig. 1.
Let $\mathrm{scal}_{\varrho}$ be the scalar curvature and let $H_{\varrho}$ be the mean curvature of $\Sigma_{\varrho}$. Furthermore, let $s$ be the minimum of the scalar curvature of $(M,g_{M})$ on $S$. The crucial properties of our construction are:
  • $(H_{\varrho})_{-}$ converges to zero as $\varrho\to 0$.
  • $(\mathrm{scal}_{\varrho}-s)_{-}$ converges to zero as $\varrho\to 0$.
Consequently, by performing a conformal transformation arbitrarily close to the identity, we can ensure that the mean curvature is positive for sufficiently small $\varrho$.
The detailed construction in the special case where $S$ is a point in flat Euclidean space $\mathbb{R}^{n}$ is carried out in Section 4. For general $M$ and $S$, error terms arise, which will be controlled in Section 5.
[…]
Figure 1. A cross-section of a family of Gromov–Lawson handles $\Sigma_{\varrho}$ attached to $\Omega$ along $M=\partial\Omega$. The origin corresponds to the intersection of the cross-section with $S$.
Our construction moreover provides precise control between handles of different sizes $\varrho$. In particular, the heuristic observation (see Figure 1) that the metrics on $\Sigma_{\varrho}$ are monotonically increasing in $\varrho$ can be made rigorous, while maintaining a uniform lower bound on the scalar curvature of $\Sigma_{\varrho}$. This feature is crucially exploited in the proof of Theorem 1.2, as described below. Our computations are fully explicit, and the handles are constructed from first principles, yielding an independent and unifying proof of both [17, Theorem A] and [20, Theorem 3.1].
Owing to these explicit geometric bounds, our results also encompass variants of the conclusions of several earlier works, including [26, 27, 7, 15, 9, 8].
1.2. Proof outline
By topological arguments, we can assume that we can pass from $M$ to the sphere $S^{n}$ through surgeries of codimension $k\geq 3$. At the same time, we pass from a fill-in of $M$ to a fill-in of $S^{n}$ via the corresponding handle attachments. Working with handles of sufficiently small width, this ensures that the total mean curvature of the original fill-in of $M$ transfers to the total mean curvature of a fill-in of $S^{n}$, up to a small error. An application of Shi–Wang–Wei’s total mean curvature estimate to $S^{n}$ then concludes the proof of Theorem 1.2.
Relying on the classical surgery result of Lawson–Michelsohn [20], this strategy can be carried out as stated when restricting to fill-ins $\Omega$ of $M$ with $H_{g_{\Omega}}\geq\mu$ for some fixed $\mu>0$ and working with a constant $\Lambda$ that depends on $M$, $g_{M}$, $\sigma$ and $\mu$. We present the proof of this special case of Theorem 1.2 in Section 2.
In the general case, however, there is a significant caveat. If the point-wise lower mean curvature bound of different fill-ins $\Omega$ is not uniformly bounded below by a positive constant, the handles have to be chosen thinner and thinner to ensure a non-negative mean curvature bound for the resulting fill-in of $S^{n}$. This leads to infinitely many metrics on the resulting $S^{n}$ and consequently, Shi–Wang–Wei’s result does not yield a uniform total mean curvature bound. At this stage, it is therefore unclear whether the constant $\Lambda$ can be chosen independently of $\mu$.
This difficulty will be overcome by the following strategy (see Section 3): Shi–Wang–Wei’s upper bound can be chosen uniformly for a family of metrics on $S^{n}$, provided that each member of the family can be connected to a fixed metric on $S^{n}$ along a monotonically increasing path of Riemannian metrics with lower scalar curvature bounds that are independent of the given metric in the family (see Proposition 3.2 for the precise statement). Our quantitative surgery construction in Section 1.1 provides the required control of the surgery handles. Proceeding by induction on the number of surgery steps finishes the proof in the general case.
Theorem 1.2 also allows us to define a new quasi-local mass with several particularly appealing properties. This will be discussed in Section 7.
Acknowledgements: BH and SH are grateful to the University of Augsburg, to Columbia University, the Lonavala Geometry Festival, and to the Chennai Mathematical Institute for their hospitality and ideal working conditions. BH acknowledges support by the DFG-SPP 2026 “Geometry at Infinity”. SH was supported in part by an AMS–Simons Travel Grant (award no. 210022).
2. Total mean curvature bounds with fixed positive lower mean curvature bound
Here we will prove the following weak version of Theorem 1.2.
Theorem 2.1.
Let $(M,g_{M})$ be a closed Riemannian manifold admitting a spin structure. Let $\mu>0$.
Then, for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma,\mu)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying:
  • $\mathrm{scal}_{g_{\Omega}}\geq\sigma$,
  • $H_{g_{\Omega}}\geq\mu$,
  • if $\dim M\geq 7$, then $\Omega$ is a spin manifold and there is a spin structure preserving isometry $(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$,
one has:
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma,\mu).$
In the remainder of this section, we prove this result. Put $n:=\dim M$. Let $\sigma\in\mathbb{R}$ and let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ satisfying the properties of Theorem 2.1. For simplicity of exposition, we assume $(M,g_{M})=\partial(\Omega,g_{\Omega})$. We start with some preliminary topological modifications. By taking products of $M$ and of $\Omega$ with standard round spheres, we can assume without loss of generality that $n\geq 5$. Upon passing to the oriented double cover of $(\Omega,g_{\Omega})$ and to the trivial double cover of $(M,g_{M})$, we can furthermore assume that $\Omega$ is orientable.
The kernel of the forgetful map from spin bordism to oriented bordism is a $2$-torsion group by [1, Corollary 2.6]. Replacing $M$ by $M\sqcup M$ and $\Omega$ by $\Omega\sqcup\Omega$, we can hence assume that $M$ bounds a connected spin manifold $B$. Furthermore, if $\dim M\geq 7$, then the given spin structure on $\Omega$ extends over $\Omega\cup_{M}B$.
Removing a small open disc from the interior of $B$, we obtain a compact connected spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{1}(W)$ and $\pi_{2}(W)$ by surgeries in the interior of $W$ (and using $n\geq 5$), we can assume that $W$ is $2$-connected. In particular, the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $2$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $3$.
Let $\ell\geq 0$ be the number of handles in this handle decomposition of $W$. Using this handle decomposition, $S^{n}$ is obtained from $M$ by $\ell$ spin surgeries of codimension at least $3$. The proof of Theorem 2.1 is by induction on $\ell$. The case $\ell=0$ is covered by [25, Theorem 4.1 and Remark 4.2]. Now assume by induction that Theorem 2.1 holds for some $\ell\geq 0$ and assume that the number of codimension $3$-handles in $W$ is equal to $\ell+1$.
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$ be the surgery datum for the first codimension $3$-handle in $W$ and let $\widetilde{M}$ be obtained from $M$ by surgery along $\psi$. Note that the induction hypothesis applies to $\widetilde{M}$. There exist smooth embeddings $\psi_{i}\colon S^{n-k}\times D^{k}\hookrightarrow M$, $i=1,2$, that are both isotopic to $\psi$ and have disjoint images, see Fig. 2. The point of this “doubling construction” is to estimate the total mean curvature after the surgery by a pigeonhole argument, see (2.2) below.
[…]
Figure 2. The disjoint embeddings $\psi_{1}$ and $\psi_{2}$ which are isotopic to $\psi$ as small displacements.
Choose a smooth collar $M\times[0,1)\hookrightarrow W$ of $M\subset\partial W$ in $W$. With respect to this collar, we identify $M\times[0,1)$ with a smooth submanifold of $W$. Choose a Riemannian metric $h$ on $W$ which restricts to the metric
(2.1) $dt^{2}+\left(1+\tfrac{t}{n-1}\mu\right)g_{M}\text{ on }M\times[0,1)\subset W,$
where $g_{M}$ and $\mu$ are from the assumptions in Theorem 2.1. Note that $(W,h)$ only depends on $(M,g_{M})$, but not on the fill-in $(\Omega,g_{\Omega})$. The second fundamental form of $(W,h)$ along $M\times\{0\}$ with respect to the exterior normal is equal to $\tfrac{\mu}{2(n-1)}\cdot g_{M}$.
From now on, let $i\in\{1,2\}$. By [20, Theorem 3.1], we can find a submanifold $T_{i}\subset W$ which intersects $M$ in $\mathrm{im}(\psi_{i})$, which is isotopic to the handle attached along the surgery datum $\psi_{i}$, such that $\Omega\cup T_{i}$ is smooth and such that the mean curvature
$\widetilde{M}_{i}:=\partial\left(\Omega\cup T_{i}\right)$
is bounded below by some constant $0<\mu_{i}<\tfrac{\mu}{2}$, see Fig. 3. Here $\partial\left(\Omega\cup T_{i}\right)$ is regarded as a smooth submanifold of $W$ and the mean curvature is computed with respect to the metric $h$ and the normal pointing inside $W$. Note that $\widetilde{M}_{i}$ carries a spin structure induced by the one on $W$.
[…]
Figure 3. The cobordism $W$ consisting of one handle and the embedded handle
Let $\widetilde{g}_{i}$ be the metric on $\widetilde{M}_{i}$ induced by $h$. Note that the induction hypothesis applies to $(\widetilde{M}_{i},\widetilde{g}_{i})$, as $\widetilde{M}_{i}$ is diffeomorphic to $\widetilde{M}$.
Let $U_{i},V_{i},Z_{i}$, $i=1,2$, be open neighborhoods of $\mathrm{im}(\psi_{i})\subset M$ such that $\overline{U}_{i}\subset V_{i}$, $\overline{V}_{i}\subset Z_{i}$ and $\overline{Z}_{1}\cap\overline{Z}_{2}=\emptyset$.
By [4, Theorem 3.7], we can smoothly deform the metric $g_{\Omega}$ on $\Omega$ to a new metric on $\Omega$, called $g_{i}$ for short, with the following properties (see Fig. 4):
  • The metrics on $\partial\Omega$ induced by $g_{i}$ are equal to $g_{M}$,
  • $H_{g_{i}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$,
  • $H_{g_{i}}\geq\tfrac{\mu}{2}$ everywhere,
  • On $V_{i}$, the second fundamental form of $g_{i}$ is given by
    $\mathrm{II}_{g_{i}}=\tfrac{\mu}{2(n-1)}g_{M},$
  • $\mathrm{scal}_{g_{i}}>\sigma-1$.
[…]
Figure 4. Deformation of $g_{\Omega}$ near $\partial\Omega$ (grey region)
Let $\delta>0$ and let $\mathscr{C}:=\partial\Omega\times(-\delta,0]\subset\Omega$ be an open collar of $\partial\Omega$ in $\Omega$ induced by the normal exponential map along $\partial\Omega$ for $g_{i}$. Put $\mathscr{V}_{i}:=V_{i}\times(-\delta,0]\subset\mathscr{C}$. Equip the topological manifold
$\mathscr{V}_{i}\cup_{V_{i}\times 0}(V_{i}\times[0,1))=V_{i}\times(-\delta,1)\subset\Omega\cup_{M}W$
with the product smooth structure. Set $\widetilde{\Omega}_{i}:=\Omega\cup_{M}T_{i}\subset\Omega\cup_{M}W$. We have
$\widetilde{\Omega}_{i}\subset\Omega\cup(U_{i}\times(-{\delta},1))\cup{\rm int}(W).$
By construction, the smooth structures on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-{\delta},1))$ and on $\widetilde{\Omega}_{i}\cap{\rm int}(W)$ are compatible with each other. Hence, we obtain an induced smooth structure on $\widetilde{\Omega}_{i}$. We have $\partial\widetilde{\Omega}_{i}=\widetilde{M}_{i}$ as smooth manifolds. Furthermore, if $n\geq 7$, then $\widetilde{\Omega}_{i}$ is equipped with a spin structure inducing the one on $\widetilde{M}_{i}$.
We claim that there exists a metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ that satisfies
  1. (i)
    $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$,
  2. (ii)
    $H_{g_{\widetilde{\Omega}_{i}}}\geq\mu_{i}$ everywhere,
  3. (iii)
    $\mathrm{scal}(g_{\widetilde{\Omega}_{i}})>{\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$.
Let us for now assume this metric $g_{\widetilde{\Omega}_{i}}$ exists and let $g_{\widetilde{M}_{i}}$ denote the induced metric on $\widetilde{M}_{i}$. Using $Z_{1}\cap Z_{2}=\emptyset$ and properties (i) and (ii), we obtain
(2.2) $\begin{split}\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq{}&\int_{\partial\Omega\setminus Z_{1}}H_{g_{\Omega}}{\rm dvol}_{g_{M}}+\int_{\partial\Omega\setminus Z_{2}}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\\ \leq{}&\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\widetilde{M}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\widetilde{M}_{2}}}\\ \leq{}&2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\widetilde{M}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\widetilde{M}_{2}}}\right\}.\end{split}$
By induction, for $i\in\{1,2\}$, Theorem 2.1 holds for the manifold $(\widetilde{M}_{i},\widetilde{g}_{i})$, the lower scalar curvature bound ${\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ and the lower mean curvature bound $\mu_{i}$. This concludes the induction step.
It remains to construct the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. For this, we first note that the restriction $h|_{V_{i}\times[0,1)}$ extends to a smooth metric $h^{\prime}$ on $V_{i}\times(-\delta,1)$ and that the $1$-jets of the metrics $g_{i}$ and $h^{\prime}$ along $V_{i}\times\{0\}\subset V_{i}\times(-\delta,1)$ coincide by the construction of $g_{i}$. Let $\chi\colon\partial\Omega\to[0,1]$ be a smooth cut-off function that is equal to $1$ on $U_{i}$ and equal to 0 on $\partial\Omega\setminus V_{i}$. Consider the smooth path of symmetric $(0,2)$-tensor fields $\psi\colon[0,1]\to C^{\infty}(T^{*}\mathscr{C}\otimes T^{*}\mathscr{C})$,
$\psi(t)(x,r):=(1-\chi(x)\cdot t)g_{i}+\chi(x)\cdot t\cdot h^{\prime}.$
Since $\chi(x)=0$ for $x\in\partial\Omega\setminus V_{i}$, this is well-defined. Possibly after passing to a smaller $\delta$ (hence to a smaller $\mathscr{C}$), the path $\psi$ consists of Riemannian metrics on $\mathscr{C}$, see Fig. 5.
[…]
Figure 5. The local deformation $\psi$ for $t=1$.
By the choice of $\chi$, the $1$-jet of $\psi(t)$ along $\partial\Omega$ is independent of $t$. In particular, the scalar curvature of $\psi(t)$ along $\partial\Omega$ is given by
$\mathrm{scal}_{\psi(t)}=(1-\chi(x)\,t)\mathrm{scal}_{g_{i}}+\chi(x)\,t\cdot\mathrm{scal}_{h}>\min\{\sigma,\mathrm{scal}_{h}\}-1.$
The last lower bound is independent of the fill-in $(\Omega,g)$. By continuity, possibly after passing to a smaller $\delta$, which may depend on $(\Omega,g_{\Omega})$, we have, on $\mathscr{C}$,
$\mathrm{scal}_{\psi(t)} >\min\{\sigma,\mathrm{scal}_{h}\}-1,\quad t\in[0,1].$
We apply the local flexibility lemma [3, Theorem 1.2] to the local deformation $\psi$ on the open neighborhood $\mathscr{C}$ of $\partial\Omega$ in $\Omega$ and to the second order open partial differential relation of being a Riemannian metric with $\mathrm{scal}>\min\{C,\mathrm{scal}_{h}\}-1$ on symmetric $(0,2)$-tensor fields in $C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$. This results in a smooth path $\Psi\colon[0,1]\to C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$ of Riemannian metrics on $\Omega$ starting at $g_{i}$, and there exists some $0<\delta^{\prime}<\delta$, such that the restriction of $\Psi(t)$ to $\partial\Omega\times(-\delta^{\prime},0]$ coincides with $\psi(t)$, see Fig. 6.
[…]
Figure 6. The end of the local deformation $\Psi(1)$ obtained by applying the local flexibility lemma to the deformation $\psi$.
We set $g_{\widetilde{\Omega}_{i}}\coloneqq\Psi(1)$. Observe, that the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, the metric $h^{\prime}$ on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-\delta^{\prime},1))$ and the metric $h$ on $\widetilde{\Omega}_{i}\cap{\rm int}(W)$ are compatible with each other. Hence, the union of these metrics defines a smooth metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. By construction, $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{i}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$. Furthermore, we have $\mathrm{scal}_{g_{\widetilde{\Omega}_{i}}}>{\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ for $i\in\{0,1\}$. Thus, $g_{\widetilde{\Omega}_{i}}$ satisfies the properties required above.
3. Uniform upper bound of total mean curvature for families of metrics
Let $(M,g_{M})$ be a closed spin Riemannian manifold for which Theorem 1.2 holds. In this section we will prove that Theorem 1.2 still holds, with a uniform upper bound $\Lambda$, for a certain family of metrics on $M$. The proof is based on arguments in [25].
Definition 3.1.
Let $s\in\mathbb{R}$ such that $\mathrm{scal}_{g}>s$. We define $\mathscr{R}_{g_{M},s}(M)$ as the subset of Riemannian metrics $g$ on $M$ with the following property: There exists a smooth path $\lambda_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ of Riemannian metrics starting at $g$ and ending at $g_{M}$ such that, for all $t\in[0,1]$, we have
  1. (i)
    $\lambda^{\prime}_{g}(t)>0$ in the sense of symmetric $(0,2)$-tensor fields.
  2. (ii)
    $\mathrm{scal}_{\lambda_{g}(t)}>s$ on $M$.
Proposition 3.2.
For any $s\in\mathbb{R}$, there exists $\Lambda=\Lambda(M,g_{M},\sigma,s)\in\mathbb{R}_{+}$ such that the following holds: If $g\in\mathscr{R}_{g_{M},s}$ and if $(\Omega,g_{\Omega})$ is a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$, then
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma,s).$
Proof.
Let $g\in\mathscr{R}_{g_{M},s}$. Let $\lambda_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ be a smooth path from $g$ to $g_{M}$ that satisfies properties (i) and (ii). Consider the smooth metric $\overline{g}$ on $M\times[0,1]$ defined by
$\overline{g}=dt^{2}+\lambda_{g}(t).$
Let $\overline{A}_{t}$ be the second fundamental form and $\overline{H}_{t}=\operatorname{tr}_{\lambda_{g}(t)}\overline{A}_{t}$ be the mean curvature of $\Sigma_{t}:=M\times\{t\}\subset(M\times[0,1],\overline{g})$ with respect to the unit normal field $\frac{\partial}{\partial t}$. By (i), we have $\overline{A}_{t}>0$, hence $\overline{H}_{t}>0$ and $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$ for $t\in[0,1]$
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Applying [4, Proposition 3.8], we can increase the mean curvature $H_{g_{\Omega}}$ slightly, while replacing $\sigma$ by $\sigma-1$. Hence, we can asume that $H_{g_{\Omega}}>0$. On $M\times[0,1]$, consider the Bartnik-type parabolic PDE
$\overline{H}_{t}\frac{\partial u}{\partial t}$ $=u^{2}\Delta_{\lambda_{g}(t)}u+\frac{1}{2}(s-\mathrm{scal}_{\lambda_{g}(t)})u^{3}+\frac{1}{2}(\mathrm{scal}_{\lambda_{g}(t)}-\mathrm{scal}_{\overline{g}})u$
$u(\cdot,0)$ $=\frac{\overline{H}_{0}}{H_{g_{\Omega}}}>0.$
Since, by (ii), we have $s-\mathrm{scal}_{\lambda_{g}(t)}<0$ for $t\in[0,1]$, this equation has a positive solution $u\colon M\times[0,1]\to\mathbb{R}_{>0}$. Consider the Riemannian metric $\widetilde{g}=u^{2}dt^{2}+\lambda_{g}(t)$ on $M\times[0,1]$.
Let $H_{t}=u^{-1}\overline{H}_{t}>0$ be the mean curvature on $\Sigma_{t}$ induced by $\widetilde{g}$ with respect to the unit normal field $u^{-1}\frac{\partial}{\partial t}$. Note that $H_{0}=H_{g_{\Omega}}$. By the computation in the proof of [25, Theorem 4.1] and using that $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$, we have
$\frac{\partial}{\partial t}\int_{\Sigma_{t}}H_{t}\,\rm{dvol}_{\lambda_{g}(t)}>0.$
In particular,
(3.1) $\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\lambda_{g}(1)}>\int_{\Sigma_{0}}H_{0}{\rm dvol}_{\lambda_{g}(0)}=\int_{M}H_{g_{\Omega}}{\rm dvol}_{g}.$
Furthermore, we have $\mathrm{scal}_{\widetilde{g}}={\rm const.}=s$.
Consider the topological manifold
$\widetilde{\Omega}:=\Omega\bigcup_{M\times\{0\}\approx\partial\Omega}M\times[0,1].$
It carries a unique smooth structure induced by the smooth structures on $\Omega$ and $M\times[0,1]$ and the collar neighborhoods along $\partial\Omega\subset\Omega$ and $M\times\{0\}\subset M\times[0,1]$ induced by $g_{\Omega}$ and $\widetilde{g}$. If $n\geq 7$, the manifold $\widetilde{\Omega}$ carries a spin structure restricting to the given spin structure on $\Omega$. By [4, Theorem 4.11] and by the choice of $u(\cdot,0)$, the mean convex singularity at the gluing region $\partial\Omega=M\times\{0\}$ can be smoothened, that is, there exists a smooth Riemannian metric $g_{\widetilde{\Omega}}$ on $\widetilde{\Omega}$ which is equal to $\widetilde{g}$ near $M\times\{1\}$ and whose scalar curvature is bounded below by $\min\{s,\sigma\}-1$.
[…]
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g}<\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\lambda_{g}(1)}=\int_{\partial\widetilde{\Omega}}H_{g_{\widetilde{\Omega}}}{\rm dvol}_{g_{M}}.$
Now the claim of Proposition 3.2 follows from the assumption on $(M,g_{M})$, with $\sigma$ replaced by ${\rm min}\{\sigma,s\}-1$. ∎
4. Surgery bending in Euclidean space
In this section we construct hypersurfaces $\overline{\Sigma}_{\varrho}$ in Euclidean space which we will later use to define surgery bendings with appropriate curvatures. For this purpose we equip $B_{1}^{n-k}(0)\times[0,1]\subset\mathbb{R}^{\,n-k+1}$ with the flat metric $g_{\mathbb{R}^{n-k+1}}=\mathrm{d}x_{1}^{2}+\dots+\mathrm{d}x_{n-k}^{2}+\mathrm{d}t^{2}$. Let us now describe the construction of $\overline{\Sigma}_{\varrho}$.
Construction 4.1.
Let $\varrho\in(0,\frac{1}{10}]$ be fixed and consider the function
$\xi_{\varrho}(r)\coloneqq-\frac{4}{3}\varrho\left(\sqrt{r}-\varrho^{2}\right)^{3/2}-4\varrho^{3}\sqrt{\sqrt{r}-\varrho^{2}}$
from Lemma A.1. The handle $\overline{\Sigma}_{\varrho}$ will essentially be given as the graph of the function $(x_{1},\dots,x_{n-k})\mapsto\xi_{\varrho}(\sqrt{x_{1}^{2}+\dots+x_{n-k}^{2}})+a_{\varrho}$, where $a_{\varrho}\coloneqq-\xi_{\varrho}(2\sqrt{\varrho})$. However, there are a few adjustments to make: Since we want to glue this hypersurface into a given manifold, we need $\overline{\Sigma}_{\varrho}$ to agree with $B_{1}^{n-k}(0)\times\{0\}$ for large radii, and we need it to be cylindrical near the top.
In order to achieve this, we fix once and for all a cutoff-function $\eta\colon\mathbb{R}\to[0,1]$, that is $\eta\equiv 1$ on $(-\infty,0]$, $\eta\equiv 0$ on $1,\infty)$, and $2\leq\eta^{\prime}\leq 0$, $|\eta^{\prime\prime}|\leq 10$. We now define (see Fig. 7) $\eta_{\varrho}(r)=\eta(\tfrac{r-\sqrt{\varrho}}{\sqrt{\varrho}})$ and
$f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr),$
and we consider the graph of the function
$B_{1}^{n-k}(0)\setminus B_{\varrho^{4}}^{n-k}(0)\to\mathbb{R},\qquad(x_{1},\dots,x_{n-k})\mapsto f_{\varrho}\left(\sqrt{x_{1}^{2}+\dots+x_{n-k}^{2}}\right).$
[…]
Figure 7. The function ${f}_{\varrho}$.
Note that the derivative of $f_{\varrho}$ converges to -$\infty$ for $r\to\varrho^{4}$ and therefore, we can attach a straight cylinder of radius $\varrho^{4}$ to this graph at the level $t=a_{\varrho}=f(\varrho^{4})$ and obtain a $C^{1}$-hypersurface which is smooth away from $t=a_{\varrho}$.
[…]
Figure 8. The hypersurface $\overline{\Sigma}_{\varrho}$ as the graph of a spherically symmetric function.
Remark 4.2.
Let us motivate the choice of the function $\xi_{\varrho}$. The graph of the functions $\ell_{\varrho}(r)=-2\sqrt{2\varrho}\sqrt{r-2\varrho}$ give rise to spherically symmetric surfaces satisfying
$2\lambda_{1}=\lambda_{2}=\dots=\lambda_{n-k}$
for their principal curvatures, cf. Lemma A.3. In $n=3$ these are the (spatial) Schwarzschild metrics of mass $m=\varrho$. In this case the scalar curvature vanishes which does not allow us absorb the error terms appearing in the handle construction. Therefore, we solve the ODE
$4\lambda_{1}=\lambda_{2}=\dots=\lambda_{n-k}$
which leads to a spherically symmetric surface with strictly positive scalar curvature. The corresponding graph function is given precisely by $\xi_{\varrho}$.
Let $\overline{\nu}_{\varrho}$ be the unit normal vector field of $\overline{\Sigma}_{\varrho}$ and denote by $\overline{\nu}_{\varrho}^{\intercal}$ the projection of $\overline{\nu}_{\varrho}$ onto the tangent space of the slices $B_{1}^{n-k}(0)\times\{t\}$, i.e.,
(4.1) $\overline{\nu}^{\intercal}_{\varrho}=\overline{\nu}_{\varrho}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle\partial_{t}.$
Note that $\langle\overline{\nu}_{\varrho},\partial_{t}\rangle\geq 0$ since $\overline{\Sigma}_{\varrho}$ is a graph. We have the following properties.
Proposition 4.3.
The hypersurfaces $\overline{\Sigma}_{\varrho}$ from 4.1 satisfy the following properties.
  1. (i)
    $\overline{\Sigma}_{\varrho}$ is $\mathrm{O}(n-k)$-invariant and satisfies
    $\overline{\Sigma}_{\varrho}\cap(B_{1}^{n-k}(0)\times[2\varrho,1])$ $=\partial B_{\varrho^{4}}^{n-k}(0)\times[2\varrho,1],$
    $\overline{\Sigma}_{\varrho}\cap\Bigl(\bigl(B_{1}^{n-k}(0)\setminus B_{2\sqrt{\varrho}}^{n-k}(0)\bigr)\times[0,1]\Bigr)$ $=\bigl(B_{1}^{n-k}(0)\setminus B_{2\sqrt{\varrho}}^{n-k}(0)\bigr)\times\{0\}.$
  2. (ii)
    For $r\in[\varrho^{4},\sqrt{\varrho}]$ and away from $t=a_{\varrho}$ we have
    $H(\overline{\Sigma}_{\varrho})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho})|\quad\text{if }n-k\geq 2$
    $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}\quad\text{if }n-k\geq 3.$
  3. (iii)
    For $r\in[\varrho^{4},\varrho^{2}]$, $n-k\geq 3$ and away from $t=a_{\varrho}$, we have
    $\min\left\{H(\overline{\Sigma}_{\varrho})^{2},\mathrm{scal}(\overline{\Sigma}_{\varrho})\right\}>\tfrac{1}{2}\varrho^{-3}.$
For the final two properties let $\overline{C}\coloneqq 18(n-k)+{432}$.
  1. (iv)
    For $r\geq\varrho^{2}$, we have $|\overline{\nu}_{\varrho}^{\intercal}|<\overline{C}\,\sqrt{\varrho}$.
  2. (v)
    For $r\geq\sqrt{\varrho}$, we have $|A(\overline{\Sigma}_{\varrho})|<\overline{C}\,\varrho^{3/8}$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>-\overline{C}\,\varrho^{3/4}$.
Proof.
The first assertion follows directly from the construction, the only thing to verify is that $a_{\varrho}<2\varrho$ for $\varrho\leq\frac{1}{10}$:
$a_{\varrho}=-\xi_{\varrho}(2\sqrt{\varrho})=\frac{4}{3}\varrho\underbrace{\left(\sqrt{2\sqrt{\varrho}}-\varrho^{2}\right)^{3/2}}_{<1}+4\varrho^{3}\underbrace{\sqrt{\sqrt{2\sqrt{\varrho}}-\varrho^{2}}}_{<1}<\frac{4}{3}\varrho+\frac{4}{100}\varrho<2\varrho.$
Concerning (ii), we note that on this inner annulus the cutoff function $\eta_{\varrho}$ equals $1$, so $f_{\varrho}=\xi_{\varrho}+a_{\varrho}$ and therefore $f_{\varrho}^{\prime}=\xi_{\varrho}^{\prime}$ as well as $f_{\varrho}^{\prime\prime}=\xi_{\varrho}^{\prime\prime}$. By Lemma A.1, we have
$4\,\xi_{\varrho}^{\prime\prime}=-\frac{1}{r}\,\xi_{\varrho}^{\prime}(1+\xi_{\varrho}^{\prime 2}).$
Combined with Lemma A.3, gives
$\lambda_{2}=\dots=\lambda_{n-k}=-4\,\lambda_{1}.$
on the graph of $f_{\varrho}$ on $\varrho^{4},\sqrt{\varrho}]$. Therefore, we get $\lambda_{1}=-(16(n-k)-15)^{-1/2}|A|$ and thus
$H(\overline{\Sigma}_{\varrho})$ $=\lambda_{1}-4(n-k-1)\lambda_{1}=\frac{4(n-k)-5}{\sqrt{16(n-k)-15}}|A|$
$>\frac{1}{2}|A|$ $\text{if }n-k\geq 2$
$\mathrm{scal}(\overline{\Sigma}_{\varrho})$ $=16(n-k-1)(n-k-2)\lambda_{1}^{2}-8(n-k-1)\lambda_{1}^{2}$
$=8(n-k-1)(2(n-k)-5)\lambda_{1}^{2}$
$=\frac{8\left((n-k)-1\right)\left(2(n-k)-5\right)}{16(n-k)-15}|A|^{2}$
$>\frac{1}{3}|A|^{2}$ $\text{if }n-k\geq 3.$
On the other hand, the cylinder of radius $\varrho^{4}$ has principal curvatures: $\lambda_{1}=0$ and $\lambda_{2}=\dots=\lambda_{n-k}=\varrho^{-4}\not=0$ and therefore
(4.2) $\begin{aligned} H(\overline{\Sigma}_{\varrho})&=\sqrt{n-k-1}\,|A|>\frac{1}{2}|A|&&\text{if }n-k\geq 2\\ \mathrm{scal}(\overline{\Sigma}_{\varrho})&=(n-k-2)|A|^{2}>\frac{1}{3}|A|^{2}&&\text{if }n-k\geq 3.\end{aligned}$
Next, we show (iii). On the cylindrical part, we have $|A|=\sqrt{n-k-1}\varrho^{-4}$, so the desired inequality holds by 4.2. On the “graphical part”, we observe that both $\xi_{\varrho}^{\prime}$ and the function $x\mapsto\frac{-x}{\sqrt{1+x^{2}}}$ are monotone decreasing and thus, for $r\leq\varrho^{2}$:
$\lambda_{2}(r)$ $=-\frac{\xi_{\varrho}^{\prime}(r)}{r\sqrt{1+\xi_{\varrho}^{\prime}(r)^{2}}}\geq-\frac{\xi_{\varrho}^{\prime}(\varrho^{2})}{\varrho^{2}\sqrt{1+\xi_{\varrho}^{\prime}(\varrho^{2})^{2}}}$
$=\frac{1}{\varrho^{2}\sqrt{1+\frac{\varrho^{2}}{\varrho-\varrho^{2}}}}\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}=\frac{\sqrt{\varrho-\varrho^{2}}}{\varrho\sqrt{\varrho}}\frac{1}{\sqrt{\varrho-\varrho^{2}}}=\frac{1}{\varrho^{3/2}}.$
Hence, we obtain
$H={}$ $-\frac{1}{4}\lambda_{2}+(n-k-1)\lambda_{2}=(n-k-\tfrac{5}{4})\lambda_{2}$
$\geq{}$ $(n-k-\tfrac{5}{4})\varrho^{-3/2}$
$\mathrm{scal}={}$ $(n-k-1)(n-k-2)\lambda_{2}^{2}-\frac{1}{2}(n-k-1)\lambda_{2}^{2}$
$\geq{}$ $(n-k-1)(n-k-\tfrac{5}{2})\varrho^{-3}$
and if $n-k\geq 3$, then $H>\varrho^{-3/2}$ and $\mathrm{scal}\geq\varrho^{-3}$, which proves (iii).
Next, we show the estimate from (iv) for $r\leq\sqrt{\varrho}$. Since $f_{\varrho}^{\prime}=\xi_{\varrho}^{\prime}$ on $\varrho^{2},\sqrt{\varrho}]$ and thus $f_{\varrho}^{\prime}$ is monotone increasing, we have
$0\geq f_{\varrho}^{\prime}(r)\geq\xi_{\varrho}^{\prime}(\varrho^{2})=-\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}\geq-2\sqrt{\varrho}.$
by Lemma A.1. Hence, $|f_{\varrho}^{\prime}(r)|\leq 2\sqrt{\varrho}$ for $r\in[\varrho^{2},\sqrt{\varrho}]$. Moreover,
$|\overline{\nu}_{\varrho}^{\intercal}|=\frac{|\nabla f_{\varrho}|}{\sqrt{1+|\nabla f_{\varrho}|^{2}}}\leq|\nabla f_{\varrho}|=|f_{\varrho}^{\prime}|\leq 2\sqrt{\varrho}.$
which proves the claim for $\varrho^{2}\leq r\leq\sqrt{\varrho}$.
Next, let us prove (iv) and (v) for $r\geq\sqrt{\varrho}$. Before we can do so, we need some estimates for the derivatives of $f_{\varrho}$ on $\sqrt{\varrho},2\sqrt{\varrho}]$. To this end, recall that
$\xi_{\varrho}^{\prime}(r)=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},\qquad 4\xi_{\varrho}^{\prime\prime}(r)=-\frac{1}{r}\xi_{\varrho}^{\prime}(r)\left(1+\xi_{\varrho}^{\prime}(r)^{2}\right).$
by Lemma A.1. Furthermore, since $\varrho$ was assumed to be smaller than $\tfrac{1}{10}$, we have
$\varrho^{1/4}-\varrho^{2}=\frac{1}{2}\varrho^{1/4}+\varrho^{1/4}\left(\frac{1}{2}-\varrho^{3/4}\right)\geq\frac{1}{2}\varrho^{1/4},$
and hence
$|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|={}$ $\frac{\varrho}{\sqrt{\varrho^{1/4}-\varrho^{2}}}\leq\sqrt{2}\,\varrho^{7/8},$
$|\xi_{\varrho}^{\prime\prime}(\sqrt{\varrho})|\leq{}$ $\frac{1}{4\sqrt{\varrho}}\,|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|\left(1+\xi_{\varrho}^{\prime}(\sqrt{\varrho})^{2}\right)\leq\frac{\sqrt{2}}{4}\,\varrho^{3/8}\left(1+2\varrho^{14/8}\right)\leq\varrho^{3/8}$
Moreover, $\xi_{\varrho}$ is strictly decreasing and thus, for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$, the fundamental theorem of calculus yields
$0\leq{}$ $\xi_{\varrho}(r)+a_{\varrho}=-\int_{r}^{2\sqrt{\varrho}}\xi_{\varrho}^{\prime}(s)ds$
$\leq{}$ $(2\sqrt{\varrho}-r)|\xi_{\varrho}^{\prime}(r)|\leq\sqrt{\varrho}\,|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|\leq\sqrt{2}\,\varrho^{11/8}.$
Therefore, since $\xi_{\varrho}^{\prime}$ is negative and strictly increasing we obtain for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$,
$|f_{\varrho}^{\prime}(r)|\leq{}$ $|\eta(r)\xi_{\varrho}^{\prime}(r)+\eta_{\varrho}^{\prime}(r)(a_{\varrho}+\xi_{\varrho}(r))|$
$\leq{}$ $-\xi_{\varrho}^{\prime}(\sqrt{\varrho})+\frac{2}{\sqrt{\varrho}}\,(a_{\varrho}+\xi_{\varrho}(\sqrt{\varrho}))$
$\leq{}$ $\sqrt{2}\,\varrho^{7/8}+\frac{2\sqrt{2}}{\sqrt{\varrho}}\,\varrho^{11/8}<3\sqrt{2}\,\varrho^{7/8}.$
By the differential equation $4\xi_{\varrho}^{\prime\prime}=-\frac{1}{r}\xi_{\varrho}^{\prime}\left(1+\xi_{\varrho}^{\prime 2}\right)$ and the monotonicity of $\xi_{\varrho}^{\prime}$, we observe that $\xi_{\varrho}^{\prime\prime}$ is positive and strictly decreasing and therefore,
$|f_{\varrho}^{\prime\prime}(r)|\leq{}$ $|\eta_{\varrho}(r)\xi_{\varrho}^{\prime\prime}(r)|+|2\eta_{\varrho}^{\prime}(r)\xi_{\varrho}^{\prime}(r)|+|\eta_{\varrho}^{\prime\prime}(r)(a_{\varrho}+\xi_{\varrho}(r))$
$\leq{}$ $\xi_{\varrho}^{\prime\prime}(\sqrt{\varrho})-\frac{4}{\sqrt{\varrho}}\,\xi_{\varrho}^{\prime}(\sqrt{\varrho})+\frac{10}{\varrho}\,(a_{\varrho}+\xi_{\varrho}(\sqrt{\varrho}))$
$\leq{}$ $\varrho^{3/8}+4\sqrt{2}\,\varrho^{3/8}+10\sqrt{2}\,\varrho^{3/8}<15\sqrt{2}\,\varrho^{3/8}$
for $r\geq\sqrt{\varrho}$.
Now, let us turn to proving (iv) and (v) for $r\geq\sqrt{\varrho}$. We observe that as before, we have $|\overline{\nu}_{\varrho}^{\intercal}|\leq|f_{\varrho}^{\prime}|$ and therefore
$|\overline{\nu}_{\varrho}^{\intercal}|\leq 3\sqrt{2}\,\varrho^{7/8}=3\sqrt{2}\,\varrho^{3/8}\sqrt{\varrho}<3\sqrt{2}\sqrt{\varrho}.$
Furthermore, by Lemma A.3, the principal curvatures of the graph of $f_{\varrho}$ are given by
$|\lambda_{1}|={}$ $\frac{|f_{\varrho}^{\prime\prime}|}{(1+f_{\varrho}^{\prime 2})^{3/2}}\leq 15\sqrt{2}\,\varrho^{3/8}$
$\lambda_{2}=\dots=\lambda_{n-k}={}$ $-\frac{f_{\varrho}^{\prime}}{r\sqrt{1+f_{\varrho}^{\prime 2}}}{\leq}\frac{3\sqrt{2}\,\varrho^{7/8}}{\sqrt{\varrho}}=3\sqrt{2}\,\varrho^{3/8}$
for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In particular, we obtain
(4.3) $|A(\overline{\Sigma}_{\varrho})|(r)=\sqrt{\lambda_{1}^{2}+(n-k-1)\lambda_{2}^{2}}\leq 3\sqrt{2(n-k+24)}\,\varrho^{3/8}.$
The scalar curvature estimate can be derived from the Gauß-equation
$\mathrm{scal}(\overline{\Sigma}_{\varrho})=H(\overline{\Sigma}_{\varrho})^{2}-|A(\overline{\Sigma}_{\varrho})|^{2}\geq-|A(\overline{\Sigma}_{\varrho})|^{2},$
which together with (4.3) yields the desired inequality and hence concludes the proof of (v). Lastly, we point out that the analysis still holds after smoothing out the $\overline{\Sigma}_{\varrho}$ at the gluing region $t=a_{\varrho}$, see for instance [3]. ∎
We close our analysis of the Euclidean handle with the following monotonicity result.
Lemma 4.4.
For every sufficiently small $R$ and $\varrho_{1}\in(0,R^{2}/4]$, there exists a family of diffeomorphisms $\Phi_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\to\overline{\Sigma}_{\varrho}$, $\varrho\in(0,\varrho_{1}]$, such that $\Phi_{\varrho_{1},\varrho_{1}}=\operatorname{\mathrm{id}}_{\overline{\Sigma}_{\varrho_{1}}}$ and for the metrics $g_{\varrho,\delta}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$, the assignment
$\varrho\mapsto e^{\tfrac{1}{2}\varrho}\cdot\Phi_{\varrho_{1},\varrho}^{\ast}(g_{\varrho,\delta})$
is strictly monotonically increasing.
We remark that the exponent $\tfrac{1}{2}$ can be replaced by an arbitrarily small positive constant.
Before we can prove Lemma 4.4, we need some preparation. Since the Euclidean handle $\overline{\Sigma}_{\varrho}\subseteq B_{1}^{n-k}(0)\times[0,1]$ is rotationally symmetric, it can be obtained by rotating a curve $\gamma_{\varrho}$. Hence, we may write the metric $g_{\overline{\Sigma}_{\varrho}}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$ in warped product form
$g_{\varrho,\delta}=\mathrm{d}\tau^{2}+r_{\varrho}(\tau)^{2}g_{S^{n-k-1}},$
where $\tau$ is the arc-length along the generating curve $\gamma_{\varrho}$ and $r_{\varrho}(\tau)$ is the radius-function. For a fixed $R\in(0,\tfrac{2}{\sqrt{10}}]$ and all $\varrho<R^{2}/4$, we denote by $d_{\varrho}$ the length of the curve $\gamma_{\varrho}$ until radius $R$. Since $\Sigma_{\varrho}$ ends in a straight cylinder, the radial function satisfies $r_{\varrho}|_{[d_{\varrho}-1+a_{\varrho},d_{\varrho}]}=\varrho^{4}$. We have the following estimate on $d_{\varrho}$ and its $\varrho$-derivative.
Lemma 4.5.
There exists a sufficiently small $R>0$ depending only on $n$, such that for every $\varrho<R^{2}/4$, we have
$|\partial_{\varrho}d_{\varrho}|<\varrho^{1/4}\qquad\text{and}\qquad 1<d_{\varrho}<1+R.$
For the following proof, we use the notation $A\lesssim B$ to mean that there exists a positive constant $C$ just depending on $n$ such that $A\leq C\cdot B$.
Proof.
Recall the function $f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr)$. The length $d_{\varrho}$ of $\gamma_{\varrho}$ is given
$d_{\varrho}=\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+\int_{4\varrho^{4}}^{R}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+(1-a_{\varrho})$
and we thus need to compute
$\partial_{\varrho}d_{\varrho}={}$ $\underbrace{\partial_{\varrho}\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr}_{\eqqcolon A_{0}}\underbrace{-16\varrho^{3}\sqrt{1+f_{\varrho}^{\prime}(4\varrho^{4})^{2}}}_{\eqqcolon A_{1}}+\underbrace{\int_{4\varrho^{4}}^{R}\frac{f_{\varrho}^{\prime}(r)\partial_{\varrho}f_{\varrho}^{\prime}(r)}{\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}}dr}_{\eqqcolon A_{2}}-\partial_{\varrho}a_{\varrho}$
We estimate $A_{0}$, $A_{1}$ and $A_{2}$ and $\partial_{\varrho}a_{\varrho}$ individually, starting with $A_{0}$. We note that an anti-derivative of $\sqrt{1+f^{\prime}_{\varrho}(r)^{2}}$ is given by
$r^{1/4}\sqrt{\sqrt{r}-\varrho^{2}}\left(\sqrt{r}+\tfrac{3}{2}\varrho^{2}\right)+\frac{3}{2}\varrho^{4}\,\operatorname{arsinh}\left(\frac{\sqrt{\sqrt{r}-\varrho^{2}}}{\varrho}\right).$
Since $\sqrt{\varrho^{4}}-\varrho^{2}=0$ and $\sqrt{\sqrt{4\varrho^{4}}-\varrho^{2}}=\varrho$, we get
$|A_{0}|=\left|\partial_{\varrho}\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr\right|\lesssim\varrho^{3}.$
Next, we turn to $A_{1}$. Since
$f_{\varrho}^{\prime}(4\varrho^{4})=\xi_{\varrho}^{\prime}(4\varrho^{4})=-\frac{\varrho}{\sqrt{\sqrt{4\varrho^{4}}-\varrho^{2}}}=-1,$
we obtain $|A_{1}|=|-16\sqrt{2}\varrho^{3}|\lesssim\varrho^{3}$.
Next, we compute
$\begin{split}\partial_{\varrho}a_{\varrho}={}&\partial_{\varrho}\left(\frac{4}{3}\varrho\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{3/2}+4\varrho^{3}\sqrt{\sqrt{2}\varrho^{1/4}-\varrho^{2}}\right)\\ ={}&\frac{4}{3}\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{3/2}+2\varrho\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{1/2}\Big(\frac{\sqrt{2}}{4}\varrho^{-3/4}-2\varrho\Big)\\ &+12\varrho^{2}\sqrt{\sqrt{2}\varrho^{1/4}-\varrho^{2}}+2\varrho^{3}\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{-1/2}\Big(\frac{\sqrt{2}}{4}\varrho^{-3/4}-2\varrho\Big)\\ \leq&11\varrho^{3/8}+4\varrho^{1+1/8-3/4}+24\varrho^{2+1/8}+2\varrho^{3-1/8-3/4}.\end{split}$
Now we turn to $A_{2}$. We have
$|A_{2}|\leq\int_{4\varrho^{4}}^{R}|f_{\varrho}^{\prime}(r)||\partial_{\varrho}f_{\varrho}^{\prime}(r)|dr.$
Note that $f_{\varrho}=\xi_{\varrho}$ for $r\leq\sqrt{\varrho}$. We obtain
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $1\cdot 2\varrho^{-1}$ $\text{for $r\in[4\varrho^{4},\varrho^{2}]$},$
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $2\sqrt{\varrho}\cdot 2\varrho^{-1/2}$ $\text{for $r\in[\varrho^{2},\varrho]$},$
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $2\varrho^{3/4}\cdot 2\varrho^{-1/4}$ $\text{for $r\in[\varrho,\sqrt{\varrho}]$}.$
Moreover, we have for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$:
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|$
$\leq$ $\Big(|\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\eta_{\varrho}||\xi_{\varrho}^{\prime}|\Big)\cdot\Big(|\partial_{\varrho}\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\partial_{\varrho}\eta_{\varrho}||\xi_{\varrho}^{\prime}|+|\eta_{\varrho}^{\prime}||\partial_{\varrho}\xi_{\varrho}+\partial_{\varrho}a_{\varrho}|+|\eta_{\varrho}||\partial_{\varrho}\xi_{\varrho}^{\prime}|\Big)$
$\lesssim$ $\Big(\varrho^{-1/2}\cdot\varrho^{5/4}+1\cdot\varrho^{3/4}\Big)\cdot\Big(\varrho^{-2}\cdot\varrho^{5/4}+\varrho^{-3/2}\cdot\varrho^{3/4}+\varrho^{-1/2}\cdot(1+1)+1\cdot\varrho^{-1/8}\Big)$
$\lesssim$ $\varrho^{3/4}\cdot\varrho^{-3/4}.$
where we used $\partial_{\varrho}a_{\varrho}\leq 1$ and that on $\sqrt{\varrho},2\sqrt{\varrho}]$ by the fundamental theorem of calculus
$|f_{\varrho}|\leq\sqrt{\varrho}|\xi^{\prime}(\sqrt{\varrho})|\leq 2\sqrt{\varrho}\cdot\varrho^{3/4}=2\varrho^{5/4}.$
Finally, for $r\geq 2\varrho$, $f_{\varrho}\equiv 0$ and therefore, putting the pieces together and integrating, we obtain
$|A_{2}|\lesssim\sqrt{\varrho}.$
Hence,
$|\partial_{\varrho}d_{\varrho}|\lesssim\sqrt{\varrho}.$
Since the curve $\gamma_{\varrho}$ is shorter than the one consisting of staight parallel to the coordinate-axes and longer than the straight line connecting its end points, we get
$1\leq\sqrt{1+R^{2}}<d_{\varrho}<1+R\qed$
Proof of Lemma 4.4.
Fix the bump function $\alpha(x)=\eta(4x)\eta(-4x)$, where $\eta$ is the cut-off function introduced earlier satisfying $2\leq\eta^{\prime}\leq 0$ and $|\eta^{\prime\prime}|<10$ and we note that $\operatorname{supp}\alpha\subseteq[-\frac{1}{4},\frac{1}{4}]$. Next, we define a smooth family or reparameterizations $\varphi_{\varrho_{1},\varrho}:[0,d_{\varrho_{1}}]\to[0,d_{\varrho}]$ for $\varrho\in(0,\varrho_{1}]$ by
$\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\alpha\left(-d_{\varrho}+\tfrac{1}{4}+s\right)ds$
where $q_{\varrho_{1},\varrho}=(d_{\varrho_{1}}-d_{\varrho})\left(\int_{\mathbb{R}}\alpha(s)ds\right)^{-1}$. We now define a diffeomorphism
$\overline{\Phi}_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\longrightarrow\overline{\Sigma}_{\varrho},\quad(\tau,x)\mapsto(\varphi_{\varrho_{1},\varrho}(\tau),x)$
where $x$ denotes the spherical coordinates in the warped product. Note that
$\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)=(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))^{2}\mathrm{d}\tau^{2}+r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))^{2}g_{S^{n-k-1}}.$
A direct inspection of the warping functions $r_{\varrho}$ gives for any $\varrho\geq\widetilde{\varrho}$
$r_{\varrho}\left(\varphi_{\varrho_{1},\varrho}(\tau)\right)\leq r_{\widetilde{\varrho}}\left(\varphi_{\varrho_{1},\widetilde{\varrho}}(\tau)\right),\qquad\tau\in[0,d_{\varrho_{1}}].$
This implies
$\partial_{\varrho}r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))=$ $(\partial_{\varrho}r_{\varrho})(\varphi_{\varrho_{1},\varrho}(\tau))+r_{\varrho}^{\prime}(\varphi_{\varrho_{1},\varrho}(\tau))\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)\geq 0$
where we used that $r_{\varrho}^{\prime}$ vanishes on the support of $\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)$. Next, we observe that
$\partial_{\tau}\varphi(\tau)=1-(d_{\varrho_{1}}-d_{\varrho})\alpha\left(-d_{\varrho}+\tfrac{1}{4}+\tau\right)$
and therefore
$\partial_{\varrho}\partial_{\tau}\varphi(\tau)=\alpha(-d_{\varrho}+\tfrac{1}{4}+\tau)\partial_{\varrho}d_{\varrho}-(d_{\varrho_{1}}-d_{\varrho})\alpha^{\prime}(-d_{\varrho}+\frac{1}{4}+\tau)\partial_{\varrho}d_{\varrho}>-\frac{1}{2}\partial_{\tau}\varphi(\tau)$
where we used that $(d_{\varrho_{1}}-d_{\varrho})\ll 1$ and that $\alpha^{\prime}$ is bounded. Consequently, we obtain for $\varrho\in[\varrho_{0},\varrho_{1}]$
$\partial_{\varrho}\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)\geq-\frac{1}{2}\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)$
which finishes the proof. ∎
5. Surgery bending in arbitrary backgrounds
In this section, we will construct the general surgery bending. To this end, let $S$ be a compact submanifold of codimension $k$ in an $n$-dimensional Riemannian manifold $(M,g)$. We denote its normal bundle by $\mathcal{N}_{S}$. Recall the constant $R\in(0,1]$ from the previous section, which we may choose smaller if necessary, so that the normal exponential map of $\mathcal{N}_{S}$ along $S$ is a diffeomorphism on the sub-bundle of vectors with length less than $R$. By abuse of notation, we call this sub-bundle $\mathcal{N}_{S}$ as well, and we denote by $D(S)$ the tubular neighbourhood of $S$ given by the image of $\mathcal{N}_{S}$ under the normal exponential map.
In order to define the submanifold $\Sigma_{\varrho}\subset M\times[0,1]$ we first choose Fermi coordinates around $S$, that is, we have an identification $U\times B^{n-k}_{R}(0)\cong D(S)|_{U}=\exp^{\perp}(\mathcal{N}_{S}|U)$, where $U\subseteq S$ is an open set over which $\mathcal{N}_{S}$ is trivial. We define for $\varrho\in(0,R^{2}/4)$
$\Sigma_{\varrho}|_{U}\coloneqq U\times\overline{\Sigma}_{\varrho}$
for $\overline{\Sigma}_{\varrho}\subset B^{n-k}_{1}(0)\times[0,1]$ as in 4.1. Note that the transition maps of our local trivialisation are given by families of orthogonal matrices acting on the normal fibers. Thus, by the O(nk)O(n-k)-invariance of $\overline{\Sigma}_{\varrho}$ (cf. Proposition 4.3, (i)), the local definitions can be glued together into a global definition of $\Sigma_{\varrho}\subset M\times[0,1]$. Moreover, since $2\sqrt{\varrho}<R$, the hypersurface $\Sigma_{\varrho}$ can be extended by $(M\setminus D(S))\times\{0\}$ to a smooth hypersurface in $M\times[0,1]$.
The main result of this section is the following theorem, which states the existence of a background metric on $D(S)\times[0,1]$ that extends $g$ on $D(S)\times\{0\}$ such that for $\varrho$ small enough, the scalar curvature of $\Sigma_{\varrho}$ can be uniformly bounded in terms of the scalar curvature of the original metric $g$ and for $\varrho$ even smaller the mean curvature of $\Sigma_{\varrho}$ is positive.
Theorem 5.1.
Let $n-k\geq 3$. Then, there exists a constant $0<\varrho_{1}<R^{2}/4$ such that for every $\varrho\in(0,\varrho_{1}]$, every $\mu\in(0,1]$ and for the metric
$g_{\mu}=u^{\frac{4}{\,n-2\,}}(g+dt^{2}),\qquad u(x,t)=u_{\mu}(x,t)=1+t\mu\eta\left(3t-1\right)$
on $D(S)\times[0,1]$ the scalar curvature of the induced metric $g_{\mu,\varrho}$ on $\Sigma_{\varrho}$ satisfies
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\varrho^{1/4}.$
Furthermore, there exists a constant $\varrho_{0}\in(0,\varrho_{1})$ such that the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\mu}$ is positive for $\varrho\in(0,\varrho_{0}]$.
Remark 5.2.
  1. (i)
    We emphasize that $\varrho_{1}$ is independent of $\mu$, while $\varrho_{0}$ will depend on $\mu$.
  2. (ii)
    The mean curvature of the slice $M\times\{0\}\subset M\times[0,1]$ with respect to the ambient metric $g_{\mu}$ is given by
    $H(M\times\{0\},g_{\mu})=\frac{2n}{n-2}\mu.$
  3. (iii)
    If nk=2n-k=2 and $n\geq 3$, then the second statement of Theorem 5.1 still holds true, that is, the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\mu}$ is still positive. However, the scalar curvature estimate might be incorrect.
Essentially, the content of Theorem 5.1 is that the handle construction of Proposition 4.3 extends to arbitrary Riemannian manifolds and to surgeries of any codimension $n-k\geq 3$. This leads to a number of error terms. Fortunately, they can be grouped into three different categories:
  1. (i)
    Error terms which approach zero as $\varrho\to 0$. These are not problematic.
  2. (ii)
    Error terms of the form $\epsilon|\overline{A}_{\varrho}|$, $\epsilon\ll 1$, where we note that the second fundamental form $|A_{\varrho}|$ blows up as $\varrho\to 0$. Using the pinching estimates $H(\overline{\Sigma}_{\varrho})>\tfrac{1}{2}|A(\overline{\Sigma}_{\varrho})|$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\tfrac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}$ we can ensure that they can be absorbed.
  3. (iii)
    Error terms which involve $|\overline{\nu}^{\intercal}_{\varrho}|$. They can be controlled by the dichotomy $\min\{H(\overline{\Sigma}_{\varrho})^{2},\mathrm{scal}(\overline{\Sigma}_{\varrho})\}>\tfrac{1}{2}\varrho^{-3}$ for $r\in[\varrho^{4},\varrho^{2}]$ and $|\overline{\nu}_{\varrho}^{\intercal}|<\overline{C}\,\sqrt{\varrho}$ for $r\geq\varrho^{2}$.
Before diving into the proof of Theorem 5.1, let us introduce its main driving force, which are estimates for scalar and mean curvature, as well as the appropriate conformal transformation formulae.
For the following lemma, we use the notation $A\lesssim B$ to mean that there exists a positive constant $C$ independent of $\varrho$ such that $A\leq C\cdot B$.
Lemma 5.3.
There exists a constant $\varepsilon>0$ such that for every $\varrho\leq\varepsilon$ and for the metric $g_{\varrho}$ on $\Sigma_{\varrho}$ induced by $g+\mathrm{d}t^{2}$, the following inequality holds:
(5.1) $\mathrm{scal}(\Sigma_{\varrho},{}$ $g_{\varrho})-\mathrm{scal}(M,g)-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\gtrsim-(\varrho+\|\overline{\nu}_{\varrho}^{\intercal}\|)\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}$
Here, $\|\overline{\nu}_{\varrho}^{\intercal}\|$ denotes the norm of the tangential part of the normal vector of the hypersurface $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{n-k+1}$ constructed in 4.1 with respect to the Euclidean metric (see also (4.1)), and $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$ denotes the scalar curvature of this hypersurface with respect to the metric $g_{\varrho,\delta}$ induced by the Euclidean ambient metric $\delta+\mathrm{d}t^{2}$. Furthermore, we have the following inequality concerning the mean curvature
(5.2) $|H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}}),$
and we can compare the $\partial_{t}$-part of the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ with respect to $g+\mathrm{d}t^{2}$ and the normal vector $\overline{\nu}_{\varrho}$ of $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{n-k+1}$ taken with respect to the Euclidean metric $\delta+\mathrm{d}t^{2}$ as follows:
(5.3) $|\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}|\lesssim\varrho.$
Lemma 5.4.
Let $u$, $g_{\mu}$ and $g_{\varrho,\mu}$ be as defined in Theorem 5.1. Then we have the following equalities comparing the scalar and mean curvature before and after conformal transformation by $u^{4/(n-2)}$:
[…]
(5.5) $\begin{split}H(\Sigma_{\varrho},g_{\mu})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)\end{split}$
where the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ and the inner product with $\partial_{t}$ are taken with respect to the metric $g+\mathrm{d}t^{2}$.
Lemma 5.4 is well-known, and we will postpone the proof of Lemma 5.3 for now and instead turn to the proof of Theorem 5.1.
Proof of Theorem 5.1.
We define $\varrho_{1}$ as follows:
$\varrho_{1}\coloneqq{}\frac{1}{2}\min\biggl\{$ $\varepsilon,\,\frac{1}{10},\,\frac{1}{36C},\,\frac{1}{12\cdot(33|\mathrm{scal}(M,g)|+360(n-1))},\,$
$\frac{1}{(9C\overline{C})^{2}},\,\frac{1}{2C\overline{C}(14n-5)},\frac{1}{(20|\mathrm{scal}(M,g)|^{2})},\,\frac{1}{8C\overline{C}+4(n-1)+2C}\biggr\}$
where $C$ and $\varepsilon$ are the constants from Lemma 5.3 and $\overline{C}=18(n-k)+432$ is the constant from Proposition 4.3.
In order to estimate the scalar curvature and mean curvature of $g_{\varrho,\mu}$, we first note that at points with radial coordinate $r\geq 2\sqrt{\varrho}$ we have $\Sigma_{\varrho}=M$ and $g_{\varrho}=g$ So, the scalar curvature estimate is trivially true here, whereas the estimate on the mean curvature follows from Lemma 5.4 using $\partial_{t}u=\mu$, $u=1=\langle\nu_{\varrho},\partial_{t}\rangle$ and $H(M\times\{0\},g+\mathrm{d}t^{2})=0$. Hence, it suffices to consider $r\leq 2\sqrt{\varrho}$. We will distinguish $3$ cases: $r\in[\varrho^{4},\varrho^{2}]$, $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In every case we will first estimate the scalar and mean curvature with respect to the ambient metric $g+\mathrm{d}t^{2}$ using Lemma 5.3 before investigating the effect of conformal change on scalar curvature employing Lemma 5.4. At the very end of the proof, we will prove positivity of the mean curvature after conformal change after passing to a possibly smaller $\varrho_{0}$, depending on $\mu$.
Let us start by considering $r\in[\varrho^{4},\varrho^{2}]$, where Proposition 4.3 yields
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq\frac{1}{2}\varrho^{-3}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2},$
which also yields $\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{6}\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|$. Furthermore, we can estimate $\|\overline{\nu}_{\varrho}^{\intercal}\|\leq 1$ and $\varrho\leq 1$, and we thus obtain from (5.1) from Lemma 5.3:
(5.6) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad={}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-C(\varrho+\|\overline{\nu}_{\varrho}^{\intercal}\|)\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}\\ &\qquad={}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}\\ &\qquad\geq{}\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C+\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-4C|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\\ &\qquad\qquad+\frac{1}{3}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\varrho|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}\varrho^{-3}\underbrace{(\frac{1}{6}-2C\varrho^{3})}_{\geq\frac{1}{6}-2C\varrho\geq\frac{1}{12}}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{18}-4C\varrho^{3/2}\right)}_{\geq(\frac{1}{6}-4C\varrho)\geq 0}\\ &\qquad\qquad+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\underbrace{\left(\frac{1}{9}-2C\varrho^{2}\right)}_{\geq\frac{1}{9}-2C\varrho\geq 0}\\ &\qquad\geq{}\frac{1}{12}\varrho^{-3}.\end{split}$
For the mean curvature, we have
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\quad\text{and}\quad H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq\frac{1}{2}\varrho^{-3/2},$
by Proposition 4.3. Therefore, (5.2) becomes:
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-2C-C\varrho\cdot|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}}$
$\geq{}$ $\underbrace{\frac{1}{4}\varrho^{-3/2}-2C}_{\geq\frac{1}{8}\varrho^{-3/2}}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{4}-C\varrho\right)}_{\geq 0}\geq\frac{1}{8}\varrho^{-3/2}.$
To close the first case of the proof for the scalar curvature estimate, let us consider the conformal transformation formula (5.4) from Lemma 5.4:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})={}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\frac{4(n-1)}{n-2}(1-\langle\nu_{\varrho},\partial_{t}\rangle^{2})\partial^{2}_{t}u$
$\qquad\qquad+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\partial_{t}u\Bigr)$
If $t\geq 2\varrho$, then $\Sigma_{\varrho}$ is given by a cylinder, so $\langle\nu_{\varrho},\partial_{t}\rangle=0$. Furthermore, $u\leq 2$ and $\partial^{2}_{t}u=3\mu\eta^{\prime}(3t-1)+9t\mu\eta^{\prime\prime}(3t-1)\leq 90$ by the choice of $\eta$. Therefore, the above formula gives the following estimate to
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $2^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)+\underbrace{\frac{1}{12}\varrho^{-3}-\frac{360(n-1)}{n-2}}_{\geq 33|\mathrm{scal}(M,g)|\geq 0}\Bigr)\geq|\mathrm{scal}(M,g)|.$
On the other hand, if $t\leq 2\varrho$, then $u=1+t\mu$, $\partial_{t}u=\mu$ and $\partial_{t}^{2}u=0$. As $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ is positive, so is the entire final summand, and we get
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $2^{-\frac{n+2}{n-2}}\underbrace{\Bigl(\mathrm{scal}(M,g)+\frac{1}{12}\varrho^{-3}\Bigr)}_{\geq 32|\mathrm{scal}(M,g)|\geq 0}\geq|\mathrm{scal}(M,g)|,$
which finishes the first case.
Next, we investigate the region $r\in[\varrho^{2},\sqrt{\varrho}]$, where Proposition 4.3 gives the following estimates:
$\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\quad\text{and}\quad\|\nu_{\varrho}^{\intercal}\|\leq\overline{C}\sqrt{\varrho}$
for $\overline{C}=18(n-k)+432$ the constant from Proposition 4.3. Furthermore, we have $\varrho\leq\overline{C}\sqrt{\varrho}$ and therefore, (5.1) from Lemma 5.3 yields
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)$
$\quad\geq\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\overline{C}\sqrt{\varrho}\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}$
$\quad\geq\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\overline{C}\sqrt{\varrho}-4C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-2C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
If $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq 1$, then this becomes $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})\geq\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}$ since $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq 0$. On the other hand, if $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\geq 1$, then $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$ and we obtain
(5.7) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}-2C\overline{C}\sqrt{\varrho}-6C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}-2C\overline{C}\sqrt{\varrho}+\underbrace{\left(\frac{1}{3}-6C\overline{C}\sqrt{\varrho}\right)}_{\geq 0}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}-2C\overline{C}\sqrt{\varrho},\end{split}$
so $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})\geq\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}$. Concerning the mean curvature, Proposition 4.3 yields
$H(\overline{\Sigma}_{\varrho},\delta)\geq\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta)|,$
and we can hence estimate
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-C\overline{C}\sqrt{\varrho}\cdot(2+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}})$
$\geq{}$ $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{2}-C\sqrt{\varrho}\right)}_{\geq 0}-2C\overline{C}\sqrt{\varrho}\geq-2C\overline{C}\sqrt{\varrho}$
By employing the conformal change formula from Lemma 5.4, we obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})={}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\mu\Bigr)$
$\geq$ ${}u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(8C\overline{C}\varrho^{1/4}+C\overline{C}\frac{8(n-1)}{n-2}\varrho^{1/4}\right)}_{\leq\tfrac{1}{2}}\Bigr)$
$\geq$ ${}u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\right).$
Since $u=1+t\mu\leq 1+2\varrho\mu$, we have $1\geq u^{-1}\geq u^{-(n+2)/(n-2)}\geq u^{{-5}}\geq 1-5t\mu\geq 1-10\varrho$. If $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\leq 0$, then $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}$, so it suffices to consider the case that $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\geq 0$, where we can argue similarly:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq$ ${}(1-10\varrho)\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4})\right)$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\Bigl(10|\mathrm{scal}(M,g)|\varrho^{3/4}+\frac{1}{2}\Bigr)}_{\leq 1}\geq\mathrm{scal}(M,g)-\varrho^{1/4}.$
Finally, we consider $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. By Proposition 4.3, we have
$\|\nu_{\varrho}^{\intercal}\|\leq\overline{C}\sqrt{\varrho},\qquad|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq\overline{C}\varrho^{3/8}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho})\geq-\overline{C}\varrho^{3/4}.$
By consulting (5.1) from Lemma 5.3 one final time, we deduce that
(5.8) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad\geq-\overline{C}\varrho^{3/4}-C\varrho-C\cdot\Bigl(3\overline{C}\varrho^{1/2}+2\overline{C}\varrho^{7/8}+\overline{C}\varrho^{14/8}\Bigr)\\ &\qquad\geq{}\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}.\end{split}$
Furthermore, we have
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $\underbrace{H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})}_{\geq-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|}-C\left(\|\nu_{\varrho}^{\intercal}\|+\varrho^{2}(1+\|\nu_{\varrho}^{\intercal}\|+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$\geq{}$ $-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-2C\sqrt{\varrho}$
$\geq{}$ $-n\varrho^{3/8}-2C\sqrt{\varrho}.$
An application of Lemma 5.4 together with the observation that $\partial_{t}u=\mu$ and $\partial_{t}^{2}u=0$ now gives:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(8C\overline{C}\varrho^{1/4}+\frac{4(n-1)}{n-2}n\varrho^{1/8}+2C\varrho^{1/4}\right)}_{\leq\frac{1}{2}}\Bigr)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\right).$
If $\mathrm{scal}(M,g)-\tfrac{1}{2}\varrho^{1/4}<0$ is negative, then $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\tfrac{1}{2}\varrho^{1/4}$, because $u^{-(n+2)/(n-2)}\leq 1$. On the other hand, if $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\geq 0$, we again use $u^{-(n+2)/(n-2)}\geq 1-10\varrho$ to obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $(1-10\varrho)(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4})$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(10|\mathrm{scal}(M,g)|\varrho^{3/4}+\frac{1}{2}\right)}_{\leq 1}$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}$
which finishes the proof of the first half of Theorem 5.1 concerning the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\mu})$.
It remains to show positivity of the mean curvature with respect to the conformally transformed metric after passing to a smaller radius bound $\varrho_{0}$. To be precise, let
$\varrho_{0}\coloneqq{}$ $\frac{1}{2}\min\left\{\frac{n-2}{104n-16},\,\frac{1}{(3C)^{4}},\,\frac{1}{(n+2C)^{8}},\,\frac{1}{5^{8}},\mu^{8}\right\}$
Let us recall the conformal transformation formula:
$H(\Sigma_{\varrho},g_{\mu})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)$
We will again distinguish radial cases.
If $r\in[\varrho^{4},\varrho^{2}]$, we have already seen that
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq\frac{1}{8}\varrho^{-3/2}$
and, since $\partial_{t}u=\mu\eta(3t-1)+3t\mu\eta^{\prime}(3t-1)\geq-6$, $u^{-1}\leq 1$, $\mu\leq 1$ and $\langle\nu_{\varrho},\partial_{t}\rangle\leq 1$, we thus obtain
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $u^{-\frac{2}{n-2}}\underbrace{\left(\frac{1}{8}\varrho^{-3/2}-\frac{12n}{n-2}\right)}_{\geq 1}>0,$
which is positive and independent of $\mu$.
For the final case $r\in[\varrho^{2},2\sqrt{\varrho}]$, we first observe, that we either have
$H(g+\mathrm{d}t^{2})\geq-3C\sqrt{\varrho}=-\varrho^{1/4}\underbrace{(3C\varrho^{1/4})}_{\leq 1}\geq-\varrho^{1/4}$
if $r\leq\sqrt{\varrho}$ or
$H(g+\mathrm{d}t^{2})\geq-n\varrho^{3/8}-2C\sqrt{\varrho}=-\varrho^{1/4}\underbrace{(n\varrho^{1/8}-2C\varrho^{1/4})}_{\leq 1}\geq-\varrho^{1/4}$
if $r\geq\sqrt{\varrho}$. Furthermore, we also have
$\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\geq\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}-2C\varrho^{2}\geq 1-\sqrt{\varrho}-2C\varrho^{2}\geq\frac{1}{2}.$
We can now plug both of these, together with $\partial_{t}u=\mu$, $1\geq u^{-1}\geq 1-2\varrho\mu$ and $u^{-\frac{2}{n-2}}\geq 1-4\varrho\mu$ into the transformation formula to obtain
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $u^{-\frac{2}{n-2}}\Bigl(-\varrho^{1/4}+\underbrace{\frac{2n}{n-2}\frac{1}{2}}_{\geq 1}(1-2\varrho\mu)\mu\Bigr)$
$\geq{}$ $(1-4\varrho\mu)\left(\mu-\varrho^{1/4}-\varrho\mu^{2}\right)$
$\geq{}$ $\mu-\varrho^{1/8}(4\mu^{2}\varrho^{7/8}+\varrho^{1/8}).$
Since $\mu^{2}\varrho^{6/8}\leq 1$, we have $(4\mu^{2}\varrho^{7/8}+\varrho^{1/8})\leq 5\varrho^{(1/8)}\leq 1$ and thus
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $\mu-\varrho^{1/8}$
which is positive by assumption $\varrho_{0}\leq\frac{1}{2}(\mu)^{8}$ and thus finishes the proof of Theorem 5.1. ∎
Let us now delve into the proof of Lemma 5.3.
Proof of Lemma 5.3.
By the Gauß-Equation we have
(5.9) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\underbrace{\mathrm{scal}(M\times[0,1],g+\mathrm{d}t^{2})}_{=\mathrm{scal}(M,g)}\\ &\qquad={}H^{2}(\Sigma_{\varrho},g+\mathrm{d}t^{2})-|A(\Sigma_{\varrho},g+\mathrm{d}t^{2})|^{2}-2\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho}),\end{split}$
where $\nu_{\varrho}$ is the upwards unit normal vector of $\Sigma_{\varrho}$ with respect to the ambient metric $g+\mathrm{d}t^{2}$. In order to estimate the summands from the right-hand side of (5.9), consider the symmetric $2$-tensor $\gamma$ on $S$ defined as the first order Taylor expansion of $g$ in normal direction, that is
(5.10) $\gamma_{\alpha\beta}=g_{\alpha\beta}|_{S}+\sum_{i=1}^{n-k}x^{i}g_{\alpha\beta,i}|_{S}$
which we extend canonically to $B_{R}^{n-k}$. Since the tensor $\gamma$ precisely describes the $1$-jet of the metric $g$ along $S$, there exists an $\varepsilon>0$ such that the restriction of $\gamma$ to the disk-bundle of radius $\varepsilon$ is positive definite and hence a Riemannian metric. If $r$ denotes the radial component of the fiber of the tubular neighbourhood $D(S)$ in $M$, we have
$|g-\gamma|_{\gamma}\leq C\cdot r^{2},$
since $g_{\alpha\beta}(x)=g_{\alpha\beta}(p)+x^{k}g_{\alpha\beta,k}(p)+O(r^{2})$ by the Taylor remainder estimate. The advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 4.
The proof now consists of investigating the difference in second fundamental form, mean curvature and normal vector when passing from the background metric $g+\mathrm{d}t^{2}$ to $\gamma+\mathrm{d}t^{2}$ and afterwards relating those terms for $\gamma+\mathrm{d}t^{2}$ to the corresponding ones for the Euclidean background metric $\delta+\mathrm{d}t^{2}$ in the fibers. Let us abbreviate
$A_{g}\coloneqq{}$ $A(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ $A_{\gamma}\coloneqq A(\Sigma_{\varrho},\gamma+\mathrm{d}t^{2})$ $A_{\delta}\coloneqq A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})$
$H_{g}\coloneqq{}$ $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ $H_{\gamma}\coloneqq H(\Sigma_{\varrho},\gamma+\mathrm{d}t^{2})$ $H_{\delta}\coloneqq H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})$
Furthermore, we denote by $\nu_{g}$, $\nu_{\gamma}$ and $\nu_{\delta}$ the respective upward unit normal vector fields of $\Sigma_{\varrho}$ or $\overline{\Sigma}_{\varrho}$. We have the following comparison formulae
$|\nu_{g}-\nu_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $r^{2},\qquad$ $|A_{g}-A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot r^{2},$
where the latter identity follows from $A_{g}(e_{i},e_{j})=-\langle\nabla^{g+dt^{2}}_{e_{i}}e_{j},\nu_{g}\rangle_{g}$ and $|\Gamma^{k}_{ij}|_{g}\lesssim r^{2}$ for the Christoffel symbols $\Gamma_{ij}^{k}=\tfrac{1}{2}g^{kl}(\nabla^{\gamma}_{i}g_{lj}+\nabla^{\gamma}_{j}g_{li}+\nabla^{\gamma}_{l}g_{ij})$.
Since the difference of $\Sigma_{\varrho}$ and $M\times\{0\}$ is supported on $\{r\leq 2\sqrt{\varrho}\}$, we thus obtain
$|H_{g}-H_{\gamma}|={}|\operatorname{tr}(A_{g})-\operatorname{tr}(A_{\gamma})|\lesssim{}\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|\lesssim{}\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
and therefore
$\bigl||A_{g}|_{g+\mathrm{d}t^{2}}^{2}$ $-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}\bigr|$
$={}$ $\bigl|(|A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})(|A_{g}|_{g+\mathrm{d}t^{2}}+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\bigr|$
$={}$ $\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\leq{}$ $\varrho^{2}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\leq{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\left(\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\right)$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
$|H_{g}^{2}-H_{\gamma}^{2}|={}$ $\bigl|(H_{g}-H_{\gamma})(H_{g}+H_{\gamma})\bigr|\leq|H_{g}-H_{\gamma}|^{2}+2|H_{\gamma}|\cdot|H_{g}-H_{\gamma}|$
$\leq{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}\cdot\varrho^{2}+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$={}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho((1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\lesssim{}$ $\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}.$
Putting these together, we obtain
$\bigl|(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}\bigr|$ $\lesssim\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
[…]
If we write $\nu_{g}^{\intercal}\coloneqq\nu_{g}-\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\partial_{t}$ and $\nu_{\gamma}^{\intercal}\coloneqq\nu_{\gamma}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\partial_{t}$, then we observe
$\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\bigr|}_{\leq\|\nu_{g}-\nu_{\gamma}\|_{g+\mathrm{d}t^{2}}\lesssim\varrho}+\underbrace{\bigl|\langle\nu_{\gamma},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\leq\|g-\gamma\|_{\gamma}\lesssim\varrho}\lesssim{}\varrho$
$\left|\|\nu_{g}^{\intercal}\|_{g}-\|\nu_{\gamma}^{\intercal}\|_{\gamma}\right|$
$\qquad\leq{}\bigl|\|\nu_{g}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}\bigr|+\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\|\nu_{g}^{\intercal}-\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}+\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\|\nu_{g}-\nu_{\gamma}\|_{g+\mathrm{d}t^{2}}+\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\varrho}+\underbrace{\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\leq\|g-\gamma\|_{\gamma+\mathrm{d}t^{2}}\lesssim{}\varrho}$
$\qquad\lesssim{}\varrho$
Regarding the final summand in (5.9), we note that $\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\partial_{t},\partial_{t})=0$ and therefore, we obtain
$\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{g},\nu_{g})={}$ $\|\nu_{g}^{\intercal}\|^{2}\cdot\mathrm{Ric}(M,g)\left(\frac{\nu_{g}^{\intercal}}{\|\nu_{\gamma}^{\intercal}\|},\frac{\nu_{\gamma}^{\intercal}}{\|\nu_{\gamma}^{\intercal}\|}\right)$
$\lesssim{}$ $\|\nu_{g}^{\intercal}\|^{2}\lesssim\|\nu_{\gamma}^{\intercal}\|$
We note that the constant in the above inequality depends on the maximum of the Ricci-curvature of $g$ and hence, in particular on $g$, but is independent of $\varrho$.
As mentioned above, the advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 4. To make this precise, we first observe that the projection
$\pi:D(S)\times[0,1]\longrightarrow B_{R}^{n-k}(0)\times[0,1]\subseteq\mathbb{R}^{n-k}\times[0,1]$
satisfies $D\pi(\nu_{\gamma})=\nu_{\delta}$. Furthermore, the restriction of $\pi$ to the fibers is a radial isometry, so
$\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}=\|\nu_{\delta}^{\intercal}\|_{\delta+\mathrm{d}t^{2}}\quad\text{and}\quad\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}=\langle\nu_{\delta},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}$
In particular, we immediately obtain $|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}|\lesssim{}\varrho$, proving the final estimate from Lemma 5.3.
If we fix a point $x\in\Sigma_{\varrho}$ and choose an adapted $\gamma$-orthonormal frame $\{e_{i},e_{a}\}$ of $\Sigma_{\varrho}$ at $x$, where $\{e_{i}\}$ are normal to $S$ and $\{e_{a}\}$ tangent to $S$, we see that $\{D\pi(e_{i})\}$ form normal coordinates at $\pi(x)\in\overline{\Sigma}_{\varrho}$. By definition of the second fundamental form,
$A_{\gamma}(e_{i},e_{j})={}$ $-\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla^{\delta+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\delta}\rangle=A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})(e_{i},e_{j}).$
Next, we note that $\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{a}}e_{j},e_{i}\rangle$ and $\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{a}}e_{b},e_{i}\rangle$ are components of the second fundamental form of the fibers of the tubular neighbourhood, hence they are independent of $\varrho$, and we get $\langle\nabla_{e_{a}}e_{j},e_{i}\rangle,\langle\nabla_{e_{a}}e_{b},e_{i}\rangle\lesssim 1$. Moreover, both $\nabla_{e_{a}}e_{j}$ and $\nabla_{e_{a}}e_{b}$ are orthogonal to $\partial_{t}$, so we obtain
$A_{\gamma}(e_{a},e_{j})={}$ $-\langle\nabla_{e_{a}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla_{e_{a}}e_{j},\nu_{\gamma}^{\intercal}\rangle$
$={}$ $-\sum_{i}\langle\nabla_{e_{a}}e_{j},e_{i}\rangle\langle\nu^{\intercal}_{\gamma},e_{i}\rangle\lesssim\|\nu^{\intercal}_{\gamma}\|=\|\nu^{\intercal}_{\delta}\|$
and $A_{\gamma}(e_{a},e_{b})\lesssim\|\nu_{\delta}^{\intercal}\|$, analogously. Thus, we have obtained
$|A_{\gamma}(e_{i},e_{a})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad|A_{\gamma}(e_{a},e_{b})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad\text{and}\quad A_{\gamma}(e_{i},e_{j})=A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})(e_{i},e_{j}).$
Using these, we immediately get
$|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|^{2}$
$|H_{\gamma}-H_{\delta}|\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|$
which, together with $|H_{g}-H_{\gamma}|\lesssim{}\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$ combines to
$|H_{g}-H_{\delta}|\lesssim\|\nu_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}),$
proving the desired mean curvature estimate.
It remains to verify the scalar curvature estimate. We have
$|H_{\gamma}^{2}-H_{\delta}^{2}|={}$ $2|H_{\delta}|\cdot|H_{\gamma}-H_{\delta}|+(H_{\gamma}-H_{\delta})^{2}$
$\lesssim{}$ $|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\cdot\|\nu_{\delta}^{\intercal}\|+\|\nu_{\delta}^{\intercal}\|^{2}$
$\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
where we used $|H_{\delta}|\leq n|A_{\delta}|$}. Combining these, we get
$(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})-(H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2})\gtrsim{}$ $-\|\nu_{\delta}^{\intercal}\|(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}).$
Furthermore, there exists a constant $C$ such that
$|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\leq{}$ $\sqrt{|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}+C\|\nu_{\delta}^{\intercal}\|^{2}}$
$\leq{}$ $\sqrt{|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}+2\sqrt{C}\|\nu_{\delta}^{\intercal}\|\cdot|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+C\|\nu_{\delta}^{\intercal}\|^{2}}$
$\leq{}$ $|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+C\|\nu_{\delta}^{\intercal}\|$
and therefore $|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\lesssim\|\nu_{\varrho}^{\intercal}\|\leq 1$. Note that $H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$ is precisely the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\delta})$, where $g_{\varrho,\delta}$ is induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$. Thus, we obtain
$(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\qquad=(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}+H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+\|\nu_{\delta}^{\intercal}\|)^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}$
Putting all the pieces together, we therefore obtain the following inequality, where the constant $C$ might change from line to line, independently of $\varrho$:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\qquad=\underbrace{H^{2}_{g}-|A_{g}|^{2}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})}_{\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}}-2\underbrace{\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho})}_{\lesssim\|\nu_{\delta}^{\intercal}\|}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+\|\nu_{\delta}^{\intercal}\|)^{2}-\|\nu_{\delta}^{\intercal}\|$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}$
which finishes the proof. ∎
Next, we compare $\Sigma_{\varrho}$ for different values of $\varrho$. Lemma 4.4 immediately yields:
Proposition 5.5.
For every $\mu$ and every $\varrho\in[\varrho_{0},\varrho_{1}]$ there exists diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$ such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$ and for $g_{\mu,\varrho}$ the metric on $\Sigma_{\varrho}$ induced by $g_{\mu}$ (see Theorem 5.1) the assignment
$\varrho\mapsto e^{\varrho}\cdot\Phi_{\varrho}^{\ast}(g_{\varrho,\mu})$
is strictly monotonically increasing.
Finally, we deform the metric $e^{\varrho_{1}}\cdot g_{\mu,\varrho_{1}}$ to a metric which is independent of $\mu$ by linear interpolation of conformal factors:
Proposition 5.6.
For $\varrho_{1}>0$ and every $\mu\in(0,1]$, define the path $g_{s}$, $s\in[0,1]$ of metrics on $\Sigma_{\varrho_{1}}$ to be induced by the family
$s\mapsto e^{\varrho_{1}}((1-s)\cdot u+2s)^{\frac{4}{n-2}}\cdot(g+\mathrm{d}t^{2})$
of metrics on $M\times[0,1]$, where $u=(1+t\mu\eta(3t-1))$ as in Theorem 5.1. Then, $g_{s}$ is strictly monotonically increasing in $s$ and satisfies
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq-e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4}).$
Proof.
Since $u_{s}=(1-s)\cdot u+2s$ is the linear path and $u\leq 2$, the monotonicity of $g_{s}$ follows immediately. Concerning the scalar curvature estimate, we recall the conformal transformation formula for scalar curvature we have
$\mathrm{scal}(\Sigma_{\varrho_{1}},u_{s}^{\frac{4}{n-2}}\cdot e^{\varrho_{1}}g_{\varrho_{1}})={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u_{s}-\frac{4(n-1)}{n-2}(\Delta_{g}u_{s})\right)$
$={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\Bigl((1-s)\underbrace{\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u-\frac{4(n-1)}{n-2}(\Delta_{g}u)\right)}_{=u^{(n+2)/(n-2)}\mathrm{scal}\left(\Sigma_{\varrho_{1}},g_{\varrho_{1},\mu}\right)}$
$\qquad\qquad\qquad\qquad\qquad+2s\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\Bigr).$
Note that
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\mu})\geq\mathrm{scal}(M,g)-\varrho_{1}^{1/4}\geq-|\mathrm{scal}(M,g)|-\varrho_{1}^{1/4}.$
Furthermore, the proof of Theorem 5.1 reveals that $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\geq-|\mathrm{scal}(M,g)|-\varrho_{1}^{1/4}$ as well, see (5.6), (5.7) and (5.8). Combining this with $u_{s}\geq 1$ which gives $u_{s}^{-(n+2)/(n-2)}\geq-1$, we obtain
$\mathrm{scal}(\Sigma_{\varrho_{1}},u_{s}^{\frac{4}{n-2}}\cdot e^{\varrho_{1}}g_{\varrho_{1}})\ \geq{}$ $-u_{s}^{-\frac{n+2}{n-2}}e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4})$
$\geq{}$ $-e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4})$
which is bounded independently of $\mu$. ∎
6. Proof of Theorem 1.2
Let $\sigma\in\mathbb{R}$, $(M,g_{M})$ and $(\Omega,g_{\Omega})$ be as in Theorem 1.2. Using the arguments from the beginning of Section 2, we can assume without loss of generality that $\Omega$ is orientable and admits a spin structure for $\dim M\geq 7$. Furthermore, we can assume that $S^{n}$ can be obtained from $M$ by a sequence of surgeries of codimension at least $3$. Let $\ell$ be the minimal number of surgeries required. The proof is by induction on $\ell$, where the case $\ell=0$ has been treated in [25].
By a small deformation of $g_{\Omega}$ near $\partial\Omega\subset\Omega$, leaving the induced metric on $\partial\Omega$ unchanged and decreasing the total mean curvature and scalar curvature by an arbitrarily small amount, we can assume (replacing $\sigma$ with $\sigma-1$) that $H_{g_{\Omega}}>0$ along $\partial\Omega$. In this situation the proof of Section 2 no longer applies, since the width of the attached handles becomes smaller when $H_{g_{\Omega}}$ approaches 0. Instead, we use the following argument in the surgery step.
Let $\widetilde{M}$ be obtained from $M$ by a surgery in codimension $k\geq 3$ such that $S^{n}$ can be obtained from $\widetilde{M}$ by $\ell$ surgeries of codimension at least $3$. Put d:=nkd:=n-k.
Let $\psi\colon S^{d}\times D^{k}\hookrightarrow M$ be the surgery datum for $\widetilde{M}$. Then, there exists a bordism $T\colon M\leadsto\widetilde{M}$ consisting of a single $\ell$-handle $D^{d+1}\times S^{k}$, which is glued onto $M$ along $\psi$.
Let $\mu>0$ and consider the metric $g_{\mu}$ from Theorem 5.1 on $M\times[0,1]$. This extended metric can then be further extended to a metric $\widetilde{g}_{\mu}$ on the handle $M\times[0,1]\cup_{M\times\{1\}}T$ in an arbitrary fashion. By Theorem 5.1 and Proposition 5.6, there exist constants $0<\varrho_{0}<\varrho_{1}<\frac{1}{10}$, a family of submanifolds $\Sigma_{\varrho}\subset\widetilde{\Omega}$ for $\varrho\in(0,\varrho_{1}]$ and a family of diffeomorphisms $\Phi_{\varrho}\colon\Sigma_{\varrho_{1}}\to\Sigma_{\varrho_{0}}$ such that
  1. (i)
    $\Sigma_{\varrho}$ is diffeomorphic to $\widetilde{M}$ and divides $M\times[0,1]\cup T$ into two manifolds $\widetilde{T}$ and $\widetilde{U}$, where $\widetilde{T}$ is diffeomorphic to $M\times[0,1]\cup T$.
  2. (ii)
    $\mathrm{scal}(\Sigma_{\varrho})>\mathrm{scal}(M)-\varrho^{1/4}$ for all $\varrho$.
  3. (iii)
    $H_{g_{\mu}}|_{\Sigma_{\varrho_{0}}}>0$.
  4. (iv)
    $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$.
  5. (v)
    $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}(g_{\mu}|_{\Sigma_{\varrho}})$ is strictly increasing.
Note that apriori the surfaces $\Sigma_{\varrho}$ constructed in Section 5 have a boundary, however, they can be capped off in a standard fashion while preserving all geometric bounds.
[…]
Figure 9. Schematic depiction of the quantitative surgery construction.
We also note that $\varrho_{1}$ is independent of $\mu$. By Proposition 5.6, there exists an increasing path $g_{t}$ from $e^{\varrho_{1}}\widetilde{g}|_{\Sigma_{\varrho_{1}}}$ to $2^{4/(n-2)}{g^{\prime}}|_{\Sigma_{\varrho_{1}}}$, where $g^{\prime}$ is some fixed extension of the metric $g\cup g|_{M}+\mathrm{d}t^{2}$. Furthermore, the metric $g^{\prime}$ is independent of $\mu$ and the scalar curvature of every metric of this path can be uniformly bounded independent of $\mu$.
Let us abbreviate $\Sigma\coloneqq\Sigma_{\varrho_{1}}$. We will now apply the criterion provided by Proposition 3.2. Let
$g_{\Sigma}$ $\coloneqq 2^{4/(n-2)}{g^{\prime}}|_{\Sigma}$
$s$ $\coloneqq\min\{\min(\mathrm{scal}(g_{t})),\mathrm{scal}(g)-\varrho_{1}^{1/4}\}.$
Then, by the above, $\Phi_{\varrho}^{\ast}(\widetilde{g}_{\mu}|_{\Sigma_{\varrho}})\in\mathscr{R}_{g_{\Sigma},s}(\Sigma_{\varrho_{1}})$ for every $\varrho\leq\varrho_{1}$ (see Section 3 for the definition of $\mathscr{R}_{g_{\Sigma},s}$).
First, we note that by induction hypothesis, the theorem holds for the metric $g_{\Sigma}$ on $\Sigma$. If $(\Omega,g_{\Omega})$ is a fill-in as in Theorem 1.2, then $\Omega\cup\widetilde{T}$ is a fill-in of $\Sigma$. Since we assumed $H_{g_{\Omega}}>0$ and $M$ is compact, there exists a positive lower bound $0<\mu<H_{g_{\Omega}}$. Moreover, by performing a smoothing procedure similar to the one in the proof of Theorem 2.1 we can extend the metric $\widetilde{g}_{\mu}$ to this fill-in. Since the Riemannian manifolds $M$ and $\Sigma_{\varrho_{0}}$ only differ on a small neighbourhood, we may assume that
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq 2\int_{\Sigma_{\varrho_{0}}}H_{\widetilde{g}_{\mu}}{\rm dvol}_{\widetilde{g}_{\mu}}|_{\Sigma_{\varrho_{0}}}.$
The extended metric $\widetilde{g}$ on $\widetilde{\Omega}\coloneqq\Omega\cup T$ constitutes a fill-in of $\Sigma_{\varrho_{0}}$ with positive mean curvature. Since $(\Sigma_{\varrho_{0}},\widetilde{g}|_{\Sigma_{\varrho_{0}}})$ is isometric to $(\Sigma_{\varrho_{1}},\Phi_{\varrho}^{\ast}(\widetilde{g}|_{\Sigma_{\varrho}})$ and $\Phi_{\varrho}^{\ast}(\widetilde{g}|_{\Sigma_{\varrho}})\in\mathscr{R}_{g_{\Sigma},s}$, Proposition 3.2 implies that there exists a constant $\Lambda=\Lambda(\Sigma,g_{\Sigma})$ such that
$\int_{\Sigma_{\varrho_{0}}}H_{\widetilde{g}_{\mu}}{\rm dvol}_{\widetilde{g}_{\mu}}|_{\Sigma_{\varrho_{0}}}\leq\Lambda.$
Since $\Lambda$ is independent of $\mu$ and hence independent of the fill-in, this finishes the proof.
7. A new quasi-local mass
Theorem 1.2 has important applications in Physics. In 1982, Penrose [23] listed important open problems in General Relativity. Finding a suitable notion of quasi-local mass was listed first.
In Newtonian gravity, the mass density function $\rho$ satisfies Poisson’s equation
$\Delta V=4\pi\rho,$
where $V$ denotes the gravitational potential. The total mass contained in a region $\Omega$ is then given by
$m(\Omega)=\int_{\Omega}\rho\,d\mu=\frac{1}{4\pi}\int_{\partial\Omega}\nabla_{\nu}V\,d\mu,$
where $\nabla_{\nu}$ denotes the outward normal derivative. However, in general relativity, the geometric nature of spacetime makes it impossible to express the mass as an integral of a local density depending only on the metric and its derivatives. To overcome this difficulty, two main approaches have been developed:
The first is to define the total mass for an isolated system, namely the ADM mass, associated with an asymptotically flat manifold $(N,g_{N})$. Here, asymptotically flat initial data sets $(N,g_{N},k)$ model isolated gravitational systems, such as stars or galaxies. The mass $m(N,g_{N},k)$ is well-defined and enjoys many desirable properties such as positivity and rigidity. However, it has the drawback that it applies only to entire non-compact manifolds, rather than to bounded regions within them. For instance, if $(N,g_{N},k)$ models two black holes contained in its interior, the ADM mass only measures the total mass of the system, without distinguishing the contributions of the individual black holes.
In order to define the mass of a finite region–known as quasi-local mass–various proposals have been made, including the Hawking mass [19], the Bartnik mass [6], the Brown–York mass [13], and the Wang–Yau mass [28]. Each of these notions has attractive geometric features but also important limitations. A striking common feature is that all of them are particularly well-suited for spherical topology: The Hawking mass often underestimates the true mass and is most effectively used together with inverse mean curvature flow, whose monotonicity holds only for $M\cong S^{2}$. The Brown–York and Wang–Yau masses rely on isometric embeddings into Euclidean or Minkowski space, and classical embedding results such as the Weyl theorem require positive Gauss curvature. The Bartnik mass, perhaps the most mathematically natural proposal, has so far been explicitly computed only for spherical boundaries.
This contrasts with the increasing importance of spacetimes with more complicated topology including Emparan–Reall’s black ring [16], the Kottler metric [21], and the Horowitz–Myers soliton [11].
Motivated by [22] we define a new notion of quasi-local mass which has no such topological shortcomings. For this purpose, let $(N,g_{N})$ be a complete manifold with $\mathrm{scal}_{g_{N}}\geq\sigma$, $\sigma\in\mathbb{R}$, and let $\Omega\subseteq N$ with $M=\partial\Omega$. We set
$\textbf{m}_{\sigma}(M)=\Lambda(M,g_{M},\sigma)-\int_{M}H_{g}{\rm dvol}_{g_{M}}.$
This should be compared with Bray’s inner mass [10] which also relies on fill-ins. Theorem 1.2 immediately yields
Corollary 7.1.
Let $M$ be spin. Then the quasi-local mass $\textbf{m}_{\sigma}(M)$ is well-defined. Moreover, $\textbf{m}_{\sigma}(M)\geq 0$.
Here $\tfrac{1}{2}\sigma$ corresponds to the cosmological constant. We point out that the ADM mass for asymptotically flat manifolds has been successfully extended to the setting $\sigma<0$ [29], though there is no analogue for $\sigma>0$ [12]. On the other hand $\textbf{m}_{\sigma}$ is well-defined and non-negative for any choice of $\sigma\in\mathbb{R}$.
When $n=3$ and $M\cong S^{2}$, this recovers the notion introduced by Mantoulidis–Miao [22, Theorem 1.5]. As in their setting, we obtain:
Corollary 7.2.
Suppose that $n=3$, $M=S^{2}$ and $K(M)>0$ for the Gaussian curvature. Then
$\textbf{m}_{0}(M)=m_{BY}(M)$
for the Brown–York mass. In particular, we have in this case $\textbf{m}_{0}(M)=0$ if and only if $\Omega\subseteq\mathbb{R}^{3}$.
Appendix A Collection of auxiliary results
This collection of elementary results is used to construct handles in Euclidean space around points.
Lemma A.1.
Let $0<\varrho<1$ and consider on the interval $\varrho^{4},1]$ the function
$\xi_{\varrho}(r)=-\frac{4}{3}\varrho\left(\sqrt{r}-\varrho^{2}\right)^{3/2}-4\varrho^{3}\sqrt{\sqrt{r}-\varrho^{2}}.$
Then the derivatives satisfy
$\xi_{\varrho}^{\prime}(r)$ $=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},$
$4\xi_{\varrho}^{\prime\prime}(r)$ $=-\frac{1}{r}\,\xi_{\varrho}^{\prime}(r)\left(1+(\xi_{\varrho}^{\prime}(r))^{2}\right).$
The inverse function $\xi_{\varrho}^{-1}$ is given by
$\xi_{\varrho}^{-1}(t)=\left(\varrho^{2}+4\varrho^{2}B^{2}\right)^{2},\qquad B=\sinh\left(\frac{1}{3}\operatorname{arsinh}\left(-\frac{3t}{8\varrho^{4}}\right)\right),$
and its derivatives are
$\dot{\xi}_{\varrho}^{-1}(t)$ $=-2B,$
$\ddot{\xi}_{\varrho}^{-1}(t)$ $=\frac{1}{4\varrho^{4}(1+4B^{2})}.$
[…]
Proof.
A direct computation yields the derivative identities for $\xi_{\varrho}$. For the inverse, we set $r=\varrho^{4}(1+4B^{2})^{2}$ and compute
$\xi_{\varrho}(r)=-\frac{32}{3}\varrho^{4}B^{3}-8\varrho^{4}B.$
Hence, the triple angle formula
$\sinh(3u)=3\sinh(u)+4\sinh^{3}(u)$
implies the claim that $\xi_{\varrho}^{-1}$ inverts $\xi_{\varrho}$. Differentiating $B(t)=\sinh\left(\frac{1}{3}\operatorname{arsinh}\left(-\tfrac{3t}{8\varrho^{4}}\right)\right)$ gives for $u=\left(\frac{1}{3}\operatorname{arsinh}\left(-\frac{3t}{8\varrho^{4}}\right)\right)$
$\dot{B}(t)=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\sqrt{1+\left(\tfrac{3t}{8\varrho^{4}}\right)^{2}}}$
$=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\sqrt{1+\sinh^{2}(3u)}}$
$=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\cosh(3u)}.$
Using the triple angle formula
$\cosh(3u)=4\cosh^{3}u-3\cosh u=\cosh(u)(4(1+\sinh^{2}u)-3),$
we obtain
$\dot{B}(t)=-\frac{1}{8\varrho^{4}(1+4B^{2})},$
from which the derivative identities for $\dot{\xi}_{\varrho}^{-1}$ and $\ddot{\xi}_{\varrho}^{-1}$ follow directly. ∎
Remark A.2.
The function $B$ can also be expressed more explicitly with the help of the formula
$2\sinh\left(\tfrac{1}{3}\operatorname{arsinh}u\right)=\left(u+\sqrt{u^{2}+1}\right)^{1/3}-\left(\sqrt{u^{2}+1}-u\right)^{1/3}$
which follows from repeated usage of the triple angle formulas.
The next result is well-known and we will omit its proofs.
Lemma A.3.
Let $\ell:\mathbb{R}^{n}\to\mathbb{R}$ be a spherically symmetric function, i.e. $\ell(x)=\ell(r)$ where $r=|x|$. Then the principal curvatures of the graph of $\ell$ are
$\lambda_{1}=-\frac{\ell^{\prime\prime}}{(1+\ell^{\prime 2})^{3/2}},\qquad\lambda_{2}=\dots=\lambda_{n}=-\frac{\ell^{\prime}}{r\sqrt{1+\ell^{\prime 2}}}$
where $\lambda_{1}$ is the principal curvature in radial direction.