### v2

# Surgery and total mean curvature

Abstract.
We prove Gromov’s conjecture on the total mean curvature of fill-ins in various cases. Our methods are based on surgery to reduce the statement to fill-ins of spheres, which can be treated by instances of the positive mass theorem. For spin fill-ins, where we permit the mean curvature to take negative values, we build on a classical surgery result of Lawson–Michelsohn and a recent positive mass theorem with creases by Kazaras–Khuri–Lin. For non-spin fill-ins of spin manifolds, where we assume the mean curvature to be non-negative, we develop a novel quantitative surgery process to reduce the general situation to a result of Shi–Wang–Wei. We also treat the case of fill-ins of non-spin manifolds, provided there is a fixed positive lower bound on the mean curvature.
Key words and phrases:
Fill-ins, scalar curvature, total mean curvature, surgery; positive mass theorem
1. Introduction
Let $(M,g_{M})$ be a closed Riemannian manifold, possibly disconnected, with scalar curvature $\mathrm{scal}_{g_{M}}\colon M\to\mathbb{R}$. A fill-in of $(M,g_{M})$ is a compact Riemannian manifold $(\Omega,g_{\Omega})$ such that $(\partial\Omega,g_{\partial\Omega})$ is isometric to $(M,g_{M})$, where $g_{\partial\Omega}$ denotes the metric induced by $g_{\Omega}$ on $\partial\Omega$. We write $H_{g_{\Omega}}\colon\partial\Omega\to\mathbb{R}$ for the mean curvature of $\partial\Omega\subset(\Omega,g_{\Omega})$ with respect to the exterior normal. By convention, this means that the mean curvature of the closed unit ball in flat Euclidean space $\mathbb{R}^{n}$ is equal to n1n-1.
Inspired by the positive mass theorem in general relativity, Gromov conjectured that for fixed $(M,g_{M})$, the total mean curvature of $(\Omega,g_{\Omega})$ cannot be arbitrarily large under a uniform lower scalar curvature bound of $(\Omega,g_{\Omega})$:
Conjecture 1.1 ([Gro23, p. 232]).
Let $(M,g_{M})$ be a closed Riemannian manifold. Then for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda(M,g_{M},\sigma)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$, one has
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq\Lambda(M,g_{M},\sigma).$
The following result verifies this conjecture for various cases, with a constant that also depends on a lower mean curvature bound.
Theorem.
Let $(M,g_{M})$ be a closed Riemannian manifold of dimension $n$ and let $\sigma,\kappa\in\mathbb{R}$. Then there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma,\kappa)>0$ with the following property: If $(\Omega,g_{\Omega})$ is a fill-in of $(M,g_{M})$ such that:
  • $H_{g_{\Omega}}\geq\kappa$,
  • $\mathrm{scal}(g_{\Omega})\geq\sigma$,
and such that one of the following conditions holds:
  1. (A)
    $\kappa>0$ and $n\leq 6$,
  2. (B)
    $\kappa=0$, $n\leq 6$ and $M$ admits a spin structure,
  3. (C)
    $\Omega$ admits a spin structure,
then
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq\Lambda(M,g_{M},\sigma,\kappa).$
The dependence of $\Lambda$ on the Riemannian metric $g_{M}$ is unavoidable, even if $\mathrm{vol}(M)$ and $\mathrm{diam}(M)$ are uniformly bounded, see Fig. 1 and Appendix B. The explicit dependence of $\Lambda$ in terms of $\sigma$ and $\kappa$ is worked out in (3.8).
Figure 1. A Riemannian manifold with large $\Lambda(M,g_{M},0,0)$.
Assuming $\kappa\geq 0$, this result has been proven before in the following cases: By Shi–Wang–Wei [SWW22] if $M$ is diffeomorphic to the $n$-sphere and for orientable fill-ins, based upon Shi–Tam’s proof for the positivity of the Brown–York mass [ST02] and the hyperboloidal positive mass theorem (PMT) [Wan01, CJŁ04, ACG08]; and by Wang [Wan24] for tori $T^{n}$ and fill-ins $D^{2}\times S^{n-1}$, using Brendle–Hung’s [BH24] work on the Horowitz–Myers PMT.
Christian Bär in [Bär26] has presented a different proof of Theorem (C) using spinorial methods. Theorem (C) and Bär’s result provide the first upper bounds on the total mean curvature to treat fill-ins whose mean curvature assumes negative values.
Convention.
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ as in Theorem (A), Theorem (B) or Theorem (C). Then the orientation double cover $(\overline{\Omega},g_{\overline{\Omega}})$ of $(\Omega,g_{\Omega})$ is a fill-in of the orientation double cover of $(M,g_{M})$ and the assumptions in the respective theorem are still satisfied. Furthermore, $\int_{\partial\overline{\Omega}}H_{g_{\overline{\Omega}}}\mathrm{dvol}_{g_{\partial\overline{\Omega}}}=2\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}$. Therefore, in our proofs, we can restrict to oriented $(M,g_{M})$ and $(\Omega,g_{\Omega})$ with an orientation preserving Riemannian isometry $(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$. These assumptions will be used throughout the rest of the paper.
1.1. Proof outline
Our approach is different from the previous ones. By topological arguments, we can assume that we can pass from $M$ to the sphere $S^{n}$ through surgeries of codimension $k\geq 2$. At the same time, we pass from a fill-in of $M$ to a fill-in of $S^{n}$ via the corresponding handle attachments. Working with handles of sufficiently small width and a pigeonhole argument, this ensures that the total mean curvature of the original fill-in of $M$ transfers to the total mean curvature of a fill-in of $S^{n}$, up to a controlled error.
Relying on the classical surgery result of Lawson–Michelsohn [LM84] for positive mean curvature, this strategy can be carried out as stated when restricting to fill-ins $\Omega$ of $M$ with $H_{g_{\Omega}}\geq\kappa$ for some fixed $\kappa>0$. Using the result of Shi–Wang–Wei [SWW22], this leads to a proof of Theorem (A), which we present in Section 2.
The same strategy also applies to Theorem (C), when working with a positive mass theorem for manifolds with creases due to [Lin23, KKL25]. The surgery part is even simpler in this case, as the result of Lawson–Michelsohn is not needed.
From a technical point of view, the proof of Theorem (B) is the most challenging task. The majority of this paper is devoted to addressing this issue. Using surgery, we again reduce the general case to $M=S^{n}$ treated by Shi–Wang–Wei. However, if the pointwise lower mean curvature bound of different fill-ins $\Omega$ is not uniformly bounded below by a positive constant, the handles have to be chosen thinner and thinner to ensure that the mean curvature for the resulting fill-in of $S^{n}$ is nonnegative, which is required in [SWW22]. This leads to infinitely many metrics on the resulting $S^{n}$ and consequently, Shi–Wang–Wei’s result does not yield a uniform total mean curvature upper bound.
This difficulty is overcome by the following strategy (see Section 4): Shi–Wang–Wei’s upper bound can be chosen uniformly for a family of metrics on $S^{n}$, provided that each member of the family can be connected to a fixed metric on $S^{n}$ along a monotonically increasing path of Riemannian metrics with lower scalar curvature bounds that are independent of the given metric in the family (see Proposition 4.2 for the precise statement). We will develop a quantitative surgery construction, see Section 1.2, that provides the required control of the surgery handles. Proceeding by induction on the number of surgery steps finishes the proof for Theorem (B).
Remark 1.2.
Theorem (A) and Theorem (B) hold for any $n$ for which the hyperboloidal PMT holds for oriented manifolds in dimension $n+1$.
1.2. Quantitative surgery
Our main tool for establishing Theorem (B) is a new quantitative surgery procedure. This encompasses and strengthens the classical surgery result of Gromov–Lawson [GL80] and has connections to the extrinsic surgery result of Lawson–Michelsohn [LM84].
To explain this construction, let $\mathcal{S}$ be a compact submanifold in $(M,g_{M})$ of codimension $k$, and let $\varrho>0$ be a small parameter. A Gromov–Lawson handle of size $\varrho$ is a hypersurface $\Sigma_{\varrho}$ contained in $(M\times[0,1],dt^{2}+g_{M})$ which coincides with $M=M\times\{0\}$ outside the $2\sqrt{\varrho}$-tubular neighbourhood of $\mathcal{S}$ in $M$, becomes cylindrical of width $\varrho^{4}$ at height $2\varrho$ and is invariant under the fibrewise $\mathrm{O}(k)$-symmetry of the normal bundle of $\mathcal{S}$ in $M$, see Fig. 2.
Let $\mathrm{scal}_{\varrho}$ be the scalar curvature of $\Sigma_{\varrho}$ and let $H_{\varrho}$ be the mean curvature of $\Sigma_{\varrho}$. Furthermore, let $s$ be the minimum of the scalar curvature of $(M,g_{M})$ on $\mathcal{S}$. The crucial properties of our construction are:
  • if $k\geq 2$, then $(H_{\varrho})_{-}$ converges to zero as $\varrho\to 0$,
  • if $k\geq 3$, then $(\mathrm{scal}_{\varrho}-s)_{-}$ converges to zero as $\varrho\to 0$.
By performing a conformal transformation arbitrarily close to the identity, we can ensure that the mean curvature is positive for sufficiently small $\varrho$.
The detailed construction in the special case where $\mathcal{S}$ is a point in flat Euclidean space $\mathbb{R}^{n}$ is carried out in Section 5. For general $M$ and $\mathcal{S}$, error terms arise, which will be controlled in Section 6.
[…]
Figure 2. A cross-section of a family of Gromov–Lawson handles $\Sigma_{\varrho}$ attached to $\Omega$ along $M=\partial\Omega$. The origin corresponds to the intersection of the cross-section with $\mathcal{S}$.
Our construction moreover provides precise control between handles of different sizes $\varrho$. In particular, the heuristic observation (see Figure 2) that the metrics on $\Sigma_{\varrho}$ are monotonically increasing in $\varrho$ can be made rigorous, while maintaining a uniform lower bound on the scalar curvature of $\Sigma_{\varrho}$. The last property requires working with handles of codimension at least $3$, which accounts for the spin assumption for $M$ in Theorem (B).
Our computations are fully explicit, and the handles are constructed from first principles, yielding an independent approach to [GL80, Theorem A] with explicit geometric bounds. Consequently, our results also encompass variants of the conclusions of several earlier works, including [Swe25, Swe25a, BDS18, Dod20, BS21, BKS20]; compare Remark 7.2.
Our results also allows us to define a new quasi-local mass in general relativity with several particularly appealing properties. This will be discussed in Section 8.
Acknowledgements: BH and SH wish to thank Simon Brendle for useful discussions and his interest in this work. BH and SH are grateful to the University of Augsburg, to Columbia University, to the Lonavala Geometry Festival, and to the Chennai Mathematical Institute for their hospitality and ideal working conditions. BH acknowledges financial support by Columbia University and the DFG-SPP 2026 “Geometry at Infinity”. SH was supported in part by an AMS–Simons Travel Grant (award no. 210022).
2. Surgery for fill-ins with fixed lower mean curvature bounds
Applying a surgery procedure, we will reduce the general case of Theorem (A) to (oriented) fill-ins of the $n$-sphere. Since this case has been treated in [SWW22], this finishes the proof of Theorem (A). A similar argument applies to Theorem (C), where the case of spheres will be postponed to Section 3.
Proof of Theorem (A).
Let $(M,g_{M})$ and $\sigma\in\mathbb{R},\kappa>0$ be as in Theorem (A). Put $n:=\dim M$. By taking products with standard round spheres, we can assume without loss of generality that $n\geq 5$. Let $B$ be a fixed compact oriented manifold such that $\partial B=M$ and so that $B$ induces the reverse orientation on $M$.
Removing a small open disc from the interior of $B$, we obtain a compact oriented bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$ and $\pi_{1}(W)$ by surgeries in the interior of $W$ (using $n\geq 5$), we can assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $1$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $2$.
Let $\ell\geq 0$ be the number of handles in this handle decomposition of $W$. The proof of Theorem (A) is by induction on $\ell$. The case $\ell=0$ is covered by [SWW22, Theorem 4.1 and Remark 4.2]. Now assume by induction that Theorem (A) holds for $\ell\geq 0$ and that the number of handles in $W$ is equal to $\ell+1$.
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq\kappa$. Let $\omega\colon(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$ be an (orientation preserving) Riemannian isometry. For simplicity of exposition, we will assume that $\partial\Omega=M$ and $\omega=\operatorname{\mathrm{id}}$. In the general case, the isometry $\omega$ has to be inserted when gluing subsets of $B$ to $\Omega$. This implies that $\Omega\cup_{M}B$ is oriented.
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 2$, be the surgery datum for the first handle $T$ in $W$ and let $\widetilde{M}=\partial(\Omega\cup T)$ be obtained from $M$ by surgery along $\psi$. Note that the induction hypothesis applies to $\widetilde{M}$. There exist smooth embeddings $\psi_{i}\colon S^{n-k}\times D^{k}\hookrightarrow M$, $i=1,2$, that are both isotopic to $\psi$ and have disjoint images, see Fig. 3. With this “doubling construction” we can estimate the total mean curvature after the surgery by a pigeonhole argument, see (2.3) below.
[…]
Figure 3. The disjoint embeddings $\psi_{1}$ and $\psi_{2}$ which are isotopic to $\psi$ as small displacements.
Choose a smooth collar $M\times[0,1)\hookrightarrow W$ of $M\subset\partial W$ in $W$. With respect to this collar, we identify $M\times[0,1)$ with a smooth submanifold of $W$. Let $t$ be the canonical coordinate on $[0,1)$ and choose a Riemannian metric $h$ on $W$ which restricts to the metric
(2.1) $dt^{2}+\left(1+\tfrac{2t}{n}\kappa\right)g_{M}\text{ on }M\times[0,1)\subset W,$
Note that $(W,h)$ only depends on $(M,g_{M})$, but not on the fill-in $(\Omega,g_{\Omega})$. The second fundamental form of $(W,h)$ along $M\times\{0\}$ with respect to the exterior normal is equal to $\tfrac{\kappa}{n}\cdot g_{M}$.
From now on, let $i\in\{1,2\}$. By [LM84, Theorem 3.1], we can find a submanifold $T_{i}\subset W$ which intersects $M$ in $\mathrm{im}(\psi_{i})$, which is isotopic to the handle attached along the surgery datum $\psi_{i}$ and such that the mean curvature of
(2.2) $\widetilde{M}_{i}:=\partial\left(\Omega\cup T_{i}\right)$
is bounded below by some constant $0<\widetilde{\kappa}<\kappa$, for $i=1,2$, see Fig. 4. Here $\partial\left(\Omega\cup T_{i}\right)$ is a smooth submanifold of $W$ and the mean curvature is computed with respect to the metric $h$ and the normal pointing inside $W$.
[…]
Figure 4. The bordism $W$ consisting of one handle and the embedded handle
Let $g_{\widetilde{M}_{i}}$ be the metric on $\widetilde{M}_{i}$ induced by $h$. The Riemannian manifold $(\widetilde{M}_{i},g_{\widetilde{M}_{i}})$ is independent of the fill-in $(\Omega,g_{\Omega})$. Furthermore, since $\widetilde{M}_{i}$ is diffeomorphic to $\widetilde{M}$, the induction hypothesis applies to $\widetilde{M}_{i}$.
Set $\widetilde{\Omega}_{i}:=\Omega\cup T_{i}\subset\Omega\cup W$. We have $\partial\widetilde{\Omega}_{i}=\widetilde{M}_{i}$. Let $Z_{i}$ be open neighborhoods of $\mathrm{supp}(\psi_{i})\subset M$ such that $Z_{1}\cap Z_{2}=\emptyset$. Further below, we will construct Riemannian metrics $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ such that
  1. (i)
    the metric on $\widetilde{M}_{i}$ induced by $g_{\widetilde{\Omega}_{i}}$ coincides with $g_{\widetilde{M}_{i}}$,
  2. (ii)
    $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$,
  3. (iii)
    $H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$ everywhere,
  4. (iv)
    $\mathrm{scal}(g_{\widetilde{\Omega}_{i}})>{\rm min}\{\sigma,\min\mathrm{scal}_{h}\}-1$.
Using $Z_{1}\cap Z_{2}=\emptyset$, properties (ii), (iii) and (iv), and $\widetilde{\kappa}>0$, we obtain
(2.3) $\begin{split}\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq{}&\int_{\partial\Omega\setminus Z_{1}}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}+\int_{\partial\Omega\setminus Z_{2}}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\\ \leq{}&\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\\ \leq{}&2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}.\end{split}$
By the inductive assumption, for $i\in\{1,2\}$, Theorem (A) holds for the manifold $(\widetilde{M}_{i},\widetilde{g}_{i})$, the lower scalar curvature bound ${\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ and the lower mean curvature bound $\widetilde{\kappa}>0$. Since all of these data are independent of the fill-in $(\Omega,g_{\Omega})$, this concludes the induction step.
It remains to construct the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. Let $U_{i},V_{i}$, $i=1,2$, be open neighborhoods of $\mathrm{im}(\psi_{i})\subset M$ such that $\overline{U}_{i}\subset V_{i}$, $\overline{V}_{i}\subset Z_{i}$. By [BH23, Theorem 3.7], we can smoothly deform the metric $g_{\Omega}$ on $\Omega$ to a new metric on $\Omega$, called $g_{i}$ for short, with the following properties (see Fig. 5):
  • The metrics on $\partial\Omega$ induced by $g_{i}$ coincide with $g_{M}$,
  • $H_{g_{i}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$,
  • $H_{g_{i}}\geq\kappa$ everywhere,
  • On $V_{i}$, the second fundamental form of $g_{i}$ is given by $\mathrm{II}_{g_{i}}=\tfrac{\kappa}{n}g_{M}$,
  • $\mathrm{scal}_{g_{i}}>\sigma-1$.
[…]
Figure 5. Deformation of $g_{\Omega}$ near $\partial\Omega$ (grey region)
Let $\delta>0$ and let $\mathscr{C}_{i}:=\partial\Omega\times(-\delta,0]\subset\Omega$ be an open collar of $\partial\Omega$ in $\Omega$ induced by the normal exponential map along $\partial\Omega$ for $g_{i}$. Put $\mathscr{V}_{i}:=V_{i}\times(-\delta,0]\subset\mathscr{C}_{i}$. Equip the topological manifold
$\mathscr{V}_{i}\cup_{V_{i}\times 0}(V_{i}\times[0,1))=V_{i}\times(-\delta,1)\subset\Omega\cup W$
with the product smooth structure. We have
$\widetilde{\Omega}_{i}\subset\Omega\cup(U_{i}\times(-{\delta},1))\cup{\rm int}(W).$
By construction, the given smooth structures on mutual intersections are compatible with each other. Hence, we obtain an induced smooth structure on $\widetilde{\Omega}_{i}$.
Extend the restriction $h|_{V_{i}\times[0,1)}$ to a smooth metric $h_{i}^{\prime}$ on $V_{i}\times(-\delta,1]$ in an arbitrary way. The $1$-jets of the metrics $g_{i}$ and $h_{i}^{\prime}$ along $V_{i}\times\{0\}\subset V_{i}\times(-\delta,1)$ coincide by the construction of $g_{i}$. Let $\chi_{i}\colon\partial\Omega\to[0,1]$ be a smooth cut-off function that is equal to $1$ on $U_{i}$ and equal to 0 on $\partial\Omega\setminus V_{i}$. Consider the smooth path of symmetric $(0,2)$-tensor fields $\psi_{i}\colon[0,1]\to C^{\infty}(T^{*}\mathscr{C}_{i}\otimes T^{*}\mathscr{C}_{i})$,
$\psi_{i}(t)(x,r):=(1-\chi_{i}(x)\cdot t)g_{i}+\chi_{i}(x)\cdot t\cdot h_{i}^{\prime}.$
Since $\chi_{i}(x)=0$ for $x\in\partial\Omega\setminus V_{i}$, this is well-defined. Possibly after passing to a smaller $\delta$ (hence to a smaller $\mathscr{C}_{i}$), the path $\psi_{i}$ consists of Riemannian metrics on $\mathscr{C}_{i}$, see Fig. 6.
[…]
Figure 6. The local deformation $\psi_{i}$ for $t=1$.
By the choice of $\chi_{i}$, the $1$-jet of $\psi_{i}(t)$ along $\partial\Omega$ is independent of $t$. In particular, the scalar curvature of $\psi_{i}(t)$ along $\partial\Omega$ is given by
$\mathrm{scal}_{\psi_{i}(t)}=(1-\chi_{i}(x)\,t)\cdot\mathrm{scal}_{g_{i}}+\chi_{i}(x)\,t\cdot\mathrm{scal}_{h^{\prime}_{i}}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1.$
By continuity, possibly after passing to a smaller $\delta$, we have, on $\mathscr{C}_{i}$,
$\mathrm{scal}_{\psi_{i}(t)}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1,\quad t\in[0,1].$
We apply the local flexibility lemma [BH22, Theorem 1.2] to the local deformation $\psi_{i}$ on the open neighborhood $\mathscr{C}_{i}$ of $\partial\Omega$ in $\Omega$ and to the second order open partial differential relation of being a Riemannian metric with $\mathrm{scal}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1$ on symmetric $(0,2)$-tensor fields in $C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$. This results in a smooth path $\Psi_{i}\colon[0,1]\to C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$ of Riemannian metrics on $\Omega$ starting at $g_{i}$, and there exists some $0<\delta^{\prime}<\delta$, such that the restriction of $\Psi_{i}(t)$ to $\partial\Omega\times(-\delta^{\prime},0]$ coincides with $\psi_{i}(t)$, see Fig. 7.
[…]
Figure 7. The end of the global deformation $\Psi_{i}(1)$ obtained by applying the local flexibility lemma to the local deformation $\psi_{i}$.
The metric $\Psi_{i}(1)$ on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, the metric $h_{i}^{\prime}$ on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-\delta^{\prime},1))$ and the metric $h$ on $\widetilde{\Omega}_{i}\cap W$ are compatible with each other. Hence, the union of these metrics defines a smooth metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. By construction, the following holds:
  • the metric on $\partial\widetilde{\Omega}_{i}$ induced by $g_{\widetilde{\Omega}_{i}}$ coincides with $g_{\widetilde{M}_{i}}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{i}}=H_{g_{\Omega}}$ on $\Omega\setminus Z_{i}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}\geq\kappa$ on $Z_{i}\setminus U_{i}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$ everywhere else on $\widetilde{M}_{i}$, by the choice of $T_{i}$,
  • $\mathrm{scal}_{g_{\widetilde{\Omega}_{i}}}>{\rm min}\{\sigma,\min\mathrm{scal}_{h}\}-1$.
Therefore, $g_{\widetilde{\Omega}_{i}}$ possesses all the required properties. This finishes the proof of Theorem (A). ∎
In preparation for Section 3, we prove the following result by a similar argument.
Lemma 2.1.
Assume that Theorem (C) holds in the case when $M$ is diffeomorphic to the $n$-sphere. Then Theorem (C) holds in general.
Proof.
Let $(M,g_{M})$ and $\sigma,\kappa\in\mathbb{R}$ be as in Theorem (C). Put $n:=\dim M$. As before, we can assume $n\geq 5$. Since $M$ carries only finitely many spin structures, we can assume without loss of generality that $M$ is equipped with a fixed spin structure and we are only considering spin fill-ins $\Omega$ of $M$ with isometries $\partial\Omega\approx M$ that are spin structure preserving. Let $B$ be a spin manifold with $\partial B=M$ such that $B$ induces the inverse spin structure on $M$.
Removing a small open disc from the interior of $B$, we obtain a compact spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$ and $\pi_{1}(W)$ by surgeries in the interior of $W$ (using $n\geq 5$), we can assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $1$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $2$.
Let $\ell\geq 0$ be the number of handles in this handle decomposition. The proof of Theorem (C) is by induction on $\ell$. The case $\ell=0$ is covered by the assumption of Lemma 2.1. Assume by induction that Theorem (C) holds for $\ell\geq 0$ and that the number of handles of $W$ is equal to $\ell+1$.
Let $(\Omega,g_{\Omega})$ be a spin fill-in of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$, $H_{g_{\Omega}}\geq\kappa$. Let $\omega\colon(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$ be a spin preserving Riemannian isometry. For simplicity of exposition, we will assume from now on that $\partial\Omega=M$ and $\omega=\operatorname{\mathrm{id}}$. This implies that $\Omega\cup_{M}B$ is a spin manifold.
We now proceed as in the proof of Theorem (A), but with one difference: For the construction of the surgery handles $T_{i}\subset W$ before (2.2), intersecting $M$ in $\mathrm{im}(\psi_{i})$, we do no longer refer to [LM84]. This implies that the lower bound $\widetilde{\kappa}$ for the mean curvatures of $\widetilde{M}_{i}$, $i=1,2$, in (2.2) may become negative. Note that, as before, $\widetilde{\kappa}$ does not depend on the fill-in $(\Omega,g_{\Omega})$ and that we can assume $\widetilde{\kappa}<\kappa$.
Set $\widetilde{\Omega}_{i}:=\Omega\cup T_{i}$. Note that $\widetilde{\Omega}_{i}$ carries an induced spin structure. By the construction below Eq. 2.3, which does not require $\widetilde{\kappa}>0$, we obtain metrics $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ with properties (i), (ii), (iii) and (iv). Using that $Z_{1}\cap Z_{2}=\emptyset$ and that $H_{g_{\Omega}},H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$, we hence obtain
(2.4) $\begin{split}&\int_{\partial\Omega}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}\\ &\qquad\leq{}\int_{\partial\Omega\setminus Z_{1}}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}+\int_{\partial\Omega\setminus Z_{2}}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}\\ &\qquad\leq{}\int_{\partial\widetilde{\Omega}_{1}}(H_{g_{\widetilde{\Omega}_{1}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}(H_{g_{\widetilde{\Omega}_{2}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\\ &\qquad\leq{}2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}(H_{g_{\widetilde{\Omega}_{1}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}(H_{g_{\widetilde{\Omega}_{2}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}.\end{split}$
Here we take advantage of the fact that all integrands are non-negative.
Since $\widetilde{\kappa}$ and the metrics on $\partial\Omega$ and $\partial\widetilde{\Omega}_{i}$, $i=1,2$, are independent of $(\Omega,g_{\Omega})$, we obtain a constant $C$, which is independent of $(\Omega,g_{\Omega})$, such that
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq{}2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}+C.$
This concludes the induction step for Theorem (C). ∎
3. Spin fill-ins and creased initial data sets
This section concludes the proof of Theorem (C). According to Lemma 2.1, the only remaining case is when $M$ is diffeomorphic to the $n$-sphere. Let $(\Omega_{0},g_{0})$ be a spin fill-in of $(M,g_{M})$ such that
  • $\mathrm{scal}(g_{0})\geq\sigma_{0}$ for some $\sigma_{0}<0$,
  • $H_{g_{0}}\geq\kappa$ along $M$, w.r.t. the outward-pointing normal, for some $\kappa<0$.
In this section we adopt the sign-convention commonly used in general relativity, whereby the mean curvatures, denoted by $H(M\subseteq W)$, are computed with respect to the “infinity-pointing normal”, that is if $W\colon M\leadsto N$ is a cobordism from $M$ to $N$, then $H(M\subseteq W)$ is computed with respect to the interior unit normal, whereas $H(N\subseteq W)$ is computed with respect to the exterior unit normal.
Now, let $M,N$ be diffeomorphic to the sphere and consider the positive Lipschitz continuous function $h$ on $M$ given by
(3.1) $h_{0}=|\kappa|\sinh(1)+\frac{1}{3}H(M\subseteq\Omega_{0})_{+}>0.$
After possibly Replacing $h_{0}$ by a smooth approximation we may without loss of generality assume that $h_{0}$ is smooth. According to the proof of [SWW22, Lemma 3.1 and Theorem 4.1], there exists a Riemannian metric $g_{1}$ on $M\times[0,1]$ such that the cobordism $(\Omega_{1}=M\times[0,1],g_{1})$ connects $(M\approx S^{n},g_{M})$ to $N:=S^{n}_{r_{0}}$, equipped with the round metric $g_{\circ,{r_{0}}}$ of radius ${r_{0}}\gg 1$ independent of $h$, with the following properties:
  1. (i)
    $H(M\subseteq\Omega_{1})=h_{0}$ and $H(N\subseteq\Omega_{1})>0$ for the mean curvatures of $M,N$,
  2. (ii)
    $\mathrm{scal}(g_{1})=\sigma_{1}$ for some constant $\sigma_{1}<0$ independent of the function $h_{0}$,
  3. (iii)
    the total mean curvatures are related by $\int_{M}H(M\subseteq\Omega_{1})\leq\int_{N}H(N\subseteq\Omega_{1})$.
We note that the final inequality together with $H(M\subseteq\Omega_{1})\geq\tfrac{1}{3}H(M\subseteq\Omega_{0})$ implies
(3.2) $\int_{M}H(M\subseteq\Omega_{0})\leq 3\int_{N}H(N\subseteq\Omega_{2}).$
We define constants $c_{i}$, $i=0,1,2$ via
$c_{0}=\max\left\{4|\kappa|,\sqrt{\frac{n}{n+1}|\sigma_{0}|},(\cosh 1)^{-1}\sqrt{\frac{n}{n+1}|\sigma_{1}|}\right\},\qquad c_{1}=c_{2}=\cosh(1)c_{0},$
and we set
$\sigma_{2}=-c_{2}^{2}\frac{n+1}{n},$
which again is independent of $h_{0}$. We observe that
$c_{0}\geq 4|\kappa|,\qquad|\sigma_{2}|>\max\{|\sigma_{0}|,|\sigma_{1}|\},\qquad c_{i}^{2}\frac{n+1}{n}\geq-\sigma_{i}\qquad\text{for }i=0,1,2.$
As in the proof of [SWW22, Lemma 3.2 and Theorem 4.1], there exists an asymptotic hyperboloidal extension $(\Omega_{2},g_{2})$ of $N$ with $H(N\subseteq\Omega_{1})=H(N\subseteq\Omega_{2})$ and $\mathrm{scal}(g_{2})=\sigma_{2}$. Note that $\sigma_{2}=\sigma_{1}$ in [SWW22], while in our setting $\sigma_{2}<\sigma_{1}$ to compensate for the mean curvature jump at $M$. Consider now the manifold $(\Omega,g):=(\Omega_{0},g_{0})\cup_{M}(\Omega_{1},g_{1})\cup_{N}(\Omega_{2},g_{2})$ with the above choices for $h_{0}$ and $\sigma_{2}$ which has two corners at $M$ and $N$. The situation is illustrated in Fig. 8.
[…]
Figure 8. At $M$ the mean curvature jumps from negative to positive. Treating $\Omega$ as an initial data set $(\Omega,g,k)$, this jump will have a favorable sign if $k$ is chosen appropriately, with an appropriate jump at $M$, which requires the scalar curvature of the asymptotic hyperboloidal extension to be sufficiently negative.
Similarly to [SWW22, Theorem 4.1 and p. 232–233] and [ST02], the functional
(3.3) $m(r)=\frac{1}{n\omega_{n}}\int_{\Sigma_{r}}\big(H(\Sigma_{r}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})-H(\Sigma_{r}\subseteq\Omega_{2})\big)\cosh(r\sqrt{-K})$
is monotonically increasing, and $m(r)\to m$ for $r\to\infty$. Here $K=\tfrac{\sigma_{2}}{n(n+1)}$ (=1=1 in [SWW22]), and $H(\Sigma_{r}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})=n\sqrt{-K}\coth(r\sqrt{-K})$ is the mean curvature of a sphere in hyperbolic space $\mathbb{H}^{n+1}_{\sigma_{2}}$ with constant scalar curvature $\sigma_{2}$. Moreover, $\Sigma_{r}$ denotes the $(r-r_{0})$-distance spheres to $N$ with respect to the hyperbolic background metric, e.g. we $\Sigma_{r_{0}}=N=S^{n}_{r_{0}}$. Finally, the hyperboloidal mass $m$ will be discussed in (3.4).
In view of (3.2) and (3.3), it suffices to prove $m\geq 0$ in order to establish Theorem (C). However, since $\kappa<0$ and $h_{0}>0$, there may be an unfavorable jump of the mean curvature at $M$, which corresponds to “infinitely negative” scalar curvature, as depicted in Fig. 8. This can be remedied by introducing a symmetric $(0,2)$-tensor $k$ on $\Omega$ which also jumps at $M$. With a suitable choice of $k$, the dominant energy condition for the resulting initial data set $(\Omega,g,k)$ is satisfied weakly across $M$. Hence, the hyperboloidal positive mass theorem can be applied in this more general context.
With the above choice of constants $c_{0},c_{1},c_{2}$, define the symmetric $(0,2)$-tensor $k_{i}=\tfrac{c_{i}}{n}g_{i}$ on $\Omega_{i}$ and let $k=k_{0}\cup k_{1}\cup k_{2}$. Consequently, we obtain an initial data set $(\Omega,g,k)$ satisfying the dominant energy condition
$\mu-|J|=\frac{1}{2}\left(\mathrm{scal}_{i}+c^{2}_{i}\frac{n+1}{n}\right)\geq\frac{1}{2}\left(\mathrm{scal}_{i}-\sigma_{i}\right)\geq 0\qquad\text{for $i=0,1,2$},$
where the energy density $\mu$ and momentum density $J$ are given by
$\mu=\frac{1}{2}\left(\mathrm{scal}+\operatorname{tr}_{g}(k)^{2}-|k|^{2}\right),\qquad J=\operatorname{div}(k-\operatorname{tr}_{g}(k)g).$
It remains to check whether the dominant energy condition of $(\Omega,g,k)$ holds weakly across $M$.
Given $\vartheta\in\mathbb{R}$, define the hyperbolic rotation matrix
$F_{\vartheta}=\begin{pmatrix}\cosh\vartheta&-\sinh\vartheta\\ -\sinh\vartheta&\cosh\vartheta\end{pmatrix}$.
Consider the spacetime mean curvature vectors
$\vec{H}(M\subseteq\Omega_{i})=\begin{pmatrix}H(M\subseteq\Omega_{i})\\ \operatorname{tr}_{M}(k_{i})\end{pmatrix}=\begin{pmatrix}H(M\subseteq\Omega_{i})\\ c_{i}\end{pmatrix}$.
Using the definition of $c_{1}$, we compute for $\vartheta=-1$
$F_{\vartheta}\vec{H}(M\subseteq\Omega_{0})-\vec{H}(M\subseteq\Omega_{1})$
$\qquad=\begin{pmatrix}\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}\\ \cosh(-1)c_{0}-\sinh(-1)H(M\subseteq\Omega_{0})\end{pmatrix}-\begin{pmatrix}H(M\subseteq\Omega_{1})\\ c_{1}\end{pmatrix}$
$\qquad=\begin{pmatrix}\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})\\ -\sinh(-1)H(M\subseteq\Omega_{0})\end{pmatrix}$.
Let $x\in M$ where $H(M\subseteq\Omega_{0})>0$. Using $c_{0}\geq 4|\kappa|$ and $2\sinh(1)>\cosh(1)$, we estimate in this case at $x$
$\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})$
$\qquad\geq\cosh(1)H(M\subseteq\Omega_{0})+(4-1)\sinh(1)|\kappa|$
$\qquad\geq\sinh(1)|H(M\subseteq\Omega_{0})|.$
Next, let $x\in M$ where $H(M\subseteq\Omega_{0})\geq 0$. Using that $\cosh(1)-\sinh(1)=e^{-1}>\tfrac{1}{3}$, we estimate in this case at $x$
$\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})$
$\qquad\geq\cosh(1)H(M\subseteq\Omega_{0})+(4-1)\sinh(1)|\kappa|-\frac{1}{3}H(M\subseteq\Omega_{0})$
$\qquad\geq\sinh(1)|H(M\subseteq\Omega_{0})|.$
This demonstrates that inequality (1.7) in [KKL25] holds at $M$, i.e. the dominant energy condition is weakly satisfied across $M$. We point out, using the notation of [KKL25], that $\nu_{+},\tau_{+}$ correspond to $\binom{1}{0},\binom{0}{1}$, and that the terms $\beta_{\pm}=k(\nu,\cdot)|_{TM}$, $\beta^{\Delta}=\beta_{+}-\beta_{-}-d\vartheta$ vanish, where $\nu$ denotes the infinity-pointing normal to $M$.
It remains to justify that the hyperboloidal positive mass theorem (PMT) can indeed now be applied, which is needed to resolve Gromov’s total mean curvature conjecture in the spin case. To this end, we first recall the standard hyperboloidal PMT for initial data sets. In what follows, let $b$ denote the standard $(n+1)$-dimensional hyperbolic metric with scalar curvature n(n+1)-n(n+1). We write $\mathcal{S}$ for the spinor bundle of the spin manifold $\Omega$, and set $\overline{\mathcal{S}}=\mathcal{S}\oplus\mathcal{S}$ for the corresponding spacetime spinor bundle, cf. [HZ25a, Definition 2.6], which admits a Clifford action of Minkowski space $\mathbb{R}^{n+1,1}=\operatorname{span}\{e_{0},e_{1},\dots,e_{n+1}\}$.
Theorem 3.1 ([CH04] and [HJZ25, Proposition 2.9]).
[…]
$\mathcal{H}_{hyp}(V)=4\int_{\Omega}\Big(\big|\nabla_{i}\psi+\frac{1}{2}k_{ij}e_{j}e_{0}\psi\big|^{2}+\frac{1}{2}\big\langle\psi,(\mu+Je_{0})\psi\big\rangle\Big)\geq 0,$
where $V=|\psi^{\infty}|^{2}_{b}$ and the hyperboloidal mass functional $\mathcal{H}_{hyp}(V)$ is defined in [HJZ25, p. 307].
By appropriate rescalings ($g_{ij}\to\lambda^{2}g_{ij}$, $k_{ij}\to\lambda k_{ij}$) an analogous result holds for initial data sets $(\Omega,g,k)$ asymptotic to $\mathbb{H}^{n+1}_{\sigma_{2}}$ instead of the standard hyperbolic space $\mathbb{H}^{n+1}=\mathbb{H}^{n+1}_{n(n+1)}$, cf. [SWW22, p. 235]. See [CJŁ04, CH04] and [HJZ25, Appendix A] for a detailed discussion and for further references on the hyperboloidal PMT. We emphasize that this result should not be confused with the asymptotically AdS PMT, in which the metric $g$ also asymptotes to the hyperbolic metric but $k$ asymptotes to zero [CMT06, HZ25].
In our situation, the initial data set has significantly more restrictive asymptotics: we have $k=c_{2}g$ at infinity, and the Bartnik extension metric has a very specific form. As pointed out in [CH04, p. 232 and Section 3] as well as [BW12, Appendix A], under these so-called Wang asymptotics the mass
(3.4) $m=\mathcal{H}_{hyp}(V_{0}),\qquad V_{0}=V_{0}(\psi_{0}^{\infty})$
reduces to the mass defined in [Wan01], for a certain choice of Killing spinor $\psi^{\infty}_{0}$. In view of (3.3), this in particular implies $m(r_{0})\geq 0$, which yields the desired total mean curvature bound.
Finally, we explain why the corners at $M$ and $N$ do not pose difficulties. The type of corner arising at $N$ is well understood; see, for instance, [BQ08]. We therefore focus on the corner along $M$. Strictly speaking, the results of [Lin23, KKL25] are formulated for asymptotically flat spin manifolds. However, their arguments are entirely local, and it is straightforward to see that they extend to the asymptotically hyperboloidal spin setting as well. More precisely, given a spinor $\psi$ on $N$, let $\psi_{\pm}$ denote its restrictions to $\Omega_{+}:=\Omega_{1}\cup_{N}\Omega_{2}$ and $\Omega_{-}:=\Omega_{0}$, respectively. Recall Witten’s standard divergence identity,
[…]
Integrating this identity by parts on both $\Omega_{-}$ and $\Omega_{+}$, and then imposing the transmission boundary condition [KKL25, Equation (1.9)]
(3.5) $\psi^{-}=\cosh\frac{\vartheta}{2}\,\psi^{+}+\sinh\frac{\vartheta}{2}\,\nu^{+}e_{0}^{+}\psi^{+},$
[…]
We note that the hyperboloidal positive mass theorem can in fact be reduced to the asymptotically flat case, as shown in [CD19]. Thus, the results of [KKL25] can also be applied directly. This finishes the proof of Theorem (C).
The hyperboloidal positive mass theorem also holds for non-spin manifolds, see [ACG08, Sak21, Lun23, CD19a]. However, the non-smooth case treated in [KKL25] does not yet have an analogue in the non-spin setting. Resolving this issue would extend the statement of Theorem (C) to the non-spin case.
We close this section with some remarks on the asymptotics of $\Lambda(M,g_{M},\sigma,\kappa)$ for fixed $(M,g_{M})$. In [Bär26] it is shown that $\Lambda(M,g_{M},\sigma,\kappa)$ can be chosen to depend linearly on $\sqrt{|\sigma|}$ and $|\kappa|$.
Our constant $\Lambda(M,g_{M},\sigma_{0},\kappa)$, in case $M\approx S^{n}$, can be explicitly estimated via
(3.6) $H(N=S^{n}_{r_{0}}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})=n\sqrt{-\frac{\sigma_{2}}{n(n+1)}}\coth\left(r_{0}\sqrt{-\frac{\sigma_{2}}{n(n+1)}}\right),$
and
(3.7) $\sigma_{2}=-\frac{n+1}{n}\cosh(1)^{2}\max\left\{16\kappa^{2},\frac{n}{n+1}|\sigma_{0}|,(\cosh 1)^{-1}\frac{n}{n+1}|\sigma_{1}|\right\}.$
For $\sigma_{0}\gg 1$ or $\kappa\gg 1$, we therefore have $H(N\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})\approx\max\{\sqrt{-\sigma_{0}},-\kappa\}$, which gives the same asymptotics as Bär’s proof.
With the methods developed in the succeeding sections, we can show that the same asymptotics hold for all closed spin manifolds. To justify this, let $(\Omega_{0},g_{0})$ be a spin fill-in of $(M,g_{M})$ such that $\mathrm{scal}_{g_{0}}\geq\sigma_{0}$ and $H_{g_{0}}\geq\kappa$ for some $\sigma_{0},\kappa<0$. Let $h_{0}\colon M\to\mathbb{R}_{>0}$ be the (smoothened) function in (3.1). According to the proof of [SWW22, Lemma 3.1 and Theorem 4.1], there exists a Riemannian metric $g_{1}$ on $M\times[0,1]$ such that the cobordism $(\Omega_{1}=M\times[0,1],g_{1})$ connects $(M,g_{M})$ to $(N:=M,2g_{M})$ satisfying properties (i), (ii) and (iii) above. This results in a spin fill-in $(\Omega=\Omega_{0}\cup_{M}\Omega_{1},g_{\Omega})$ with positive mean curvature along $(N,2g_{M})$, with a crease at $M$ and with scalar curvature equal to $\sigma_{1}=\sigma_{1}(M,g_{M})$.
[…]
  • $g_{\widetilde{\Omega}}=g_{\Omega}=g_{0}$ on $\Omega_{0}$,
  • $\mathrm{scal}(\widetilde{\Omega},g_{\widetilde{\Omega}})\geq\min\{\sigma_{0},C-\ell\}$,
[…]
By (3.2) and (3.3), we thus obtain
$\int_{N}H\big(M\subseteq\Omega_{0})\leq 2^{\ell}\cdot 3\cdot\int_{S^{n}_{r_{0}}}H(S^{n}_{r_{0}}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})$
for some $r_{0}=r_{0}(M,g_{M})>0$ and for $\sigma_{2}$ given by
$\sigma_{2}=-\frac{n+1}{n}\cosh(1)^{2}\max\left\{16\kappa^{2},-\sigma_{0}\frac{n}{n+1},-C+\ell\right\},$
similar to (3.7). Hence, there exists a $C_{0}(M,g_{M})>0$, independent of $\kappa$ and $\sigma_{0}$, (3.6) gives
(3.8) $\Lambda(M,g_{M},\sigma_{0},\kappa)\leq C_{0}(M,g_{M})\cdot\big(\max\{\sqrt{-\sigma_{0}},-\kappa\}+1\big).$
4. Fill-ins of families of metrics
Let $(M,g_{M})$ be a closed oriented Riemannian manifold. In this section we derive a uniform upper bound for the total mean curvatures of fill-ins of $M$ for a certain family of metrics on $M$. The proof is based on arguments in [SWW22].
Definition 4.1.
Let $s\in\mathbb{R}$ such that $\mathrm{scal}_{g}>s$. We define $\mathscr{R}_{g_{M},s}(M)$ as the set of Riemannian metrics $g$ on $M$ with the following property: There exists a smooth path $\zeta_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ of Riemannian metrics starting at $g$ and ending at $g_{M}$ such that, for all $t\in[0,1]$, we have
  1. (i)
    $\zeta^{\prime}_{g}(t)>0$ pointwise on $M$, in the sense of symmetric $(0,2)$-tensor fields,
  2. (ii)
    $\mathrm{scal}_{\zeta_{g}(t)}>s$ pointwise on $M$.
Proposition 4.2.
For any $\sigma,s\in\mathbb{R}$ and $g\in\mathscr{R}_{g_{M},s}(M)$ the following holds: Let $(\Omega_{0},g_{\Omega_{0}})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Then there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on
$\Omega_{1}:=\Omega_{0}\bigcup_{\partial\Omega\approx M}\big(M\times[0,1]\big)$
with the following properties:
  • $g_{\Omega_{1}}=g_{\Omega_{0}}$ outside an arbitrarily small neighborhood of $M\times[0,1]\subset\Omega_{1}$,
  • $g_{\Omega_{1}}$ induces the metric $g_{M}$ on $\partial\Omega_{1}=M\times\{1\}$,
  • $H_{g_{\Omega_{1}}}>0$ along $M\times\{1\}$,
  • $\mathrm{scal}_{\Omega_{1}}\geq\min\{\sigma,s\}-1$,
  • $\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}<\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{M}}$.
Proof.
Let $g\in\mathscr{R}_{g_{M},s}(M)$. Let $\zeta_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ be a smooth path from $g$ to $g_{M}$ that satisfies properties (i) and (ii). Consider the smooth metric $\overline{g}$ on $M\times[0,1]$ defined by
$\overline{g}=dt^{2}+\zeta_{g}(t).$
Let $\overline{A}_{t}$ be the second fundamental form and $\overline{H}_{t}=\operatorname{tr}_{\zeta_{g}(t)}\overline{A}_{t}$ be the mean curvature of $\Sigma_{t}:=M\times\{t\}\subset(M\times[0,1],\overline{g})$ with respect to the unit normal field $\frac{\partial}{\partial t}$. By (i), we have $\overline{A}_{t}>0$, hence $\overline{H}_{t}>0$ and $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$ for $t\in[0,1]$
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Without loss of generality, we can assume that $\partial\Omega=M$. Applying [BH23, Proposition 3.8], we can increase the mean curvature $H_{g_{\Omega}}$ slightly, while replacing $\sigma$ by $\sigma-\tfrac{1}{2}$. Hence, we can assume that $H_{g_{\Omega}}>0$.
On $M\times[0,1]$, consider the Bartnik-type parabolic PDE
$\overline{H}_{t}\frac{\partial u}{\partial t}$ $=u^{2}\Delta_{\zeta_{g}(t)}u+\frac{1}{2}(s-\mathrm{scal}_{\zeta_{g}(t)})u^{3}+\frac{1}{2}(\mathrm{scal}_{\zeta_{g}(t)}-\mathrm{scal}_{\overline{g}})u$
$u(\cdot,0)$ $=\frac{\overline{H}_{0}}{H_{g_{\Omega}}}>0.$
Since, by (ii), we have $s-\mathrm{scal}_{\zeta_{g}(t)}<0$ for $t\in[0,1]$, this equation has a positive solution $u\colon M\times[0,1]\to\mathbb{R}_{>0}$. Consider the Riemannian metric $\widetilde{g}=u^{2}dt^{2}+\zeta_{g}(t)$ on $M\times[0,1]$.
Let $H_{t}=u^{-1}\overline{H}_{t}>0$ be the mean curvature on $\Sigma_{t}$ induced by $\widetilde{g}$ with respect to the unit normal field $u^{-1}\frac{\partial}{\partial t}$. Note that $H_{0}=H_{g_{\Omega}}$. By the computation in the proof of [SWW22, Theorem 4.1] and using that $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$, we have
$\frac{\partial}{\partial t}\int_{\Sigma_{t}}H_{t}\,\rm{dvol}_{\zeta_{g}(t)}>0.$
In particular,
(4.1) $\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\zeta_{g}(1)}>\int_{\Sigma_{0}}H_{0}{\rm dvol}_{\zeta_{g}(0)}=\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}.$
Furthermore, we have $\mathrm{scal}_{\widetilde{g}}=s$.
Put
$\Omega_{1}:=\Omega_{0}\bigcup_{\partial\Omega=M\times\{0\}}M\times[0,1].$
By [BH23, Theorem 4.11] and by the choice of $u(\cdot,0)$, the mean convex singularity at the gluing region $\partial\Omega_{0}=M\times\{0\}$ can be smoothened, that is, there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on $\Omega_{1}$ which is equal to $\widetilde{g}$ near $M\times\{1\}$ and whose scalar curvature is bounded below by $\min\{s,\sigma\}-1$.
The manifold $(\Omega_{1},g_{\Omega_{1}})$ is a fill-in of $(M,g_{M})$. Furthermore, by (4.1), we get
$\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}<\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\zeta_{g}(1)}=\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{M}}.\qed$
5. Surgery bending in Euclidean space
In this section we construct smooth hypersurfaces $\overline{\Sigma}_{\varrho}$ of size $\varrho$ in Euclidean space which we will later use to define surgery bendings with quantitative control. Let $(x_{1},\ldots,x_{k})$ be the standard coordinates on $\mathbb{R}^{k}$ and let $t$ be the standard coordinate on $[0,1]$. Equip $B_{1}^{k}(0)\times[0,1]\subset\mathbb{R}^{\,k+1}$ with the flat metric $g_{\mathbb{R}^{k+1}}=\mathrm{d}x_{1}^{2}+\dots+\mathrm{d}x_{k}^{2}+\mathrm{d}t^{2}$. Let $r:=\sqrt{x_{1}^{2}+\ldots+x_{k}^{2}}\colon\mathbb{R}^{k}\to[0,\infty)$ denote the radial coordinate.
Construction 5.1.
(Bending hypersurface) Let $\varrho\in(0,\frac{1}{10}]$ be fixed and consider the function
$\xi_{\varrho}(r)\coloneqq-\frac{4}{3}\varrho\left(\sqrt{r}-\varrho^{2}\right)^{3/2}-4\varrho^{3}\sqrt{\sqrt{r}-\varrho^{2}}$
from Lemma A.1. Put
$a_{\varrho}\coloneqq-\xi_{\varrho}(2\sqrt{\varrho})>0.$
Fix a cutoff-function $\eta\colon\mathbb{R}\to[0,1]$ such that $\eta\equiv 1$ on $(-\infty,0]$, $\eta\equiv 0$ on $1,\infty)$, and $2\leq\eta^{\prime}\leq 0$, $|\eta^{\prime\prime}|\leq 10$. Define $\eta_{\varrho}(r)=\eta\left(\tfrac{r-\sqrt{\varrho}}{\sqrt{\varrho}}\right)$ and set
$f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr),$
see Fig. 9. Let $G_{\varrho}\subset\mathbb{R}^{k}\times\mathbb{R}$ be the graph of the function
$B_{1}^{k}(0)\setminus B_{\varrho^{4}}^{k}(0)\to\mathbb{R},\qquad(x_{1},\dots,x_{k})\mapsto f_{\varrho}(r).$
[…]
Figure 9. The function ${f}_{\varrho}$ constructed as the interpolation of $\xi_{\rho}$ and the constant function using the cutoff-function $\eta_{\varrho}$.
Finally, put
$\overline{\Sigma}_{\varrho}:=G_{\varrho}\;\bigcup\;\{r=\varrho^{4}\}\times[a_{\varrho},1]\subset B^{k}_{1}(\varrho)\times[0,1].$
Since the inverse of the function $f_{\varrho}|_{[\varrho^{4},\varrho^{2}]}$ is the smooth function $f_{\varrho}(\varrho^{2}),a_{\varrho}]\ni t\mapsto\xi_{\rho}^{-1}(t-a_{\varrho})$ with vanishing derivative at $t=a_{\varrho}$ (see Lemma A.1), $\overline{\Sigma}_{\varrho}$ is a $C^{1}$-hypersurface that is smooth everywhere except $\{t=a_{\varrho}\}$. This $C^{1}$-singularity will be smoothened out at the end of this section.
[…]
Figure 10. The hypersurface $\overline{\Sigma}_{\varrho}$ as the graph of a spherically symmetric function with a cylinder attached.
Remark 5.2.
Let us motivate the choice of the function $\xi_{\varrho}$. The graph of the functions $\ell_{\varrho}(r)=-2\sqrt{2\varrho}\sqrt{r-2\varrho}$ gives rise to spherically symmetric surfaces satisfying
$2\lambda_{1}=\lambda_{2}=\dots=\lambda_{k}$
for their principal curvatures, cf. Lemma A.3. In $n=3$ these are the (spatial) Schwarzschild metrics of mass $m=\varrho$. In this case the scalar curvature vanishes which does not allow us to absorb the error terms appearing in the handle construction. Therefore, we solve the ODE
$4\lambda_{1}=\lambda_{2}=\dots=\lambda_{k}$
which leads to a spherically symmetric surface with strictly positive scalar curvature. The corresponding graph function is given precisely by $\xi_{\varrho}$.
Let $\overline{\nu}_{\varrho}$ be the unit normal vector field of $\overline{\Sigma}_{\varrho}$ and denote by $\overline{\nu}_{\varrho}^{\intercal}$ the projection of $\overline{\nu}_{\varrho}$ onto the tangent space of the slices $B_{1}^{k}(0)\times\{t\}$, i.e.,
(5.1) $\overline{\nu}^{\intercal}_{\varrho}=\overline{\nu}_{\varrho}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle\partial_{t}.$
Note that $\langle\overline{\nu}_{\varrho},\partial_{t}\rangle\geq 0$ since $\overline{\Sigma}_{\varrho}$ is a graph. We have the following properties.
Proposition 5.3.
The hypersurfaces $\overline{\Sigma}_{\varrho}$ from 5.1 satisfy the following properties.
  1. (i)
    $\overline{\Sigma}_{\varrho}$ is $\mathrm{O}(k)$-invariant and satisfies
    $\overline{\Sigma}_{\varrho}\cap(B_{1}^{k}(0)\times[2\varrho,1])$ $=\partial B_{\varrho^{4}}^{k}(0)\times[2\varrho,1],$
    $\overline{\Sigma}_{\varrho}\cap\Bigl(\bigl(B_{1}^{k}(0)\setminus B_{2\sqrt{\varrho}}^{k}(0)\bigr)\times[0,1]\Bigr)$ $=\bigl(B_{1}^{k}(0)\setminus B_{2\sqrt{\varrho}}^{k}(0)\bigr)\times\{0\}.$
  2. (ii)
    For $r\in[\varrho^{4},\sqrt{\varrho}]$, we have
    $H(\overline{\Sigma}_{\varrho})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho})|\quad\text{if }k\geq 2$
    $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}\quad\text{if }k\geq 3.$
  3. (iii)
    For $r\in[\varrho^{4},\varrho^{2}]$, $k\geq 3$, we have
    $\min\left\{H(\overline{\Sigma}_{\varrho})^{2},\mathrm{scal}(\overline{\Sigma}_{\varrho})\right\}>\tfrac{1}{2}\varrho^{-3}.$
For the final two properties let $\overline{C}\coloneqq 18k+{432}$.
  1. (iv)
    For $r\geq\varrho^{2}$, we have $|\overline{\nu}_{\varrho}^{\intercal}|<\overline{C}\,\sqrt{\varrho}$.
  2. (v)
    For $r\geq\sqrt{\varrho}$, we have $|A(\overline{\Sigma}_{\varrho})|<\overline{C}\,\varrho^{3/8}$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>-\overline{C}\,\varrho^{3/4}$.
Proof.
The first assertion follows directly from the construction, the only thing to verify is that $a_{\varrho}<2\varrho$ for $\varrho\leq\frac{1}{10}$:
$a_{\varrho}=-\xi_{\varrho}(2\sqrt{\varrho})=\frac{4}{3}\varrho\underbrace{\left(\sqrt{2\sqrt{\varrho}}-\varrho^{2}\right)^{3/2}}_{<1}+4\varrho^{3}\underbrace{\sqrt{\sqrt{2\sqrt{\varrho}}-\varrho^{2}}}_{<1}<\frac{4}{3}\varrho+\frac{4}{100}\varrho<2\varrho.$
Concerning (ii), we note that on this inner annulus the cutoff function $\eta_{\varrho}$ equals $1$, so $f_{\varrho}=\xi_{\varrho}+a_{\varrho}$ and therefore $f_{\varrho}^{\prime}=\xi_{\varrho}^{\prime}$ as well as $f_{\varrho}^{\prime\prime}=\xi_{\varrho}^{\prime\prime}$. By Lemma A.1, we have
$4\,\xi_{\varrho}^{\prime\prime}=-\frac{1}{r}\,\xi_{\varrho}^{\prime}(1+\xi_{\varrho}^{\prime 2}).$
Combined with Lemma A.3, gives
$\lambda_{2}=\dots=\lambda_{k}=-4\,\lambda_{1}.$
On the graph of $f_{\varrho}$ on $\varrho^{4},\sqrt{\varrho}]$. Therefore, we get $\lambda_{1}=-(16k-15)^{-1/2}|A|$ and thus
$H(\overline{\Sigma}_{\varrho})$ $=\lambda_{1}-4(k-1)\lambda_{1}=\frac{4k-5}{\sqrt{16k-15}}|A|$
$>\frac{1}{2}|A|$ $\text{if }k\geq 2$
$\mathrm{scal}(\overline{\Sigma}_{\varrho})$ $=16(k-1)(k-2)\lambda_{1}^{2}-8(k-1)\lambda_{1}^{2}$
$=8(k-1)(2k-5)\lambda_{1}^{2}$
$=\frac{8\left(k-1\right)\left(2k-5\right)}{16k-15}|A|^{2}$
$>\frac{1}{3}|A|^{2}$ $\text{if }k\geq 3.$
On the other hand, the cylinder of radius $\varrho^{4}$ has principal curvatures: $\lambda_{1}=0$ and $\lambda_{2}=\dots=\lambda_{k}=\varrho^{-4}\not=0$ and therefore
(5.2) $\begin{aligned} H(\overline{\Sigma}_{\varrho})&=\sqrt{k-1}\,|A|>\frac{1}{2}|A|&&\text{if }k\geq 2\\ \mathrm{scal}(\overline{\Sigma}_{\varrho})&=(k-2)|A|^{2}>\frac{1}{3}|A|^{2}&&\text{if }k\geq 3.\end{aligned}$
Next, we show (iii). On the cylindrical part, we have $|A|=\sqrt{k-1}\varrho^{-4}$, so the desired inequality holds by (5.2). On the “graphical part”, we observe that both $\xi_{\varrho}^{\prime}$ and the function $x\mapsto\frac{-x}{\sqrt{1+x^{2}}}$ are monotone decreasing and thus, for $r\leq\varrho^{2}$:
$\lambda_{2}(r)$ $=-\frac{\xi_{\varrho}^{\prime}(r)}{r\sqrt{1+\xi_{\varrho}^{\prime}(r)^{2}}}\geq-\frac{\xi_{\varrho}^{\prime}(\varrho^{2})}{\varrho^{2}\sqrt{1+\xi_{\varrho}^{\prime}(\varrho^{2})^{2}}}$
$=\frac{1}{\varrho^{2}\sqrt{1+\frac{\varrho^{2}}{\varrho-\varrho^{2}}}}\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}=\frac{\sqrt{\varrho-\varrho^{2}}}{\varrho\sqrt{\varrho}}\frac{1}{\sqrt{\varrho-\varrho^{2}}}=\frac{1}{\varrho^{3/2}}.$
Hence, we obtain
$H={}$ $-\frac{1}{4}\lambda_{2}+(k-1)\lambda_{2}=(k-\tfrac{5}{4})\lambda_{2}$
$\geq{}$ $(k-\tfrac{5}{4})\varrho^{-3/2}$
$\mathrm{scal}={}$ $(k-1)(k-2)\lambda_{2}^{2}-\frac{1}{2}(k-1)\lambda_{2}^{2}$
$\geq{}$ $(k-1)(k-\tfrac{5}{2})\varrho^{-3}$
and if $k\geq 3$, then $H>\varrho^{-3/2}$ and $\mathrm{scal}\geq\varrho^{-3}$, which proves (iii).
Next, we show the estimate from (iv) for $r\leq\sqrt{\varrho}$. Since $f_{\varrho}^{\prime}=\xi_{\varrho}^{\prime}$ on $\varrho^{2},\sqrt{\varrho}]$ and thus $f_{\varrho}^{\prime}$ is monotone increasing, we have
$0\geq f_{\varrho}^{\prime}(r)\geq\xi_{\varrho}^{\prime}(\varrho^{2})=-\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}\geq-2\sqrt{\varrho},$
by Lemma A.1. Hence, $|f_{\varrho}^{\prime}(r)|\leq 2\sqrt{\varrho}$ for $r\in[\varrho^{2},\sqrt{\varrho}]$. Moreover,
$|\overline{\nu}_{\varrho}^{\intercal}|=\frac{|\nabla f_{\varrho}|}{\sqrt{1+|\nabla f_{\varrho}|^{2}}}\leq|\nabla f_{\varrho}|=|f_{\varrho}^{\prime}|\leq 2\sqrt{\varrho},$
which proves the claim for $\varrho^{2}\leq r\leq\sqrt{\varrho}$.
Next, let us prove (iv) and (v) for $r\geq\sqrt{\varrho}$. Before we can do so, we need some estimates for the derivatives of $f_{\varrho}$ on $\sqrt{\varrho},2\sqrt{\varrho}]$. To this end, recall that
$\xi_{\varrho}^{\prime}(r)=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},\qquad 4\xi_{\varrho}^{\prime\prime}(r)=-\frac{1}{r}\xi_{\varrho}^{\prime}(r)\left(1+\xi_{\varrho}^{\prime}(r)^{2}\right),$
by Lemma A.1. Furthermore, since $\varrho$ was assumed to be smaller than $\tfrac{1}{10}$, we have
$\varrho^{1/4}-\varrho^{2}=\frac{1}{2}\varrho^{1/4}+\varrho^{1/4}\left(\frac{1}{2}-\varrho^{3/4}\right)\geq\frac{1}{2}\varrho^{1/4},$
and hence
$|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|={}$ $\frac{\varrho}{\sqrt{\varrho^{1/4}-\varrho^{2}}}\leq\sqrt{2}\,\varrho^{7/8},$
$|\xi_{\varrho}^{\prime\prime}(\sqrt{\varrho})|\leq{}$ $\frac{1}{4\sqrt{\varrho}}\,|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|\left(1+\xi_{\varrho}^{\prime}(\sqrt{\varrho})^{2}\right)\leq\frac{\sqrt{2}}{4}\,\varrho^{3/8}\left(1+2\varrho^{14/8}\right)\leq\varrho^{3/8}$
Moreover, $\xi_{\varrho}$ is strictly decreasing and thus, for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$, the fundamental theorem of calculus yields
$0\leq{}$ $\xi_{\varrho}(r)+a_{\varrho}=-\int_{r}^{2\sqrt{\varrho}}\xi_{\varrho}^{\prime}(s)ds$
$\leq{}$ $(2\sqrt{\varrho}-r)|\xi_{\varrho}^{\prime}(r)|\leq\sqrt{\varrho}\,|\xi_{\varrho}^{\prime}(\sqrt{\varrho})|\leq\sqrt{2}\,\varrho^{11/8}.$
Therefore, since $\xi_{\varrho}^{\prime}$ is negative and strictly increasing we obtain for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$,
$|f_{\varrho}^{\prime}(r)|\leq{}$ $|\eta(r)\xi_{\varrho}^{\prime}(r)+\eta_{\varrho}^{\prime}(r)(a_{\varrho}+\xi_{\varrho}(r))|$
$\leq{}$ $-\xi_{\varrho}^{\prime}(\sqrt{\varrho})+\frac{2}{\sqrt{\varrho}}\,(a_{\varrho}+\xi_{\varrho}(\sqrt{\varrho}))$
$\leq{}$ $\sqrt{2}\,\varrho^{7/8}+\frac{2\sqrt{2}}{\sqrt{\varrho}}\,\varrho^{11/8}<3\sqrt{2}\,\varrho^{7/8}.$
By the differential equation $4\xi_{\varrho}^{\prime\prime}=-\frac{1}{r}\xi_{\varrho}^{\prime}\left(1+\xi_{\varrho}^{\prime 2}\right)$ and the monotonicity of $\xi_{\varrho}^{\prime}$, we observe that $\xi_{\varrho}^{\prime\prime}$ is positive and strictly decreasing and therefore,
$|f_{\varrho}^{\prime\prime}(r)|\leq{}$ $|\eta_{\varrho}(r)\xi_{\varrho}^{\prime\prime}(r)|+|2\eta_{\varrho}^{\prime}(r)\xi_{\varrho}^{\prime}(r)|+|\eta_{\varrho}^{\prime\prime}(r)(a_{\varrho}+\xi_{\varrho}(r))$
$\leq{}$ $\xi_{\varrho}^{\prime\prime}(\sqrt{\varrho})-\frac{4}{\sqrt{\varrho}}\,\xi_{\varrho}^{\prime}(\sqrt{\varrho})+\frac{10}{\varrho}\,(a_{\varrho}+\xi_{\varrho}(\sqrt{\varrho}))$
$\leq{}$ $\varrho^{3/8}+4\sqrt{2}\,\varrho^{3/8}+10\sqrt{2}\,\varrho^{3/8}<15\sqrt{2}\,\varrho^{3/8}$
for $r\geq\sqrt{\varrho}$.
Now, let us turn to proving (iv) and (v) for $r\geq\sqrt{\varrho}$. We observe that as before, we have $|\overline{\nu}_{\varrho}^{\intercal}|\leq|f_{\varrho}^{\prime}|$ and therefore
$|\overline{\nu}_{\varrho}^{\intercal}|\leq 3\sqrt{2}\,\varrho^{7/8}=3\sqrt{2}\,\varrho^{3/8}\sqrt{\varrho}<3\sqrt{2}\sqrt{\varrho}.$
Furthermore, by Lemma A.3, the principal curvatures of the graph of $f_{\varrho}$ are given by
$|\lambda_{1}|={}$ $\frac{|f_{\varrho}^{\prime\prime}|}{(1+f_{\varrho}^{\prime 2})^{3/2}}\leq 15\sqrt{2}\,\varrho^{3/8}$
$\lambda_{2}=\dots=\lambda_{k}={}$ $-\frac{f_{\varrho}^{\prime}}{r\sqrt{1+f_{\varrho}^{\prime 2}}}{\leq}\frac{3\sqrt{2}\,\varrho^{7/8}}{\sqrt{\varrho}}=3\sqrt{2}\,\varrho^{3/8}$
for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In particular, we obtain
(5.3) $|A(\overline{\Sigma}_{\varrho})|(r)=\sqrt{\lambda_{1}^{2}+(k-1)\lambda_{2}^{2}}\leq 3\sqrt{2(k+24)}\,\varrho^{3/8}.$
The scalar curvature estimate can be derived from the Gauss equation
$\mathrm{scal}(\overline{\Sigma}_{\varrho})=H(\overline{\Sigma}_{\varrho})^{2}-|A(\overline{\Sigma}_{\varrho})|^{2}\geq-|A(\overline{\Sigma}_{\varrho})|^{2},$
which together with (5.3) yields the desired inequality and hence concludes the proof of (v). Lastly, we point out that the analysis still holds after smoothing out the $\overline{\Sigma}_{\varrho}$ at the gluing region $t=a_{\varrho}$, see for instance [BH22]. ∎
We close our analysis of the Euclidean handle with the following monotonicity result.
Proposition 5.4.
For every sufficiently small $R\in(0,\tfrac{2}{\sqrt{10}}]$ and $\varrho_{1}\in(0,R^{2}/4]$, there exists a smooth path of diffeomorphisms $\Phi_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\to\overline{\Sigma}_{\varrho}$, $\varrho\in(0,\varrho_{1}]$, such that $\Phi_{\varrho_{1},\varrho_{1}}=\operatorname{\mathrm{id}}_{\overline{\Sigma}_{\varrho_{1}}}$ and for the metrics $g_{\varrho,\delta}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$, the assignment
$\varrho\mapsto e^{\tfrac{1}{2}\varrho}\cdot\Phi_{\varrho_{1},\varrho}^{\ast}(g_{\varrho,\delta})$
is strictly monotonically increasing.
We remark that the exponent $\tfrac{1}{2}$ can be replaced by an arbitrarily small positive constant.
Before we can prove Proposition 5.4, we need some preparation. Since the Euclidean handle $\overline{\Sigma}_{\varrho}\subseteq B_{1}^{k}(0)\times[0,1]$ is rotationally symmetric, it can be obtained by rotating a curve $\overline{\gamma}_{\varrho}$. Hence, we may write the metric $g_{\overline{\Sigma}_{\varrho}}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$ in warped product form
(5.4) $g_{\varrho,\delta}=d\tau^{2}+r_{\varrho}(\tau)^{2}g_{S^{k-1}},$
where $\tau$ is the arc-length along the generating curve $\overline{\gamma}_{\varrho}$ and $r_{\varrho}(\tau)$ is the radius-function. The parametrization is such that $r_{\varrho}$ is monotonically non-increasing. For $R\in(0,\tfrac{2}{\sqrt{10}}]$ and all $\varrho<R^{2}/4$, we denote by $d_{\varrho}$ the length of the curve $\overline{\gamma}_{\varrho}$ until radius $R$. Since $\Sigma_{\varrho}$ ends in a straight cylinder, the radial function satisfies $r_{\varrho}|_{[d_{\varrho}-1+a_{\varrho},d_{\varrho}]}=\varrho^{4}$. We have the following estimate on $d_{\varrho}$ and its $\varrho$-derivative.
Lemma 5.5.
There exists $R\in(0,\tfrac{2}{\sqrt{10}}]$, depending only on $n$, such that for every $\varrho<R^{2}/4$, we have
$|\partial_{\varrho}d_{\varrho}|<\varrho^{1/4}\qquad\text{and}\qquad 1<d_{\varrho}<1+R.$
For the following proof, we use the notation $A\lesssim B$ to mean that there exists a positive constant $C$ just depending on $n$ such that $A\leq C\cdot B$.
Proof.
Recall the function $f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr)$. The length $d_{\varrho}$ of $\overline{\gamma}_{\varrho}$ is given
$d_{\varrho}=\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+\int_{4\varrho^{4}}^{R}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+(1-a_{\varrho}),$
and we thus need to compute
$\partial_{\varrho}d_{\varrho}={}$ $\underbrace{\partial_{\varrho}\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr}_{\eqqcolon A_{0}}\underbrace{-16\varrho^{3}\sqrt{1+f_{\varrho}^{\prime}(4\varrho^{4})^{2}}}_{\eqqcolon A_{1}}+\underbrace{\int_{4\varrho^{4}}^{R}\frac{f_{\varrho}^{\prime}(r)\partial_{\varrho}f_{\varrho}^{\prime}(r)}{\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}}dr}_{\eqqcolon A_{2}}-\partial_{\varrho}a_{\varrho}$
We estimate $A_{0}$, $A_{1}$ and $A_{2}$ and $\partial_{\varrho}a_{\varrho}$ individually, starting with $A_{0}$. We note that an antiderivative of $\sqrt{1+f^{\prime}_{\varrho}(r)^{2}}$ is given by
$r^{1/4}\sqrt{\sqrt{r}-\varrho^{2}}\left(\sqrt{r}+\tfrac{3}{2}\varrho^{2}\right)+\frac{3}{2}\varrho^{4}\,\operatorname{arsinh}\left(\frac{\sqrt{\sqrt{r}-\varrho^{2}}}{\varrho}\right).$
Since $\sqrt{\varrho^{4}}-\varrho^{2}=0$ and $\sqrt{\sqrt{4\varrho^{4}}-\varrho^{2}}=\varrho$, we get
$|A_{0}|=\left|\partial_{\varrho}\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr\right|\lesssim\varrho^{3}.$
Next, we turn to $A_{1}$. Since
$f_{\varrho}^{\prime}(4\varrho^{4})=\xi_{\varrho}^{\prime}(4\varrho^{4})=-\frac{\varrho}{\sqrt{\sqrt{4\varrho^{4}}-\varrho^{2}}}=-1,$
we obtain $|A_{1}|=|-16\sqrt{2}\varrho^{3}|\lesssim\varrho^{3}$.
Next, we compute
$\begin{split}\partial_{\varrho}a_{\varrho}={}&\partial_{\varrho}\left(\frac{4}{3}\varrho\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{3/2}+4\varrho^{3}\sqrt{\sqrt{2}\varrho^{1/4}-\varrho^{2}}\right)\\ ={}&\frac{4}{3}\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{3/2}+2\varrho\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{1/2}\Big(\frac{\sqrt{2}}{4}\varrho^{-3/4}-2\varrho\Big)\\ &+12\varrho^{2}\sqrt{\sqrt{2}\varrho^{1/4}-\varrho^{2}}+2\varrho^{3}\Big(\sqrt{2}\varrho^{1/4}-\varrho^{2}\Big)^{-1/2}\Big(\frac{\sqrt{2}}{4}\varrho^{-3/4}-2\varrho\Big)\\ \leq{}&11\varrho^{3/8}+4\varrho^{1+1/8-3/4}+24\varrho^{2+1/8}+2\varrho^{3-1/8-3/4}.\end{split}$
Now we turn to $A_{2}$. We have
$|A_{2}|\leq\int_{4\varrho^{4}}^{R}|f_{\varrho}^{\prime}(r)||\partial_{\varrho}f_{\varrho}^{\prime}(r)|dr.$
Note that $f_{\varrho}=\xi_{\varrho}$ for $r\leq\sqrt{\varrho}$. We obtain
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $1\cdot 2\varrho^{-1}$ $\text{for $r\in[4\varrho^{4},\varrho^{2}]$},$
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $2\sqrt{\varrho}\cdot 2\varrho^{-1/2}$ $\text{for $r\in[\varrho^{2},\varrho]$},$
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|\leq$ $2\varrho^{3/4}\cdot 2\varrho^{-1/4}$ $\text{for $r\in[\varrho,\sqrt{\varrho}]$}.$
Moreover, we have for $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$:
$|f_{\varrho}^{\prime}(r)|\cdot|\partial_{\varrho}f_{\varrho}^{\prime}(r)|$
$\quad\leq\Big(|\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\eta_{\varrho}||\xi_{\varrho}^{\prime}|\Big)$
$\qquad\qquad\cdot\Big(|\partial_{\varrho}\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\partial_{\varrho}\eta_{\varrho}||\xi_{\varrho}^{\prime}|+|\eta_{\varrho}^{\prime}||\partial_{\varrho}\xi_{\varrho}+\partial_{\varrho}a_{\varrho}|+|\eta_{\varrho}||\partial_{\varrho}\xi_{\varrho}^{\prime}|\Big)$
$\quad\lesssim\Big(\varrho^{-1/2}\cdot\varrho^{5/4}+1\cdot\varrho^{3/4}\Big)$
$\qquad\qquad\cdot\Big(\varrho^{-2}\cdot\varrho^{5/4}+\varrho^{-3/2}\cdot\varrho^{3/4}+\varrho^{-1/2}\cdot(1+1)+1\cdot\varrho^{-1/8}\Big)$
$\quad\lesssim\varrho^{3/4}\cdot\varrho^{-3/4},$
where we used $\partial_{\varrho}a_{\varrho}\leq 1$ and that on $\sqrt{\varrho},2\sqrt{\varrho}]$ by the fundamental theorem of calculus
$|f_{\varrho}|\leq\sqrt{\varrho}|\xi^{\prime}(\sqrt{\varrho})|\leq 2\sqrt{\varrho}\cdot\varrho^{3/4}=2\varrho^{5/4}.$
Finally, for $r\geq 2\varrho$, $f_{\varrho}\equiv 0$ and therefore, putting the pieces together and integrating, we obtain
$|A_{2}|\lesssim\sqrt{\varrho}.$
Hence,
$|\partial_{\varrho}d_{\varrho}|\lesssim\sqrt{\varrho}.$
Since the curve $\overline{\gamma}_{\varrho}$ is shorter than the one consisting of straight ones parallel to the coordinate-axes and longer than the straight line connecting its end points, we get
$1\leq\sqrt{1+R^{2}}<d_{\varrho}<1+R\qed$
Proof of Proposition 5.4.
Fix the bump function $\beta(x)=\eta(4x)\eta(-4x)$, where $\eta$ is the cut-off function introduced earlier satisfying $2\leq\eta^{\prime}\leq 0$ and $|\eta^{\prime\prime}|<10$, and we note that $\operatorname{supp}\beta\subseteq[-\frac{1}{4},\frac{1}{4}]$. Next, we define a smooth family or reparameterizations $\varphi_{\varrho_{1},\varrho}:[0,d_{\varrho_{1}}]\to[0,d_{\varrho}]$ for $\varrho\in(0,\varrho_{1}]$ by
(5.5) $\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\beta\left(s-\frac{1}{2}\right)ds$
where $q_{\varrho_{1},\varrho}=(d_{\varrho_{1}}-d_{\varrho})\left(\int_{\mathbb{R}}\beta(s)ds\right)^{-1}$. We now define a diffeomorphism
$\overline{\Phi}_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\longrightarrow\overline{\Sigma}_{\varrho},\quad(\tau,x)\mapsto(\varphi_{\varrho_{1},\varrho}(\tau),x)$
where $x$ denotes the spherical coordinates in the warped product. Note that
$\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)=(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))^{2}\mathrm{d}\tau^{2}+r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))^{2}g_{S^{k-1}}.$
A direct inspection of the warping functions $r_{\varrho}$ gives for any $\varrho\geq\widetilde{\varrho}$
$r_{\varrho}\left(\varphi_{\varrho_{1},\varrho}(\tau)\right)\leq r_{\widetilde{\varrho}}\left(\varphi_{\varrho_{1},\widetilde{\varrho}}(\tau)\right),\qquad\tau\in[0,d_{\varrho_{1}}].$
This implies
$\partial_{\varrho}r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))=$ $(\partial_{\varrho}r_{\varrho})(\varphi_{\varrho_{1},\varrho}(\tau))+r_{\varrho}^{\prime}(\varphi_{\varrho_{1},\varrho}(\tau))\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)\geq 0$
where we used that $r_{\varrho}^{\prime}$ vanishes on the support of $\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)$. Next, we observe that
$\partial_{\tau}\varphi(\tau)=1-(d_{\varrho_{1}}-d_{\varrho})\beta\left(\tau-\frac{1}{2}\right)$
and therefore
$\partial_{\varrho}\partial_{\tau}\varphi(\tau)=\beta\left(\tau-\frac{1}{2}\right)\partial_{\varrho}d_{\varrho}-(d_{\varrho_{1}}-d_{\varrho})\beta^{\prime}(\tau-\frac{1}{2})>-\frac{1}{2}\partial_{\tau}\varphi(\tau)$
where we used that $(d_{\varrho_{1}}-d_{\varrho})\ll 1$ and that $\beta^{\prime}$ is bounded. Consequently, we obtain for $\varrho\in[\varrho_{0},\varrho_{1}]$
$\partial_{\varrho}\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)\geq-\frac{1}{2}\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)$
which finishes the proof. ∎
To conclude the construction of $\overline{\Sigma}$, we will now smoothen the $C^{1}$-singularity of $\overline{\Sigma}_{\varrho}$ at $\{t=a_{\varrho}\}$ while preserving Proposition 5.3 and Lemma 5.5, hence also Proposition 5.4. Fix some $R\in(0,\tfrac{2}{\sqrt{10}}]$ so that Lemma 5.5 holds for this $R$. Consider the smooth manifold
$V=\{(\varrho,t)\mid\varrho\in(0,R^{2}/4),t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset(0,R^{2}/4)\times(0,1]$
with non-compact boundary $\partial V=\{(\varrho,a_{\varrho})\mid\varrho\in(0,R^{2}/4)\}\subset V$.
Put $f_{0}\colon V\to\mathbb{R}$, $f_{0}(\varrho,t)=\xi_{\rho}^{-1}(t-a_{\varrho})$. Note that for fixed $\varrho$, the function $f_{0}(\varrho,-)\colon(f_{\varrho}(\varrho^{2}),a_{\varrho}]\to\mathbb{R}$ is the inverse of the function $\varrho^{4},\varrho^{2})\to\mathbb{R}$, $r\mapsto f_{\varrho}(r)$.
Let $\omega\colon V\to\mathbb{R}_{>0}$ be a smooth function. For each $\varrho\in(0,R^{2}/4)$, we obtain the surface of revolution
$S(\omega,\varrho):=\{r=\omega(t)\mid t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset\mathbb{R}^{n}\times(f_{\varrho}(\varrho^{2}),a_{\varrho}].$
Let $J^{2}V\to V$ be the bundle of $2$-jets of strictly positive real valued smooth functions on $V$. We consider the open partial differential relation $\mathfrak{R}\subset J^{2}V$ determined by the following pointwise conditions for $\omega$, working with the numbers $|\partial_{\varrho}d_{\varrho}|<\varrho^{1/4}$, from Lemma 5.5:
  1. (i)
    $H(S(\omega,\varrho))>\frac{1}{2}|A(S(\omega,\varrho))|$ if $k\geq 2$,
  2. (ii)
    $\mathrm{scal}(S(\omega,\varrho))>\frac{1}{3}|A(S(\omega,\varrho))|^{2}$ if $k\geq 3$,
  3. (iii)
    $\min\left\{H(S(\omega,\varrho))^{2},\mathrm{scal}(S(\omega,\varrho))\right\}>\tfrac{1}{2}\varrho^{-3}$
  4. (iv)
    $\left|\partial_{\varrho}\left(\sqrt{1+\omega^{\prime}(\varrho,-)}-\sqrt{1+f^{\prime}_{0}(\varrho,-)}\right)\right|<\varrho^{1/4}-|\partial_{\varrho}d_{\varrho}|$.
By Proposition 5.3, $f_{0}$ solves $\mathfrak{R}$ over $V$. Consider the smooth path $F\colon[0,1]\to C^{\infty}(V,\mathbb{R}_{>0})$,
$F(\tau)(\varrho,t):=(1-\tau)f_{0}(\varrho,t)+\tau\varrho^{4}.$
We have $j^{1}F(\tau)|_{\partial V}=j^{1}f_{0}|_{\partial V}$ for all $t$. Carrying out the computation in the proof of part (ii) and (iii) of Proposition 5.3 with $\lambda_{2}=\ldots=\lambda_{k}=\varrho^{4}$ and $\lambda_{1}\in[-\tfrac{1}{4}\varrho^{4},0]$, one sees that $F$ solves $\mathfrak{R}$ along $\partial V$ for all $\tau$ (for $\omega=F(\tau)$, the left-hand side of (iv) is equal to 0 along $\partial V$). By continuity, there exists an open neighborhood $\partial V\subset U\subset V$ such that $F$ solves $\mathfrak{R}$ over $U$ for all $t$.
By the local flexibility lemma [BH22, Theorem 1.2] applied to the closed subset $\partial V\subset V$, there exists an open neighborhood $\partial V\subset U_{0}\subset U$ satisfying $\overline{U}_{0}\subset V$ and a smooth path $f\colon[0,1]\to C^{\infty}(V,\mathbb{R}_{>0})$ such that $f(0)=f_{0}$, $f(\tau)=F(\tau)$ on $U_{0}$, $f(\tau)=f_{0}$ on $V\setminus U$ and $f(\tau)$ solves $\mathfrak{R}$ over $V$ for all $\tau$.
For $\varrho\in(0,R^{2}/4)$, define
  • the smooth function $\psi_{\varrho}\colon(f_{\varrho}(\varrho^{2}),a_{\varrho}]\to\mathbb{R}_{>0}$, $\psi_{\varrho}(t)=f(1)(\varrho,t)$. It is equal to $\varrho^{4}$ near $a_{\varrho}$ and equal to $f_{0}(\varrho,-)$ near $f_{\varrho}(\varrho^{2})$;
  • the surface of revolution
    $S(\varrho):=\{r=\psi_{\varrho}(t)\mid t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset B_{1}^{k}(0)\times(f_{\varrho}(\varrho^{2}),a_{\varrho}];$
  • the smooth hypersurface
    $\widetilde{\Sigma}_{\varrho}:=\{(r,t)\in\overline{\Sigma}_{\varrho}\mid t\notin(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\;\bigcup\;S(\varrho)\subset B_{1}^{k}(0)\times[0,1].$
The hypersurface $\widetilde{\Sigma}_{\rho}$ satisfies the properties in Proposition 5.3 by the choice of $\mathfrak{R}$. Note that parts (iv) and (v) of Proposition 5.3 are satisfied since the relevant parts of $\overline{\Sigma}_{\varrho}$ are not affected by the smoothing.
Let $\widetilde{d}_{\rho}$ be defined for $\widetilde{\Sigma}_{\varrho}$ as before Lemma 5.5. Using that $\overline{\Sigma}_{\varrho}=\widetilde{\Sigma}_{\varrho}$ on $\{t\notin(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}$, and that $\psi^{\prime}_{\varrho}(t)=f_{0}^{\prime}(\varrho,-)(t)$ for $t=f_{\varrho}(\varrho^{2})$ and for $t=a_{\varrho}$ (where both sides of the equation are 0), we obtain
$|\partial_{\varrho}\widetilde{d}_{\varrho}|-|\partial_{\varrho}d_{\varrho}|$ $\leq\Big|\partial_{\varrho}(\widetilde{d}_{\varrho}-d_{\varrho})\Big|$
$=\Big|\partial_{\varrho}\int_{f_{\varrho}(\varrho^{2})}^{a_{\varrho}}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)dt\Big|$
$=\Big|\int_{f_{\varrho}(\varrho^{2})}^{a_{\varrho}}\partial_{\varrho}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)dt\Big|$
$\leq\big|a_{\varrho}-f_{\varrho}(\varrho^{2})\big|\max_{(f_{\varrho}(\varrho^{2}),a_{\varrho}]}\Big|\partial_{\varrho}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)\Big|$
$<1\cdot\left(\varrho^{1/4}-|\partial_{\varrho}d_{\varrho}|\right).$
Adding $|\partial_{\varrho}d_{\varrho}|$ to both sides of this inequality proves the first inequality from Lemma 5.5. The inequality $1<\widetilde{d}_{\varrho}$ is clear, while the inequality $\widetilde{d}_{\varrho}<1+R$ is ensured if $\|f(1)-f_{0}\|_{C^{1}(V)}$ is sufficiently small, a condition that can be included in the relation $\mathfrak{R}$.
Finally, we replace $\overline{\Sigma}_{\varrho}$ with the smooth bending handle $\widetilde{\Sigma}_{\varrho}$. By construction, Proposition 5.3 and Lemma 5.5, hold for the new $\overline{\Sigma}_{\varrho}$, which we will use from now on.
6. Surgery bending in arbitrary backgrounds
In this section, we will construct the general surgery bending. To this end, let $\mathcal{S}$ be a compact submanifold of codimension $k$ in an $n$-dimensional Riemannian manifold $(M,g_{M})$. We denote its normal bundle by $\mathcal{N}_{\mathcal{S}}$. Recall the constant $R\in(0,\tfrac{2}{\sqrt{10}}]$ from the previous section, which we may choose smaller if necessary, so that the normal exponential map of $\mathcal{N}_{\mathcal{S}}$ along $\mathcal{S}$ is a diffeomorphism on the sub-bundle of vectors with length less than or equal to $R$ to the tubular neighborhood $B_{R}(\mathcal{S})$ of radius $R$ around $\mathcal{S}$. With this identification, we consider $B_{R}(\mathcal{S})$ as a fibre bundle over $\mathcal{S}$ with fibre $B_{r}^{k}(0)$.
We will construct a smooth submanifold $\Sigma_{\varrho}\subset M\times[0,1]$ describing a surgery bending in $\mathcal{N}_{\mathcal{S}}$ which for $M=\mathbb{R}^{k}$ with the Euclidean metric and $\mathcal{S}$ a point recovers 5.1. First, we choose Fermi coordinates around $\mathcal{S}$, that is, we have an identification $U\times B^{k}_{R}(0)\cong B_{R}(\mathcal{S})|_{U}$, where $U\subseteq\mathcal{S}$ is a trivialising open subset for $\mathcal{N}_{\mathcal{S}}$. We define for $\varrho\in(0,R^{2}/4)$
$\Sigma_{\varrho}|_{U}\coloneqq U\times\overline{\Sigma}_{\varrho}$
for the smooth submanifold $\overline{\Sigma}_{\varrho}\subset B^{k}_{R}(0)\times[0,1]$ as in 5.1. Since $\mathcal{N}_{\mathcal{S}}$ has structure group $\mathrm{O}(k)$, using the $\mathrm{O}(k)$-invariance of $\overline{\Sigma}_{\varrho}$ (cf. Proposition 5.3, (i)), the local definitions can be glued together into a global definition of $\Sigma_{\varrho}\subset B_{R}(\mathcal{S})\times[0,1]$. Moreover, since $2\sqrt{\varrho}<R$, the hypersurface $\Sigma_{\varrho}$ can be extended by $(M\setminus B_{R}(\mathcal{S}))\times\{0\}$ to a smooth hypersurface $\Sigma_{\varrho}\subset M\times[0,1]$.
The main result of this section is the following theorem, which states the existence of a background metric on $M\times[0,1]$ that extends $g_{M}$ on $M\times\{0\}$ such that for $\varrho$ small enough, the scalar curvature of $\Sigma_{\varrho}$ can be uniformly bounded in terms of the scalar curvature of the original metric $g_{M}$ and, for $\varrho$ even smaller the mean curvature of $\Sigma_{\varrho}$ is positive.
Theorem 6.1.
Let $k\geq 3$ and $\alpha\in(0,\tfrac{1}{3}]$. Then, there exists a constant $0<\varrho_{1}<\min(R^{2}/4,\tfrac{\alpha}{3})$ such that for all $\kappa,\varrho\in(0,\varrho_{1}]$ and for the metric
$g_{\kappa}=u^{\frac{4}{\,n-2\,}}(g_{M}+dt^{2}),\qquad u(x,t)=u_{\kappa}(x,t)=1+t\tilde{\kappa}\eta\left(\alpha^{-1}t-1\right)$
on $M\times[0,1]$, where $\tilde{\kappa}\coloneqq\tfrac{n-2}{2n}\kappa$, the scalar curvature of the induced metric $g_{\kappa,\varrho}$ on $\Sigma_{\varrho}$ satisfies
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4},$
where $\mathrm{pr}_{M}\colon M\times[0,1]\to M$ is the projection. Furthermore, there exists a constant $\varrho_{0}\in(0,\varrho_{1})$ such that the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\kappa}$ satisfies
$H(\Sigma_{\varrho},g_{\kappa})>(1-\alpha)\kappa-\varrho^{1/4}$
for all $0<\varrho\leq\varrho_{0}$.
Remark 6.2.
  1. (i)
    We emphasize that $\varrho_{1}$ is independent of $\kappa$, while $\varrho_{0}$ will depend on $\kappa$.
  2. (ii)
    The mean curvature of the slice $M\times\{0\}\subset M\times[0,1]$ with respect to the ambient metric $g_{\kappa}$ and the exterior normal is given by
    $H(M\times\{0\},g_{\kappa})=-\frac{2n}{n-2}\tilde{\kappa}=-\kappa.$
  3. (iii)
    If $k=2$ and $n\geq 3$, then the second statement of Theorem 6.1 still holds true, that is, the mean curvature of $\Sigma_{\varrho}$ with respect to $g_{\kappa}$ is still positive. However, the scalar curvature estimate might be incorrect.
Remark 6.3.
Essentially, the content of Theorem 6.1 is that the handle construction of Proposition 5.3 extends to arbitrary Riemannian manifolds and to surgeries of any codimension $k\geq 3$. This leads to a number of error terms. They can be grouped into three different categories:
  1. (i)
    Error terms which approach zero as $\varrho\to 0$. These are not problematic.
  2. (ii)
    Error terms of the form $\varepsilon|\overline{A}_{\varrho}|$, $\varepsilon\ll 1$, where we note that the second fundamental form $|A_{\varrho}|$ blows up as $\varrho\to 0$. Using the pinching estimates $H(\overline{\Sigma}_{\varrho})>\tfrac{1}{2}|A(\overline{\Sigma}_{\varrho})|$ and $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\tfrac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}$ we can ensure that they can be absorbed.
  3. (iii)
    Error terms which involve $|\overline{\nu}^{\intercal}_{\varrho}|$. They can be controlled by the dichotomy $\min\{H(\overline{\Sigma}_{\varrho})^{2},\mathrm{scal}(\overline{\Sigma}_{\varrho})\}>\tfrac{1}{2}\varrho^{-3}$ for $r\in[\varrho^{4},\varrho^{2}]$ and $|\overline{\nu}_{\varrho}^{\intercal}|<\overline{C}\,\sqrt{\varrho}$ for $r\geq\varrho^{2}$.
Before diving into the proof of Theorem 6.1, let us introduce its main driving force, which are estimates for scalar and mean curvature, as well as the appropriate conformal transformation formulae.
For the following lemma, we use the notation $A\lesssim B$ to mean that there exists a positive constant $C$ independent of $\varrho$ such that $A\leq C\cdot B$.
Lemma 6.4.
There exists a constant $\varepsilon>0$ such that for every $0<\varrho\leq\varepsilon$ and for the metric $g_{\varrho}$ on $\Sigma_{\varrho}$ induced by $g+\mathrm{d}t^{2}$, the following inequality holds:
(6.1) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\\ &\qquad\qquad\gtrsim-\varrho\cdot\bigl(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\bigr)^{2}-\|\overline{\nu}_{\varrho}^{\intercal}\|\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\end{split}$
Here, $\|\overline{\nu}_{\varrho}^{\intercal}\|$ denotes the norm of the tangential part of the normal vector of the hypersurface $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ constructed in 5.1 with respect to the Euclidean metric (see also (5.1)), and $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$ denotes the scalar curvature of this hypersurface with respect to the metric $g_{\varrho,\delta}$ induced by the Euclidean ambient metric $\delta+\mathrm{d}t^{2}$, composed with the projection onto the fiber of the tubular neighbourhood $D(\mathcal{S})$. Furthermore, we have the following inequality concerning the mean curvature
(6.2) $|H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|),$
and we can compare the $\partial_{t}$-part of the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ with respect to $g+\mathrm{d}t^{2}$ and the normal vector $\overline{\nu}_{\varrho}$ of $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ taken with respect to the Euclidean metric $\delta+\mathrm{d}t^{2}$ as follows:
(6.3) $|\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}|\lesssim\varrho.$
Lemma 6.5.
Let $u$, $g_{\kappa}$ and $g_{\varrho,\kappa}$ be as defined in Theorem 6.1. Then we have the following equalities comparing the scalar and mean curvature before and after conformal transformation by $u^{4/(n-2)}$:
[…]
(6.5) $\begin{split}H(\Sigma_{\varrho},g_{\kappa})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)\end{split}$
where the normal vector $\nu_{\varrho}$ of $\Sigma_{\varrho}$ and the inner product with $\partial_{t}$ are taken with respect to the metric $g+\mathrm{d}t^{2}$.
Lemma 6.5 is well-known, and we will postpone the proof of Lemma 6.4 for now and instead turn to the proof of Theorem 6.1.
Proof of Theorem 6.1.
In order to estimate the scalar curvature and mean curvature of $g_{\varrho,\kappa}$, we first note that at points with radial coordinate $r\geq 2\sqrt{\varrho}$ we have $\Sigma_{\varrho}=M$ and $g_{\varrho}=g$. So, the scalar curvature estimate is trivially true here, whereas the estimate on the mean curvature follows from Lemma 6.5 using $\partial_{t}u=\tilde{\kappa}$, $u=1=\langle\nu_{\varrho},\partial_{t}\rangle$ and $H(M\times\{0\},g+\mathrm{d}t^{2})=0$. Hence, it suffices to consider $r\leq 2\sqrt{\varrho}$. We will distinguish $3$ cases: $r\in[\varrho^{4},\varrho^{2}]$, $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. In every case we will first estimate the scalar and mean curvature with respect to the ambient metric $g+\mathrm{d}t^{2}$ using Lemma 6.4 before investigating the effect of conformal change on scalar curvature employing Lemma 6.5. At the very end of the proof, we will prove positivity of the mean curvature after conformal change after passing to a possibly smaller $\varrho_{0}$, depending on $\kappa$.
Let us start by considering $r\in[\varrho^{4},\varrho^{2}]$, where Proposition 5.3 yields
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{2}\varrho^{-3}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2},$
where $g_{\varrho,\delta}$ is the metric on $\overline{\Sigma}_{\varrho}\subset\mathbb{R}^{k+1}$ induced by the ambient Euclidean metric $\delta$. This also yields $\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{6}\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|$. Hence,
$\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})>\frac{1}{20}\left(\varrho^{-3}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right).$
Furthermore, we obtain from (6.1) from Lemma 6.4 that there exist constants $C_{0}>0$, independent of $\kappa$, such that:
(6.6) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}\\ &\quad>{}\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-C_{0}\Bigl(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}+\|\overline{\nu}_{\varrho}^{\intercal}\|(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\Bigr),\end{split}$
where $\mathrm{pr}_{M}\colon\Sigma_{\varrho}\subset M\times[0,1]\to M$ denotes the projection. We will omit $\circ\mathrm{pr}_{M}$ for the remainder of this proof to ease the notation. Since $\|\overline{\nu}_{\varrho}^{\intercal}\|\leq 1$, and we thus obtain
$\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\\ &\qquad>\frac{1}{20}\left(\varrho^{-3}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}+\varrho^{-3/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right)\\ &\qquad\qquad-C_{0}\Bigl(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}+(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\Bigr)\\ &\qquad\geq\varrho^{-3}\left(\frac{1}{20}-C_{0}\varrho^{3}\right)+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{20}\varrho^{-3/2}-C_{0}\right)\\ &\qquad\qquad+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\left(\frac{1}{20}-C_{0}\varrho\right)\end{split}$
Hence, for $c_{0}\coloneqq\tfrac{1}{40}$ and $\varrho_{1}\leq\tfrac{1}{40C_{0}}$, we have
(6.7) $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})+c_{0}\cdot\varrho^{-3}.$
For the mean curvature, we first recall that by (6.2) from Lemma 6.4 there exists a constant $C_{1}$ such that
(6.8) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})-H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq-C_{1}\left(\|\overline{\nu}_{\varrho}^{\intercal}\|+\varrho(1+|A(\overline{\Sigma},\delta+\mathrm{d}t^{2})|)\right).$
Furthermore, we have
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\quad\text{and}\quad H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}\varrho^{-3/2},$
by Proposition 5.3. Therefore, $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\tfrac{1}{4}(|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|+\varrho^{-3/2})$ and (6.8) becomes:
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>{}$ $\frac{1}{4}\left(\varrho^{-3/2}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\right)$
$-C_{1}\left(1+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$={}$ $\varrho^{-3/2}\left(\frac{1}{4}-C_{1}(\varrho^{3/2}+\varrho^{5/2})\right)+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{4}-C_{1}\varrho\right)$
Hence, for $c_{1}\coloneqq\tfrac{1}{8}$ and $\varrho_{1}\leq\tfrac{1}{16C_{1}}$, we get
(6.9) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>c_{1}\cdot\varrho^{-3/2}$
To close the first case of the proof for the scalar curvature estimate, let us consider the conformal transformation formula (6.4) from Lemma 6.5:
$\begin{split}\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})={}&u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\frac{4(n-1)}{n-2}(1-\langle\nu_{\varrho},\partial_{t}\rangle^{2})\partial^{2}_{t}u\\ &\qquad\qquad+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\partial_{t}u\Bigr)\end{split}$
If $t\geq 2\varrho$, then $\Sigma_{\varrho}$ is given by the straight cylinder of radius $\varrho^{4}$, so $\langle\nu_{\varrho},\partial_{t}\rangle=0$. Furthermore, $u\leq 2$ and
$\partial^{2}_{t}u=\alpha^{-1}\tilde{\kappa}\eta^{\prime}(\alpha^{-1}t-1)+\alpha^{-2}t\tilde{\kappa}\eta^{\prime\prime}(\alpha^{-1}t-1)\lesssim\alpha^{-2}\varrho_{1}\lesssim\varrho_{1}^{-1},$
by the choice of $\eta$ and $\kappa$. Therefore, we may assume $\partial^{2}_{t}u\leq C_{1}\varrho_{1}$ and the above formula gives the following estimate
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $2^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g_{M})+\varrho^{-3}(c_{1}-\frac{4(n-1)}{n-2}C_{0}\varrho_{1}^{2})\Bigr).$
We choose $\varrho_{1}$ such that $c_{1}-\tfrac{4(n-1)}{n-2}C_{0}\varrho_{1}^{2}\leq c_{1}/2$ and $c_{1}\varrho_{1}^{-3/2}\geq 66\max|\mathrm{scal}(M,g_{M})|$, so that we get $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>|\mathrm{scal}(M,g_{M})|$ as desired.
On the other hand, if $t\leq 2\varrho\leq\alpha$, then $u=1+t\tilde{\kappa}$, $\partial_{t}u=\tilde{\kappa}$ and $\partial_{t}^{2}u=0$. As $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ is positive, so is the entire final summand, and we get
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $2^{-\frac{n+2}{n-2}}\underbrace{\Bigl(\mathrm{scal}(M,g_{M})+c_{1}\varrho^{-3}\Bigr)}_{\geq 32\max|\mathrm{scal}(M,g_{M})|}\geq|\mathrm{scal}(M,g_{M})|,$
which finishes the first case.
Next, we investigate the region $r\in[\varrho^{2},\sqrt{\varrho}]$, where Proposition 5.3 gives the following estimates:
$\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\quad\text{and}\quad\|\nu_{\varrho}^{\intercal}\|<\overline{C}\sqrt{\varrho}$
for $\overline{C}=18k+432$ the constant from Proposition 5.3. We set $C_{2}\coloneqq 4C_{0}\cdot\overline{C}$, and observe that (6.6) yields
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})$
$\qquad>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
$\qquad\qquad-\frac{C_{2}}{4}\left(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}-\varrho^{1/2}\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$\qquad\geq-\frac{C_{2}}{4}\varrho^{1/2}-\frac{C_{2}}{2}\varrho^{1/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|+\left(\frac{1}{3}-\frac{C_{2}}{4}\varrho\right)|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
If $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq 1$, then this becomes $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})>-C_{2}\varrho^{1/2}$. On the other hand, if $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\geq 1$, then $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$ and we obtain
$\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})>-\frac{C_{2}}{4}\varrho^{1/2}+\left(\frac{1}{3}-\frac{C_{2}}{2}\varrho^{1/2}-\frac{C_{2}}{4}\varrho\right)|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\end{split}$
and thus, for $\varrho_{1}\leq(\frac{1}{3C_{2}})^{2}$, we have
(6.10) $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})-C_{2}\varrho^{1/2}$
Concerning the mean curvature, Proposition 5.3 again yields
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|,$
and we can hence estimate
$\begin{split}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>{}&\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-C_{1}\cdot\overline{C}\left(\varrho^{1/2}+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)\\ \geq{}&-C_{3}\varrho^{1/2}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{2}-C_{3}\varrho\right)\end{split}$
where $C_{3}\coloneqq 2C_{1}\cdot\overline{C}$. If $\varrho_{1}\leq\tfrac{1}{4C_{3}}$, we get
(6.11) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-C_{3}\varrho^{1/2}$
By employing the conformal change formula from Lemma 6.5, we can use $u=1+t\tilde{\kappa}$ to obtain
[…]
$>{}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g_{M})-C_{3}\varrho^{1/2}\Bigr)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right),$
if $\varrho_{1}^{1/4}C_{3}\leq\tfrac{1}{2}$. Since $u=1+t\tilde{\kappa}\leq 1+\kappa\leq 1+\varrho_{1}$, we have $1\geq u^{-1}\geq u^{-(n+2)/(n-2)}\geq u^{{-5}}\geq 1-5t\tilde{\kappa}\geq 1-5\varrho_{1}$. [^2] This follows from the inequality $(1+x)^{-n}\geq 1-nx$ for $n\in\mathbb{N}$. If $\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\leq 0$, then
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\geq\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho_{1}^{1/4},$
so it suffices to consider the case that $\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\geq 0$, where we can argue similarly:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $(1-5\varrho_{1})\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right)$
$>{}$ $\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}\Bigl(5\cdot\mathrm{scal}(M,g_{M})\varrho_{1}^{3/4}+\frac{1}{2}\Bigr),$
and, if $\varrho_{1}$ is chosen such that $5\max\mathrm{scal}(M,g_{M})\varrho_{1}^{3/4}<\frac{1}{2}$, we obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}.$
Finally, we consider $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$. By Proposition 5.3, we have
$\|\nu_{\varrho}^{\intercal}\|\lesssim\varrho^{1/2},\qquad|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\varrho^{3/8}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho})\gtrsim-\varrho^{3/4}.$
By consulting (6.1) from Lemma 6.4 one final time, we deduce that
(6.12) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\\ &\qquad\gtrsim-\varrho^{3/4}-\varrho\cdot(1+\varrho^{3/8})^{2}-\varrho^{1/2}\cdot(1+\varrho^{3/8})\gtrsim-\varrho^{1/2}\end{split}$
Furthermore, using $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\gtrsim\varrho^{3/8}$ we obtain
(6.13) $\begin{split}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\gtrsim{}&H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-\|\nu_{\varrho}^{\intercal}\|-\varrho(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\\ \gtrsim{}&-\varrho^{3/8}-\varrho^{1/2}-\varrho-\varrho^{11/8}\gtrsim-\varrho^{3/8}\end{split}$
Thus there exists a constant $C_{4}>0$ such that
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})-C_{4}\varrho^{1/2}\quad\text{and}\quad H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-C_{4}\varrho^{3/8}.$
An application of Lemma 6.5 together with the observation that $\partial_{t}u=\tilde{\kappa}\leq 1$, $\partial_{t}^{2}u=0$ and $\tfrac{4(n-1)}{(n-2)}\leq 8$ now gives:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\varrho^{1/4}C_{4}\cdot(\varrho^{1/4}+8\varrho^{1/8})\right)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right),$
provided $\varrho_{1}$ satisfies $C_{4}\cdot(\varrho^{1/4}+8\varrho^{1/8})<\tfrac{1}{2}$. Using $\varrho^{3/8}\leq\varrho^{1/4}$, the same argument as in the second case, below (6.11) shows that
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}.$
It remains to show positivity of the mean curvature with respect to the conformally transformed metric after passing to a smaller radius bound $\varrho_{0}$ which may depend on $\kappa$. Recall the conformal transformation formula:
$H(\Sigma_{\varrho},g_{\kappa})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)$
We will again distinguish radial cases.
If $r\in[\varrho^{4},\varrho^{2}]$, we have already seen that
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>c_{1}\varrho^{-3/2},$
see (6.9) Since $\partial_{t}u=\tilde{\kappa}\eta(\alpha^{-1}t-1)+\alpha^{-1}t\tilde{\kappa}\eta^{\prime}(\alpha^{-1}t-1)\geq-3\alpha^{-1}$, $u^{-1}\leq 1$, $\tilde{\kappa}\leq 1$ and $\langle\nu_{\varrho},\partial_{t}\rangle\leq 1$, we thus obtain for $C_{5}\coloneqq\tfrac{6n\alpha^{-1}}{(n-2)}$:
(6.14) $H(\Sigma_{\varrho},g_{\kappa})>u^{-\frac{2}{n-2}}\left(c_{1}\varrho^{-3/2}-C_{5}\right).$
If $\varrho_{0}$ is chosen such that $c_{1}\varrho_{0}^{-3/2}-C_{5}\geq 2^{2/(n-2)}\geq u^{2/(n-2)}\kappa$, then
$H(\Sigma_{\varrho},g_{\kappa})>\kappa>(1-\alpha)\kappa-\varrho^{1/4}$
as desired.
The cases $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$ can be treated simultaneously. We first observe, that (6.11) and (6.13) both give
(6.15) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-\varrho^{1/4},$
provided $\varrho_{0}$ is small enough. Furthermore, combining $\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}\gtrsim-\varrho$ from Lemma 6.4 and $\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}-1\gtrsim-\varrho^{1/2}$ from Proposition 5.3, we find a constant $C_{6}>0$ such that we have
(6.16) $\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\geq 1-C_{6}\varrho^{1/2}\geq 1-\frac{1}{2}\varrho^{1/4}$
if $\varrho_{0}\leq C_{6}^{-4}$. From (6.15) and (6.16), together with $\partial_{t}u=\tilde{\kappa}=\tfrac{n-2}{2n}\kappa$ and $u^{-1}\geq 1-\varrho_{1}$, we obtain
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u$ $>-\frac{1}{2}\varrho^{1/4}+(1-\varrho^{1/4})(1-\varrho_{1})\kappa$
$=(1-\varrho_{1})\kappa-\frac{1}{2}\varrho^{1/4}-\varrho^{1/4}(1-\varrho_{1})\kappa$
$\geq(1-\varrho_{1})\kappa-\varrho^{1/4}$
Note, that $u^{-\frac{2}{n-2}}\geq 1-2\varrho_{1}$, and we choose $\varrho_{0}\leq\tfrac{1}{2}(1-\varrho_{1})^{4}\kappa^{4}$ small enough, so that $(1-\varrho_{1})\kappa-\varrho^{1/4}>0$ as well. The transformation formula (6.5) then gives
(6.17) $\begin{split}H(\Sigma_{\varrho},g_{\kappa})>{}&(1-2\varrho_{1})\Bigl((1-\varrho_{1})\kappa-\varrho^{1/4}\Bigr)\\ ={}&(1-3\varrho_{1})\kappa-\varrho^{1/4}\geq(1-\alpha)\kappa-\varrho^{1/4},\end{split}$
which finishes the proof. ∎
Let us now dive into the proof of Lemma 6.4.
Proof of Lemma 6.4.
By the Gauss equation we have
(6.18) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\underbrace{\mathrm{scal}(M\times[0,1],g+\mathrm{d}t^{2})}_{=\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}}\\ &\qquad={}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})^{2}-|A(\Sigma_{\varrho},g+\mathrm{d}t^{2})|^{2}-2\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho}),\end{split}$
where $\nu_{\varrho}$ is the upwards unit normal vector of $\Sigma_{\varrho}$ with respect to the ambient metric $g+\mathrm{d}t^{2}$ and $\mathrm{pr}_{M}\colon M\times[0,1]\to M$ is the projection. Again, we will omit $\circ\mathrm{pr}_{M}$ from now on. In order to estimate the summands from the right-hand side of (6.18), consider the symmetric $2$-tensor $\gamma$ on $\mathcal{S}$ defined as the first order Taylor expansion of $g$ in normal direction, that is
(6.19) $\gamma_{\alpha\beta}=g_{\alpha\beta}|_{\mathcal{S}}+\sum_{i=1}^{k}x^{i}g_{\alpha\beta,i}|_{\mathcal{S}}$
which we extend canonically to $B_{R}^{k}$. Since the tensor $\gamma$ precisely describes the $1$-jet of the metric $g$ along $\mathcal{S}$, there exists an $\varepsilon>0$ such that the restriction of $\gamma$ to the disk-bundle of radius $\varepsilon$ is positive definite and hence a Riemannian metric. If $r$ denotes the radial component of the fiber of the tubular neighbourhood $D(\mathcal{S})$ in $M$, we have
$|g-\gamma|_{\gamma}\leq C\cdot r^{2},$
since $g_{\alpha\beta}(x)=g_{\alpha\beta}(p)+x^{i}g_{\alpha\beta,i}(p)+O(r^{2})$ by the Taylor remainder estimate. The advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 5.
The proof now consists of investigating the difference in second fundamental form, mean curvature and normal vector when passing from the background metric $g+\mathrm{d}t^{2}$ to $\gamma+\mathrm{d}t^{2}$ and afterwards relating those terms for $\gamma+\mathrm{d}t^{2}$ to the corresponding ones for the Euclidean background metric $\delta+\mathrm{d}t^{2}$ in the fibers. Let us abbreviate
$A_{g}\coloneqq{}$ $A(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ $A_{\gamma}\coloneqq A(\Sigma_{\varrho},\gamma+\mathrm{d}t^{2})$ $A_{\delta}\coloneqq A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})$
$H_{g}\coloneqq{}$ $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ $H_{\gamma}\coloneqq H(\Sigma_{\varrho},\gamma+\mathrm{d}t^{2})$ $H_{\delta}\coloneqq H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})$
Furthermore, we denote by $\nu_{g}$, $\nu_{\gamma}$ and $\nu_{\delta}$ the respective upward unit normal vector fields of $\Sigma_{\varrho}$ or $\overline{\Sigma}_{\varrho}$. We have the following comparison formulae
$\|\nu_{g}-\nu_{\gamma}\|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $r^{2},\qquad$ $|A_{g}-A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot r^{2},$
where the latter identity follows from $A_{g}(e_{i},e_{j})=-\langle\nabla^{g+dt^{2}}_{e_{i}}e_{j},\nu_{g}\rangle_{g}$ and $|\Gamma^{m}_{ij}|_{g}\lesssim r^{2}$ for the Christoffel symbols $\Gamma_{ij}^{m}=\tfrac{1}{2}g^{ml}(\nabla^{\gamma}_{i}g_{lj}+\nabla^{\gamma}_{j}g_{li}+\nabla^{\gamma}_{l}g_{ij})$.
Since the difference of $\Sigma_{\varrho}$ and $M\times\{0\}$ is supported on $\{r\leq 2\sqrt{\varrho}\}$, we thus obtain
$|H_{g}-H_{\gamma}|={}|\operatorname{tr}(A_{g})-\operatorname{tr}(A_{\gamma})|\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
and therefore
$\bigl||A_{g}|_{g+\mathrm{d}t^{2}}^{2}$ $-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}\bigr|$
$\leq{}$ $\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\lesssim{}$ $\varrho^{2}\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\left(\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\right)$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
$|H_{g}^{2}-H_{\gamma}^{2}|\leq{}$ $|H_{g}-H_{\gamma}|^{2}+2|H_{\gamma}|\cdot|H_{g}-H_{\gamma}|$
$\lesssim{}$ $\varrho^{2}\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\bigl((1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr)$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}.$
Putting these together, we obtain
$\bigl|(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}\bigr|$ $\lesssim\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
If we write $\nu_{g}^{\intercal}\coloneqq\nu_{g}-\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\partial_{t}$ and $\nu_{\gamma}^{\intercal}\coloneqq\nu_{\gamma}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\partial_{t}$, then we observe
$\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{g},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\|g-\gamma\|_{\gamma}\lesssim\varrho}+\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\|\nu_{g}-\nu_{\gamma}\|_{\gamma+\mathrm{d}t^{2}}\lesssim\varrho}\lesssim{}\varrho$
$\left|\|\nu_{g}^{\intercal}\|_{g}-\|\nu_{\gamma}^{\intercal}\|_{\gamma}\right|$
$\qquad\leq{}\bigl|\|\nu_{g}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}\bigr|+\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\|\nu_{g}^{\intercal}-\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}+\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\qquad\leq{}\|\nu_{g}-\nu_{\gamma}\|_{g+\mathrm{d}t^{2}}+\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\varrho}+\underbrace{\bigl|\|\nu_{\gamma}^{\intercal}\|_{g+\mathrm{d}t^{2}}-\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\leq\|g-\gamma\|_{\gamma+\mathrm{d}t^{2}}\lesssim{}\varrho}$
$\qquad\lesssim{}\varrho$
Regarding the final summand in (6.18), we note that $\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\partial_{t},\partial_{t})=0$ and therefore, we obtain
$\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{g},\nu_{g})={}$ $\|\nu_{g}^{\intercal}\|_{g}^{2}\cdot\mathrm{Ric}(M,g_{M})\left(\frac{\nu_{g}^{\intercal}}{\|\nu_{g}^{\intercal}\|_{g}},\frac{\nu_{g}^{\intercal}}{\|\nu_{g}^{\intercal}\|_{g}}\right)$
$\lesssim{}$ $\|\nu_{g}^{\intercal}\|_{g}^{2}\lesssim\|\nu_{\gamma}^{\intercal}\|_{\gamma}+\varrho$
We note that the constant in the above inequality depends on the maximum of the Ricci-curvature of $g$ and hence, in particular on $g$, but is independent of $\varrho$.
As mentioned above, the advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles from Section 5. To make this precise, we first observe that the projection
$\pi:B_{R}(\mathcal{S})\times[0,1]\longrightarrow B_{R}^{k}(0)\times[0,1]\subseteq\mathbb{R}^{k}\times[0,1]$
satisfies $D\pi(\nu_{\gamma})=\nu_{\delta}$. Furthermore, the restriction of $\pi$ to the fibers is a radial isometry, so
$\|\nu_{\gamma}^{\intercal}\|_{\gamma+\mathrm{d}t^{2}}=\|\nu_{\delta}^{\intercal}\|_{\delta+\mathrm{d}t^{2}}\quad\text{and}\quad\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}=\langle\nu_{\delta},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}$
In particular, we immediately obtain $|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}|\lesssim{}\varrho$, proving the final estimate from Lemma 6.4.
If we fix a point $x\in\Sigma_{\varrho}$ and choose an adapted $\gamma$-orthonormal frame $\{e_{i},e_{a}\}$ of $\Sigma_{\varrho}$ at $x$, where $\{e_{i}\}$ are normal to $\mathcal{S}$ and $\{e_{a}\}$ tangent to $\mathcal{S}$, we see that $\{D\pi(e_{i})\}$ form normal coordinates at $\pi(x)\in\overline{\Sigma}_{\varrho}$. By definition of the second fundamental form,
$A_{\gamma}(e_{i},e_{j})={}$ $-\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla^{\delta+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\delta}\rangle=A_{\delta}(e_{i},e_{j}).$
Next, we note that $\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{a}}e_{j},e_{i}\rangle$ and $\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{a}}e_{b},e_{i}\rangle$ are components of the second fundamental form of the fibers of the tubular neighbourhood, hence they are independent of $\varrho$, and we get $\langle\nabla_{e_{a}}e_{j},e_{i}\rangle,\langle\nabla_{e_{a}}e_{b},e_{i}\rangle\lesssim 1$. Moreover, both $\nabla_{e_{a}}e_{j}$ and $\nabla_{e_{a}}e_{b}$ are orthogonal to $\partial_{t}$, so we obtain
$A_{\gamma}(e_{a},e_{j})={}$ $-\langle\nabla_{e_{a}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla_{e_{a}}e_{j},\nu_{\gamma}^{\intercal}\rangle$
$={}$ $-\sum_{i}\langle\nabla_{e_{a}}e_{j},e_{i}\rangle\langle\nu^{\intercal}_{\gamma},e_{i}\rangle\lesssim\|\nu^{\intercal}_{\gamma}\|=\|\nu^{\intercal}_{\delta}\|$
and $A_{\gamma}(e_{a},e_{b})\lesssim\|\nu_{\delta}^{\intercal}\|$, analogously. Thus, we have obtained
$|A_{\gamma}(e_{i},e_{a})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad|A_{\gamma}(e_{a},e_{b})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad\text{and}\quad A_{\gamma}(e_{i},e_{j})=A_{\delta}(e_{i},e_{j}).$
Using these, we immediately get
$\bigl||A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}\bigr|\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|^{2}\leq\|\nu_{\delta}^{\intercal}\|$
$|H_{\gamma}-H_{\delta}|\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|$
which, together with $|H_{g}-H_{\gamma}|\lesssim{}\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$ combines to
$|H_{g}-H_{\delta}|\lesssim\|\nu_{\varrho}^{\intercal}\|+\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}),$
proving the desired mean curvature estimate.
It remains to verify the scalar curvature estimate. We have
$|H_{\gamma}^{2}-H_{\delta}^{2}|={}$ $|2H_{\delta}\cdot(H_{\gamma}-H_{\delta})+(H_{\gamma}-H_{\delta})^{2}|$
$\lesssim{}$ $|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\cdot\|\nu_{\delta}^{\intercal}\|+\|\nu_{\delta}^{\intercal}\|^{2}$
$\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
where we used ${|H_{\delta}|\leq n|A_{\delta}|}\lesssim|A_{\delta}|$. Combining these, we get
$(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})-(H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2})\gtrsim{}$ $-\|\nu_{\delta}^{\intercal}\|(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}).$
Furthermore, there exists a constant $C$ such that
$|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\leq{}$ $\sqrt{|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}+C\|\nu_{\delta}^{\intercal}\|^{2}}$
$\leq{}$ $\sqrt{|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}+2\sqrt{C}\|\nu_{\delta}^{\intercal}\|\cdot|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+C\|\nu_{\delta}^{\intercal}\|^{2}}$
$\leq{}$ $|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+C\|\nu_{\delta}^{\intercal}\|$
and therefore $|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\lesssim\|\nu_{\varrho}^{\intercal}\|\leq 1$ and $1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim 1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}$. Note that $H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$ is precisely the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\delta})$, where $g_{\varrho,\delta}$ is induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$. Thus, we obtain
$(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\quad=(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})+(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})-(H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2})$
$\quad\gtrsim-\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\quad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
Putting all the pieces together, we therefore obtain:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\qquad=H^{2}_{g}-|A_{g}|^{2}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2\underbrace{\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho})}_{\lesssim\|\nu_{\delta}^{\intercal}\|+\varrho}$
$\qquad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})-\|\nu_{\delta}^{\intercal}\|+\varrho$
$\qquad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
which finishes the proof. ∎
Next, we compare $\Sigma_{\varrho}$ for different values of $\varrho$.
Proposition 6.6.
For every $\kappa\in(0,\varrho_{1}]$ there exists a smooth path of diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$, $\varrho_{0}\leq\varrho\leq\varrho_{1}$, such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$ and for $g_{\kappa,\varrho}$ the metric on $\Sigma_{\varrho}$ induced by $g_{\kappa}$ (see Theorem 6.1) the assignment
$\varrho\mapsto e^{\varrho}\cdot\Phi_{\varrho}^{\ast}(g_{\varrho,\kappa})$
is strictly monotonically increasing.
Proof.
Recall from (5.5) the family of reparametrizations
$\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\beta\left(s-\frac{1}{2}\right)ds$
and their lift to $\overline{\Sigma}_{\varrho}$ and $\Sigma_{\varrho}$ which, by slight abuse of notation, we both denote by $\Phi_{\varrho_{1},\varrho}$. First, we show that the proof in Proposition 5.4 still goes through in case the Euclidean background metric $\delta+\mathrm{d}t^{2}$ is replaced with $g+dt^{2}$ restricted to the fiber of the tubular neighbourhood $B_{R}(\mathcal{S})\times[0,1]$. We denote this restriction of $g_{M}+dt^{2}$ by $\hat{g}$. As in the proof of Proposition 5.4 we have
$\Phi_{\varrho_{1},\varrho}^{\ast}g_{\overline{\Sigma}_{\varrho},\hat{g}}(x,\tau)=\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))\bigl(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau)\bigr)\mathrm{d}\tau^{2}+g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}$.
Here $g_{S^{k-1}_{r},\hat{g}}$ is the metric induced by $\hat{g}$ on a sphere of radius $r$ where the radius is measured with respect to $\delta$. For $\varrho_{1}$ sufficiently small, the metric on the spheres $S^{k}_{r}$ is increasing as long as the radius $r\leq\varrho_{1}$ increases. Consequently,
$\partial_{\varrho}g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}\geq 0.$
Moreover, we have $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\lesssim\varrho^{1/4}\ll 1$. Since the derivatives of $\hat{g}$ are bounded, this yields as in the proof of Proposition 5.4
$\partial_{\varrho}\Bigl(\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))\bigl(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau)\bigr)\mathrm{d}\tau^{2}\Bigr)$
$\qquad\geq-\frac{1}{2}\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))\mathrm{d}\tau^{2}.$
Consequently,
$\partial_{\varrho}\left[e^{\varrho/2}\left(\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))\mathrm{d}\tau^{2}+g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}\right)\right]\geq 0$
and Proposition 5.4 still holds for the background metric $\hat{g}$. Next, we note that the other components of the metric $g_{\varrho}$ are independent of $\varrho$ except for the basepoint change caused by the reparametrization $\Phi$. Since $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\lesssim\varrho^{1/4}$, and since the geometry of both $M$ and $\mathcal{S}$ are bounded, the corresponding $\varrho$-derivatives of these remaining components are negligible small. Moreover, recall that the conformal factor $u_{\kappa}(x,t)=1+t\tilde{\kappa}\eta\left(\alpha^{-1}t-1\right)\leq 1+\varrho_{1}$ is fixed and again that $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\ll 1$. Hence, the result follows. Finally, we remark that the above estimates still hold in view of the smoothing construction performed at the end of Section 5. ∎
Finally, we deform the metric $e^{\varrho_{1}}\cdot g_{\kappa,\varrho_{1}}$ to a metric which is independent of $\kappa$ by linear interpolation of conformal factors:
Proposition 6.7.
For $\varrho_{1}>0$ and every $\kappa\in(0,\varrho_{1}]$, define the path $g_{s}$, $s\in[0,1]$ of metrics on $\Sigma_{\varrho_{1}}$ to be induced by the family
$s\mapsto e^{\varrho_{1}}\bigl((1-s)\cdot u+2s\bigr)^{\frac{4}{n-2}}\cdot g_{\varrho_{1}}$
of metrics on $M\times[0,1]$, where $u=(1+t\tilde{\kappa}\eta(\alpha^{-1}t-1))$ as in Theorem 6.1. Then, $g_{s}$ is strictly monotonically increasing in $s$ and satisfies
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq-(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4}).$
Proof.
Since $u_{s}=(1-s)\cdot u+2s$ is the linear path and $u\leq 2$, the monotonicity of $g_{s}$ follows immediately. Concerning the scalar curvature estimate, we recall the conformal transformation formula for scalar curvature we have
$\mathrm{scal}(\Sigma_{\varrho_{1}},e^{\varrho_{1}}u_{s}^{\frac{4}{n-2}}\cdot g_{\varrho_{1}})={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u_{s}-\frac{4(n-1)}{n-2}(\Delta_{g_{\varrho_{1}}}u_{s})\right)$
$={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\Bigl((1-s)\underbrace{\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u-\frac{4(n-1)}{n-2}(\Delta_{g_{\varrho_{1}}}u)\right)}_{=u^{(n+2)/(n-2)}\mathrm{scal}\left(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa}\right)}$
$\qquad\qquad\qquad\qquad\qquad+2s\cdot\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\Bigr).$
Note that
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa})\geq\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}\geq-|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|-\varrho_{1}^{1/4}.$
Furthermore, the proof of Theorem 6.1 reveals that the same inequalities hold for $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})$ as well, see (6.7), (6.10) and (6.12). Thus, we obtain
$\mathrm{scal}(\Sigma_{\varrho_{1}},$ $e^{\varrho_{1}}\cdot u_{s}^{\frac{4}{n-2}}\cdot g_{\varrho_{1}})$
$\geq{}$ $-e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\underbrace{\left((1-s)\cdot u^{\frac{n+2}{n-2}}+s\cdot 2\right)}_{\leq((1-s)\cdot u+s\cdot 2)^{{(n+2)}/{(n-2)}}}(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4})$
$\geq{}$ $-(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4})\qed$
Remark 6.8.
If the scalar curvature of $g_{M}$ is positive, we can arrange for $\mathrm{scal}(g_{s})$ to be positive in Proposition 6.7: If $\varrho_{1}$ is chosen such that $\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}>0$, then $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})$ and $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa})$ can be bounded by $\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}>0$. Since $(1-s)\cdot u^{\frac{n+2}{n-2}}+s\cdot 2\geq u_{s}$, we obtain
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq e^{-\varrho_{1}}u_{s}^{-\frac{4}{n-2}}(\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho^{1/4})>0.$
Remark 6.9.
Theorem 6.1 has an analogue for initial data sets $(M,g_{M},k)$, if we replace the scalar curvature by the dominant energy scalar $\mu-|J|$, where
$\mu=\frac{1}{2}(\mathrm{scal}_{g_{\Omega}}+\operatorname{tr}(k)^{2}-|k|^{2}),\qquad J=\operatorname{div}(k-\operatorname{tr}(k)g).$
That is, if we extend $k$ trivially to $M\times[0,1]$ and multiply it by a cutoff-function supported on $0,2a_{\varrho}]$, the dominant energy scalar of the hypersurface $\Sigma_{\varrho}$ only decreases by $\varrho^{1/4}$. If one goes through the proof, one encounters additional error terms for $\tfrac{1}{2}(\operatorname{tr}_{g_{M}}(k)^{2}-|k|^{2})$, $J$ and $\operatorname{tr}_{g_{\varrho}}(k)$, which have exactly the same type as the error terms already appearing for $\mathrm{scal}$ and $H$, see Remark 6.3.
7. Non-spin fill-ins with non-negative mean curvature
In this section we will prove Theorem (B) by induction on the number of surgeries required to transform $M$ into a sphere. The following is the main result needed for the induction step.
Proposition 7.1.
Let $(M,g_{M})$ be a closed oriented Riemannian manifold. Let $\widetilde{M}$ be a closed oriented manifold such that $\widetilde{M}$ is obtained from $M$ by performing an oriented surgery in codimension at least $3$. Then there exists
  • a Riemannian metric $g_{\widetilde{M}}$ on $\widetilde{M}$,
  • a smooth bordism $T_{1}$ from $M$ to $\widetilde{M}$,
  • a constant $C\leq 0$,
with the following property: Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Then there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on
$\Omega_{1}=\Omega\bigcup_{\partial\Omega\approx M}T_{1}$
such that the following holds:
  • $g_{\Omega_{1}}=g_{\Omega}$ outside an arbitrarily small neighborhood of $T_{1}\subset\Omega_{1}$,
  • $g_{\Omega_{1}}$ induces the metric $g_{\widetilde{M}}$ on $\partial\Omega_{1}=\widetilde{M}\times\{1\}$,
  • $H_{g_{\Omega_{1}}}>0$ along $\widetilde{M}\times\{1\}$,
  • $\mathrm{scal}_{g_{\Omega_{1}}}\geq\min\{\sigma,C\}-1$,
  • $\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}<2\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{\partial\Omega_{1}}}$.
Proof.
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 3$, be the surgery datum for passing from $M$ to $\widetilde{M}$. Let $\mathcal{S}:=\psi(S^{n-k}\times 0)\subset M$ be the corresponding surgery sphere. The trace of $\psi$ is defined as
$T=T(\psi)\coloneqq(M\times[0,1])\cup_{\psi}(D^{n-k+1}\times D^{k}).$
It yields an oriented bordism from $M$ to $\widetilde{M}$. For simplicity, we will suppress the map $\psi$ from now on in our notation.
Choose a metric $h$ on $T$ that restricts to $g_{M}+\mathrm{d}t^{2}$ on $M\times[0,1]\subset T$. The metric $h$ stays fixed for the remainder of the proof. Choose $R\in(0,\frac{2}{\sqrt{10}}]$ smaller than the normal injectivity radius of $h$ along the core of the surgery
$\mathcal{S}\times[0,1]\;\bigcup\;D^{n-k+1}\times\{0\}\subset T.$
In particular, we assume that $T$ contains the tubular $R$-neighborhood of its core.
Let us choose $\alpha\in(0,\tfrac{1}{3}]$ and let $\varrho_{1}\in(0,\min(R^{2}/4,\alpha/2))$ be the first constant obtained in Theorem 6.1 for $\mathcal{S}\subset(M,g_{M})$. For $0<\varrho\leq\varrho_{1}$, let $B_{\varrho}(D^{n-k+1}\times 0)\subset T$ be the $\varrho$-tubular neighborhood of $D^{n-k+1}\times 0$ in $T$ and let $S_{\varrho}(D^{n-k+1}\times 0)\subset B_{\varrho}(D^{n-k+1}\times 0)$ be the normal sphere bundle of radius $\varrho$ around $D^{n-k+1}\times 0$. Let $h_{S_{\varrho}}$ be the metric on $S_{\varrho}(D^{n-k+1}\times 0)$ induced by $h$.
By decreasing $\alpha$ and hence $\varrho_{1}$, we may assume that for every $0<\varrho\leq\varrho_{1}$, the metric $h_{S_{\varrho}}$ has positive scalar curvature and $S_{\varrho}(D^{n-k+1}\times 0)\subset B_{\varrho}(D^{n-k+1}\times 0)$ has positive mean curvature. For $0<\varrho\leq\varrho_{1}$, let
(7.1) $\Psi_{\varrho_{1},\varrho}\colon S_{\varrho^{4}_{1}}(D^{n-k+1}\times 0)\to S_{\varrho^{4}}(D^{n-k+1}\times 0)$
be the diffeomorphism induced by radial dilatation in the normal bundle of $\mathcal{S}$. By further decreasing $\alpha$, we can assume that the smooth path $(0,\varrho_{1}]\ni\varrho\mapsto\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})$ satisfies
$\partial_{\varrho}\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})\geq-\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}}).$
This implies that the smooth path
(7.2) $\varrho\mapsto e^{\varrho}\cdot\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})$
is strictly monotonically increasing.
Let $\sigma\in\mathbb{R}$ and let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ with $\mathrm{scal}(g_{\Omega})\geq\sigma$ and $H_{g_{\Omega}}\geq 0$ on $\partial\Omega$. By a small deformation of $g_{\Omega}$ near $\partial\Omega\subset\Omega$, leaving the induced metric on $\partial\Omega$ unchanged and replacing $\sigma$ with $\sigma-\tfrac{1}{2}$, we can assume that $H_{g_{\Omega}}>0$ along $\partial\Omega$, see [BH23, Proposition 3.8]. By compactness of $\partial\Omega$, there exists a constant $\kappa=\kappa(\Omega,g_{\Omega})>0$ such that $H_{g_{\Omega}}\geq\kappa$.
We choose $\varrho_{0}=\varrho_{0}(\kappa)$ as the second constant from Theorem 6.1. By further decreasing $\varrho_{0}$ we may assume that $(1-\alpha)\kappa>\varrho_{0}^{1/4}$. By choosing $\varrho_{0}$ even smaller, we may assume that
(7.3) $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}\leq 2\int_{\partial\Omega\setminus B_{4\sqrt{\varrho_{0}}}(\mathcal{S})}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}.$
Note that $\varrho_{1}$ does not depend on the fill-in $(\Omega,g_{\Omega})$, but $\varrho_{0}$ does.
For the given $\kappa$, consider the metric $g_{\kappa}$ on $M\times[0,1]$ defined in Theorem 6.1,
$g_{\kappa}=u^{\frac{4}{\,n-2\,}}(g_{M}+dt^{2}),\qquad u(x,t)=u_{\kappa}(x,t)=1+t\widetilde{\kappa}\eta\left(\alpha^{-1}t-1\right).$
Recall that $H_{g_{\kappa}}=-\kappa$ along $M\times\{0\}$ with respect to the exterior normal. By Theorem 6.1 and our choice of $\varrho_{0}$, we have
$H(\Sigma_{\varrho_{0}}\cap(M\times[0,1]),g_{\kappa})>(1-\alpha)\kappa-\varrho_{0}^{1/4}>0.$
Since $g_{\kappa}$ restricted to $M\times[\alpha,1]$ equals $g_{M}+\mathrm{d}t^{2}$, the union of $g_{\kappa}$ and $h|_{T\setminus(M\times[0,1])}$ defines a smooth metric $g_{T}$ on $T$. Here, $h$ is the fixed metric on $T$ chosen before. Furthermore, as $g_{\kappa}\stackrel{{\scriptstyle\kappa\to 0}}{{\longrightarrow}}g_{M}+\mathrm{d}t^{2}$ in the $C^{\infty}$-topology, $\mathrm{scal}(T,g_{T})$ is bounded from below by a constant $C(M,g_{M},h)$, which is independent of $\kappa$.
For $0<\varrho\leq\varrho_{1}$ we now define the hypersurface $\Sigma_{\varrho}\subset T$ as follows: The restriction $\Sigma_{\varrho}\cap(M\times[0,1])$ is given by the manifold $\Sigma_{\varrho}\subset M\times[0,1]$ constructed in the beginning of Section 6. Since $a_{\varrho}<2\varrho\leq\alpha$, we see that $\Sigma_{\varrho}\cap\left(M\times[\alpha,1]\right)=S_{\varrho^{4}}(\mathcal{S})\times[\alpha,1]$, where $S_{\varrho^{4}}(\mathcal{S})\subset M$ is the normal sphere bundle of radius $\varrho^{4}$ with respect to the metric $g_{M}$. We extend this by the normal sphere bundle $S_{\varrho^{4}}(D^{n-k+1}\times 0)$ around the upper part $D^{n-k+1}\times\{0\}\subset T$. Let $\Sigma_{\varrho}$ be the resulting smooth manifold and denote the induced Riemannian metric by $g_{\varrho,\kappa}$, see Fig. 11. Since $g_{T}$ is given by the union of $g_{\kappa}$ and $h$, our choice of $\varrho_{1}$ and Theorem 6.1 ensures that the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\kappa})$ is bounded below by
(7.4) $\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}$
for all $0<\varrho\leq\varrho_{1}$ and that the mean curvature of $\Sigma_{\varrho_{0}}$ is positive.
[…]
Figure 11. Schematic depiction of the quantitative surgery construction.
By a smooth deformation of $g_{\Omega}$ we can assure that for some $\delta>0$, the restriction of $g_{\Omega}$ to the tubular neighborhood $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\times(-\delta,0]\subset\Omega$ of $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\subset\partial\Omega$ can be smoothly extended by the metric $g_{\kappa}$ onto $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\times(-\delta,1]$. This can be achieved without altering the induced metric $g_{\varrho,\kappa}$ on $\Sigma_{\varrho}$, $0<\varrho\leq\varrho_{1}$ or the mean curvature of $g_{\Omega}$ on $M\setminus B_{4\sqrt{\varrho_{0}}}(\mathcal{S})$, while preserving $H_{g_{\Omega}}\geq\kappa$. This deformation is described in detail in Section 2, below (2.3), where $U_{i}\subset V_{i}\subset Z_{i}$ correspond to $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\subset B_{3\sqrt{\varrho_{0}}}(\mathcal{S})\subset B_{4\sqrt{\varrho_{0}}}(\mathcal{S})$.
The manifold $\Sigma_{\varrho_{0}}$ thus bounds an oriented manifold of the form $\Omega_{0}=\Omega\cup T_{0}$, which is diffeomorphic to $\Omega\cup T$ and $g_{T}$ restricts to a smooth metric $g_{\Omega_{0}}$ on $\Omega_{0}$. Hence, we obtain a fill-in $(\Omega_{0},g_{\Omega_{0}})$ of $\Sigma_{\varrho_{0}}$. Since on $M\setminus B_{4\sqrt{\varrho_{0}}}$ the metrics $g_{M}$ and $g_{\varrho_{0},\kappa}$ agree, since $\Sigma_{\varrho_{0}}\subset(\Omega_{0},g_{\Omega_{0}})$ has positive mean curvature and using (7.3), we have
(7.5) $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}\leq 2\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}\mathrm{dvol}_{g_{\partial\Omega_{0}}}.$
In the next step, the right hand side will be bounded above by an application of Proposition 4.2. Let us abbreviate $\Sigma\coloneqq\Sigma_{\varrho_{1}}$ and consider the metric
$g_{\Sigma}\coloneqq e^{\varrho_{1}}\cdot 2^{4/(n-2)}\iota^{*}\big((g_{M}+\mathrm{d}t^{2})\cup h\big),$
where $\iota\colon\Sigma\hookrightarrow T$ denotes the inclusion. Note, that $g_{\Sigma}$ only depends on $(M,g_{M})$ and $\psi$ and that $\Sigma\approx\widetilde{M}$.
We claim that $g_{\varrho_{0},\kappa}\in\mathscr{R}_{g_{\Sigma},s}(\Sigma)$ (see Proposition 4.2) for
$s=-\max(|\mathrm{scal}(M,g_{M})|+\varrho^{1/4})\geq-\max(|\mathrm{scal}(M,g_{M})|-1.$
There exists a smooth path of diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$, $\varrho_{0}\leq\varrho\leq\varrho_{1}$, such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}$, $\Phi_{\varrho}|_{M\setminus B_{2\sqrt{\varrho_{1}}}}=\operatorname{\mathrm{id}}$ and such that the path $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}g_{\varrho,\kappa}$ is strictly increasing. On $\Sigma_{\varrho_{1}}\cap(M\times[0,1])$ these diffeomorphisms are constructed in Proposition 6.6. Since the restrictions of these diffeomorphisms to $\Sigma_{\varrho_{1}}\cap(M\times[\tfrac{3}{4},1])$ dilate the normal sphere bundle around $\mathcal{S}\subset M$ from radius $\varrho_{1}^{4}$ to $\varrho^{4}$ (this follows form the choice of $\beta$ in the proof of Proposition 6.6), we can take the union with the dilatations $\Psi_{\varrho_{1},\varrho}$ from (7.1) to define $\Phi_{\varrho_{1},\varrho}$ on $\Sigma_{\varrho}$. By construction, we get $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}$, and the path $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}g_{\varrho,\kappa}$ is strictly increasing by Proposition 6.6 and the choice of $\alpha$, see (7.2). This path ends at the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma$. Finally, we
  • rescale the conformal factor in $e^{\varrho_{1}}\cdot u^{\tfrac{4}{n-2}}(g_{M}+dt^{2})$, which induces the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma\cap(M\times[0,1])$, along the path in Proposition 6.7,
  • rescale the conformal factor in $e^{\varrho_{1}}\cdot h$, which induces the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma\setminus(M\times[0,1])$, along the path $s\mapsto e^{\varrho_{1}}(1+s)^{\tfrac{4}{n-2}}$.
This results in a strictly increasing smooth path of metrics on $\Sigma$ starting at $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ and ending at $g_{\Sigma}$, while maintaining the uniform lower curvature bound $s=-\max(|\mathrm{scal}(M,g_{M})|+\varrho^{1/4})$. This proves our claim that $g_{\varrho_{0},\kappa}\in\mathscr{R}_{g_{\Sigma},s}(\Sigma)$. Let $g_{\Omega_{1}}$ be the smooth Riemannian metric on $\Omega_{1}=\Omega_{0}\cup_{\widetilde{M}}(\widetilde{M}\times[0,1])$ constructed in Proposition 4.2. We have
$\mathrm{scal}_{g_{\Omega_{1}}}\geq\min\{\sigma,s,C(M,g_{M},h)\}-1.$
Put $T_{1}=T_{0}\cup(\widetilde{M}\times[0,1])$ and $C:=\min\{0,s,C(M,g_{M},h)\}$. We obtain
$\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}\mathrm{dvol}_{g_{\partial\Omega_{0}}}<\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}\mathrm{dvol}_{g_{\partial\Omega_{1}}}.$
Hence the claim follows from (7.5). ∎
Proof Theorem (B).
Let $(M,g_{M})$ and $\sigma\in\mathbb{R}$ be as in Theorem (B). Put $n:=\dim M$. As before, we can assume $n\geq 5$. Since $M$ carries only finitely many spin structures, we can work with a fixed spin structure on $M$.
Since each fill-in of $M$ induces a fill-in of the spin manifold $M\sqcup\overline{M}$, where $\overline{M}$ carries the reverse spin structure, we can furthermore assume that the spin manifold $M$ is spin null-bordant. Let $B$ be a compact spin manifold with $\partial B=M$.
Removing a small open disc from the interior of $B$, we obtain a compact spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$, $\pi_{1}(W)$ and $\pi_{2}(W)$ by spin surgeries in the interior of $S$ (using $n\geq 5$ and that $W$ is spin), we may assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $2$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $3$. Hence, the sphere $S^{n}$ is obtained from $M$ by oriented surgeries of codimension at least $3$.
[…]
Remark 7.2.
The construction in Proposition 7.1 allows to impose quantitative restrictions onto the classical surgery step of Gromov–Lawson [GL80]. More specifically, let $\varepsilon>0$, let $\sigma\in\mathbb{R}$, and let $(M,g_{M})$ be a Riemannian manifold satisfying $\mathrm{scal}_{g}>\sigma$. Let $\widetilde{M}$ be obtained from $M$ by surgery along an embedding $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 3$. Let $\breve{M}:=M\setminus\mathrm{im}(\psi)$. Choosing $\alpha$ in the proof of Proposition 7.1 small enough, working with a metric $h$ on $T$ of sufficiently small volume and diameter, and working with $M\times[0,\alpha]$ instead of $M\times[0,1]$, we obtain a Riemannian metric $g_{\widetilde{M}}$ on $\widetilde{M}$ with the following properties (compare Eq. 7.4):
  • $\mathrm{scal}_{g_{\widetilde{M}}}>\sigma$,
  • the restriction of $g_{\widetilde{M}}$ to $\breve{M}$ is equal to $g_{M}$,
  • the volume and diameter of the restriction of $g_{\widetilde{M}}$ to $\widetilde{M}\setminus\breve{M}$ are bounded above by $\varepsilon$.
8. A new quasi-local mass
Our results have applications in Physics. In 1982, Penrose [Pen82] listed important open problems in General Relativity. Finding a suitable notion of quasi-local mass was listed first.
In Newtonian gravity, the mass density function $\rho$ satisfies Poisson’s equation
$\Delta V=4\pi\rho,$
where $V$ denotes the gravitational potential. The total mass contained in a region $\Omega$ is then given by
$m(\Omega)=\int_{\Omega}\rho\,d\mu=\frac{1}{4\pi}\int_{\partial\Omega}\nabla_{\nu}V\,d\mu,$
where $\nabla_{\nu}$ denotes the outward normal derivative. However, in general relativity, the geometric nature of spacetime makes it impossible to express the mass as an integral of a local density depending only on the metric and its derivatives. To overcome this difficulty, two main approaches have been developed:
The first is to define the total mass for an isolated system, namely the ADM mass, associated with an asymptotically flat manifold $(N,g_{N})$. Here, asymptotically flat initial data sets $(N,g_{N},k)$ model isolated gravitational systems, such as stars or galaxies. The mass $m(N,g_{N},k)$ is well-defined and enjoys many desirable properties such as positivity and rigidity. However, it has the drawback that it applies only to entire non-compact manifolds, rather than to bounded regions within them. For instance, if $(N,g_{N},k)$ models two black holes contained in its interior, the ADM mass only measures the total mass of the system, without distinguishing the contributions of the individual black holes.
In order to define the mass of a finite region–known as quasi-local mass–various proposals have been made, including the Hawking mass [Haw68, HI01], the Bartnik mass [Bar89], the Brown–York mass [BY93, ST02], and the Wang–Yau mass [WY09]. Each of these notions has attractive geometric features but also important limitations. A striking common feature is that all of them are particularly well-suited for spherical topology: The Hawking mass often underestimates the true mass and is most effectively used together with inverse mean curvature flow, whose monotonicity holds only for $M\cong S^{2}$. The Brown–York and Wang–Yau masses rely on isometric embeddings into Euclidean or Minkowski space, and classical embedding results such as the Weyl theorem require positive Gauss curvature. The Bartnik mass, perhaps the most mathematically natural proposal, has so far been explicitly computed only for spherical boundaries.
This contrasts with the increasing importance of spacetimes with more complicated topology including Emparan–Reall’s black ring [ER02], the Kottler metric [LN15], and the Horowitz–Myers soliton [BH24].
Motivated by [MM17] we define a new notion of quasi-local mass which has no such topological shortcomings. For this purpose, let $(N,g_{N})$ be a complete oriented manifold with $\mathrm{scal}_{g_{N}}\geq\sigma$, $\sigma\in\mathbb{R}$, and let $M\subset N$ be a mean convex hypersurface and let $g_{M}$ be the metric on $M$ induced by $g_{N}$. We set
$\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))=\Lambda(M,g_{M},\sigma)-\int_{M}H_{g}{\rm dvol}_{g_{M}},$
where we define
[…]
to be the “optimal” constant for which Theorem (B) holds. This should be compared with Bray’s inner mass [Bra01] which also relies on fill-ins. Theorem (B) immediately yields
Corollary 8.1.
Suppose that $\operatorname{dim}(N)\leq 7$ and $M$ is spin. Then the quasi-local mass $\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))$ is well-defined and $\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))\geq 0$.
Here $\tfrac{1}{2}\sigma$ corresponds to the cosmological constant. We point out that the ADM mass for asymptotically flat manifolds has been successfully extended to the setting $\sigma<0$ [Wan01], though there is no analogue for $\sigma>0$ [BMN11]. On the other hand $\textbf{m}_{\sigma}$ is well-defined and non-negative for any choice of $\sigma\in\mathbb{R}$.
When $n=2$ and $M\cong S^{2}$, this recovers the notion introduced by Mantoulidis–Miao [MM17, Theorem 1.5]. As in their setting, we obtain:
Corollary 8.2.
Suppose that $n=2$, $M=S^{2}$ and $K(M)>0$ for the Gaussian curvature. Then
$\textbf{m}_{0}(M)=m_{BY}(M)$
for the Brown–York mass. In particular, we have in this case $\textbf{m}_{0}(M)=0$ if and only if $\Omega\subseteq\mathbb{R}^{3}$.
Appendix A Collection of auxiliary results
This collection of elementary results is used to construct handles in Euclidean space around points.
Lemma A.1.
Let $0<\varrho<1$ and consider on the interval $\varrho^{4},1]$ the function
$\xi_{\varrho}(r)=-\frac{4}{3}\varrho\left(\sqrt{r}-\varrho^{2}\right)^{3/2}-4\varrho^{3}\sqrt{\sqrt{r}-\varrho^{2}}.$
Then the derivatives satisfy
$\xi_{\varrho}^{\prime}(r)$ $=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},$
$4\xi_{\varrho}^{\prime\prime}(r)$ $=-\frac{1}{r}\,\xi_{\varrho}^{\prime}(r)\left(1+(\xi_{\varrho}^{\prime}(r))^{2}\right).$
The inverse function $\xi_{\varrho}^{-1}$ is given by
$\xi_{\varrho}^{-1}(t)=\left(\varrho^{2}+4\varrho^{2}B^{2}\right)^{2},\qquad B=\sinh\left(\frac{1}{3}\operatorname{arsinh}\left(-\frac{3t}{8\varrho^{4}}\right)\right),$
and its derivatives are
$\dot{\xi}_{\varrho}^{-1}(t)$ $=-2B,$
$\ddot{\xi}_{\varrho}^{-1}(t)$ $=\frac{1}{4\varrho^{4}(1+4B^{2})}.$
[…]
Proof.
A direct computation yields the derivative identities for $\xi_{\varrho}$. For the inverse, we set $r=\varrho^{4}(1+4B^{2})^{2}$ and compute
$\xi_{\varrho}(r)=-\frac{32}{3}\varrho^{4}B^{3}-8\varrho^{4}B.$
Hence, the triple angle formula
$\sinh(3u)=3\sinh(u)+4\sinh^{3}(u)$
implies the claim that $\xi_{\varrho}^{-1}$ inverts $\xi_{\varrho}$. Differentiating $B(t)=\sinh\left(\frac{1}{3}\operatorname{arsinh}\left(-\tfrac{3t}{8\varrho^{4}}\right)\right)$ gives for $u=\left(\frac{1}{3}\operatorname{arsinh}\left(-\frac{3t}{8\varrho^{4}}\right)\right)$
$\dot{B}(t)=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\sqrt{1+\left(\tfrac{3t}{8\varrho^{4}}\right)^{2}}}$
$=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\sqrt{1+\sinh^{2}(3u)}}$
$=$ $-\frac{1}{8\varrho^{4}}\cosh(u)\frac{1}{\cosh(3u)}.$
Using the triple angle formula
$\cosh(3u)=4\cosh^{3}u-3\cosh u=\cosh(u)(4(1+\sinh^{2}u)-3),$
we obtain
$\dot{B}(t)=-\frac{1}{8\varrho^{4}(1+4B^{2})},$
from which the derivative identities for $\dot{\xi}_{\varrho}^{-1}$ and $\ddot{\xi}_{\varrho}^{-1}$ follow directly. ∎
Remark A.2.
The function $B$ can also be expressed more explicitly with the help of the formula
$2\sinh\left(\tfrac{1}{3}\operatorname{arsinh}u\right)=\left(u+\sqrt{u^{2}+1}\right)^{1/3}-\left(\sqrt{u^{2}+1}-u\right)^{1/3}$
which follows from repeated usage of the triple angle formulas.
The next result is well-known and we will omit its proof.
Lemma A.3.
Let $\ell:\mathbb{R}^{n}\to\mathbb{R}$ be a spherically symmetric function, i.e. $\ell(x)=\ell(r)$ where $r=|x|$. Then the principal curvatures of the graph of $\ell$ are
$\lambda_{1}=-\frac{\ell^{\prime\prime}}{(1+\ell^{\prime 2})^{3/2}},\qquad\lambda_{2}=\dots=\lambda_{n}=-\frac{\ell^{\prime}}{r\sqrt{1+\ell^{\prime 2}}}$
where $\lambda_{1}$ is the principal curvature in radial direction.
Appendix B Example of unbounded total mean curvature
Let $n\geq 2$. Then, for every $K>0$, there exists a metric $g_{K}$ on $S^{n}$ satisfying
  • $\mathrm{vol}(S^{n},g_{K})\leq K^{-1}$,
  • $\mathrm{diam}(S^{n},g_{K})\leq K^{-1}$,
  • there is a fill-in of $(S^{n},g_{K})$ with nonnegative scalar curvature, positive mean curvature and total mean curvature larger than $K$.
Hence, $\Lambda(S^{n},g_{K},0,0)\geq K$.
By taking connected sums with $g_{K}$, we can implant these examples into an arbitrary Riemannian manifold. More precisely, let $(M,g_{M})$ be a Riemannian manifold and let $(\Omega,g_{\Omega})$ be a fill in with nonnegative scalar and mean curvature and
$\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}=\Lambda.$
Then, for $K>0$, there exists a metric $g_{M,K}$ on $M$ and a fill-in $(\Omega_{\varepsilon},g_{\Omega,\varepsilon})$ of $(M,g_{M,K})$ with nonnegative scalar curvature and positive mean curvature, which satisfies
  1. (i)
    $g_{M}|_{U}=g_{M,K}|_{U}$ for a subset $U\subset M$ with $\mathrm{vol}(U)=\mathrm{vol}(M)-K^{-1}$,
  2. (ii)
    $\mathrm{vol}(M,g_{M,K})\leq\mathrm{vol}(M,g_{M})+K^{-1}$,
  3. (iii)
    $\mathrm{diam}(M,g_{M,K})\leq{\rm diam}(M,g_{M})+K^{-1}$,
  4. (iv)
    $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}=\Lambda+K$.
To construct the metric $g_{K}$ on $S^{n}$, let $\ell\coloneqq(2c(n)K)^{2n-1}$ and $r\coloneqq(2c(n)K)^{-2}$, where $c(n)=\mathrm{vol}(S^{n},g_{\circ})$ is a dimensional constant. By increasing $K$, we may assume that $\ell\in\mathbb{N}$ and $20r\leq K^{-1}$. We embed $\ell+1$ disjoint spheres of radius $r$ into $\mathbb{R}^{n+1}$, such that one of them is centered at the origin and the straight paths connecting a given embedded sphere with this “center sphere” does not intersect any other sphere. We perform connected sums to connect all outer spheres to the center-sphere by straight tubes, and we may assume that any connecting tube has volume smaller than $\tfrac{1}{2\ell}K^{-1}$. Let us denote the resulting Riemannian manifold by $(X,g_{X})$. We observe, that $X\approx S^{n}$. By [LM84, Theorem 3.1], we may assume that $X$ has positive mean curvature with respect to the euclidean metric, since $n\geq 2$. Hence, the interior of $X$ constitutes a flat fill-in $(\widetilde{\Omega},g_{\rm flat})$ with positive mean curvature. A visualisation can be found in Fig. 12.
Figure 12. The metric $g_{K}$ on $S^{n}$.
Let us now estimate the volume and the total mean curvature of this fill-in. For the total mean curvature, we can restrict to the respective outer hemispheres, where the mean curvature is constantly equal to $n/r$. Hence, we obtain:
(B.1) $\begin{split}\int_{\partial\widetilde{\Omega}}H_{g_{\rm flat}}\mathrm{dvol}_{g_{\partial\widetilde{\Omega}}}\geq{}&\ell\int_{S^{n}_{r,+}}H_{g_{\rm flat}}\mathrm{dvol}_{g_{\partial\widetilde{\Omega}}}\geq\frac{1}{2}nc(n)\ell r^{n-1}\\ \geq{}&\frac{1}{2}nc(n)\bigl(2c(n)K\bigr)^{2n-1}\bigl(2c(n)K\bigr)^{-2(n-1)}\geq K.\end{split}$
The volume on the other hand satisfies:
$\mathrm{vol}(X)$ $\leq\ell\left(\mathrm{vol}(S^{n}_{r})+\frac{1}{2\ell}K^{-1}\right)=c(n)\bigl(2c(n)K\bigr)^{2n-1}\bigl(2c(n)K\bigr)^{-2n}+\frac{1}{2}K^{-1}$
$\leq\frac{1}{2}K^{-1}+\frac{1}{2}K^{-1}=K^{-1}$
Note, that while we cannot put the outer spheres onto a circle of arbitrarily small radius, we can still cut away portions of the tubes as non-embedded Riemannian manifolds, resulting in tubes of length less than $r\leq\tfrac{1}{12}K^{-1}$. The resulting manifold and its fill-in might not be isometric to a Riemannian submanifold of $(\mathbb{R}^{n+1},\delta)$, but the mean curvature will still be positive, the scalar curvature will remain flat and estimate (B.1) is still valid, as we do not alter the outer hemispheres by this procedure. The volume may have further decreased and the diameter of the resulting metric $g_{K}$ is smaller than $3\pi\cdot r+4\cdot r\leq 12\cdot r\leq K^{-1}$.